A Complete Bibliography of the Journal of Statistical Physics: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 1.04

Title word cross-reference

1/f [But72, Tun76c]. 32 [Run76]. _{(αbyβ)} [Pie72]. A ⇀ ↽ B [MK75]. F [Bax73b]. H [Hub76, Rés78]. μ [Onu78b]. n [AZ74, PT77]. ν [SM76]. T_{c} [SKW71]. θ ≥ T [Sch74]. v ≥ 3 [Roe78].

-dimensional [SM76]. -model [Bax73b]. -space [Onu78b]. -theorem [Hub76, Rés78]. -vector [AZ74, PT77]. -vertex [Run76].

1978 [Wid79].

2nd [May78].

34th [Ano76p]. 35th [Leb76]. 36th [Ano77o]. 37th [Ano77p]. 38th [Ano78o]. 39th [Ano78p].

40th [Ano78q]. 41st [Ano79n].

A. [Son79]. Above [AS73]. absence [GP76b]. absorbing [LK79, MS73]. absorption [Nis76]. ac [Tun76a]. Accuracy [Bra74b, Coh70]. acoustic [MT73, SJ70]. Action [Bis72]. Action-angle [Bis72]. adaptation [FG74a, FG74b]. added [Sni76]. Addendum [Blu78a, Fer74]. additive [BGV74]. adiabatic [Ale75b]. Adsorbate [Run73]. Adsorption [IF79, SP77]. Algorithms [HKK79]. allowing [NW74]. alloy [Bor74]. alternative [Gat76, Var79]. among [LK79]. Analysis [Fox77, BNW78, CGMW77, Mat77, Oon72]. Analytic [GKPR78, Nor73, Lin74b]. analytical [HS78]. Analyticity [RL76]. angle [Bis72]. angular [SLD76, TS70]. anharmonic [LLL77, Sch76]. animals [Sta78]. anisotropic [ML72, ML74]. Announcement [Ano74a, Ano76a, Ano76b, Ano77a, Ano78a, Ano72h, Ano72i, Ano73y, Ano76o, Ano77n, Ano78n, Ano79m]. Announcements [Ano76c]. Anomalous [Suz77a]. answers [BG71, Sha72, She70, Swe74, Tyk71, Whe72]. antiferromagnetic [vdBCWW78, vdBCGW79]. any [HSS77]. appearing [DB71]. applicable [KM75]. Application [CGWE73, FG76, Kre75, VR73a, WD75, HR71a, Mat78, ML77, Rai78, SJ70, Tun76a, VR73b, BN73, De 75, vK78]. Applications [MR75a, MR76, RH72, Cho75, Leu70, PD78]. applied [MA77, Tsa79, YK76]. Approach [AZ74, HR71b, Akc77, Ale72, BW77, BN73, CP75, CM74, CS76, GCdR77, Gil77, Grm71, Kei70a, Kei70b, RS75, Spo74, YOK77, dRGC78]. Approximate [BGV74, Blo70, Gro73b, Gro73c, RL75, Rés75, RL77, HR76, Wei72, Bud76, Bud77, dRGC78]. approximation [Ado71, AM79, BGV74, Blu78a, Blu78b, Bra74b, Egg74, FG74b, GPE79, HS78, Lee72, Rés75, Sch74]. approximations [Bax78, Coh70, KHY76a, KHY76b, Kih75, Rei72, Tsa79]. aqueous [BN73]. arbitrary [Blu78b, MS74, Per77, Pol72, Rid77, Sch75b, Tun76b]. area [BS72a]. arrays [WP78]. assumption [Sta78]. asymmetric [GKPR78]. Asymptotic [EHvL76a, EHvL76b, GPE79, Mon73, Shl74, Tun74, Hal76, Hil74, HMMH77, KS78b, Tun75]. Asymptotically [Bax73a]. attractor [Sim79, Son79]. augmented [DO78]. autocorrelation [CJSJT73, Cha70, FF73, LS72, Rés75, SLD76]. average [SSL76]. axiomatic [ST74].

balance [Ale77, GLS79]. band [Pli74, Sch75b]. Bandwidth [Coh75, HR71a]. barriers [De 75]. based [Ler71, SS78, Tun76c, Var79]. basis [Kle74, TR73, TS75, Vst76]. Bayesian [Sýk74]. BBGKY [FSV74, MMT78, RS77, Var79]. BCS [BM78, BM79]. Bead [Cha72]. Bead-spring [Cha72]. beam [Cho75]. behavior [Bar74, BW77, BP78, CM74, Coo79, EHvL76a, EHvL76b, GC73, KS78b, Lav78, MC74, McL76, MA77, Mon73, NM78, PW79, SLD76, Wal76, Wie73, YY79]. below [AS73]. Bethe [Egg74, RL78, SK76]. between [HH76, McL78]. bias [NW74]. bifurcations [FT79, McL76, Son79]. bimodal [PM78]. binary [BGCP73, Bor74, BFGC73, Fel72, GCFB73, Piñ74, TD71]. biomembrane [Wie75]. biophysical [LT79]. birth [Mat77]. bistable [CCR79, vK77]. Bitter [Bru70]. Block [Abr78a]. Board [Ano72d, Ano73a, Ano73c, Ano73d, Ano73h, Ano73i, Ano74c, Ano75h, Ano71a, Ano72a, Ano72b, Ano72c, Ano72e, Ano72f, Ano73b, Ano73e, Ano73f, Ano73g, Ano73j, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75i, Ano75j, Ano75k, Ano75l]. bodies [KM75]. body [BGV74, BG75, BF73]. Boltzmann [ASJS71, ABO74, BBO75, CGZ79, DL79, GVZ79, Kan77, MM78, Pin75, Ray79, Spo79, TS70, TS75]. bond [Fis78, HKK79, WP78]. bonded [AH72, AH75]. Book [FTM70, Gri70, MD79, Sat70, Sch72, TD78, Wie79, vK78, Bru78, May78]. bord [Ano74b]. Bose [FtHW70, Gir78, GIB74, Hal75, RS73, Roe78]. bound [vdBCGW79]. boundaries [Lin74b, MS73]. Boundary [Mer79, BLP79, MR75b, MMS77, VH74, VR73a]. Bounds [Roe78, Gro79]. breakdown [KvLH77]. breaking [CPC79]. broken [Mat74, Mat75]. Brownian [BZ71, CM79, CFTA79, De 75, GZW75, HML73, Hyn74, Iva73a, Mat78, Maz75, NZ75, SSM70, SS71, VH74, Wil76]. brownons [Wei73]. bubble [SKW71]. Burgers [HY75, Kol76].

calculated [Ale76]. calculation [ABO74, Fli74, LP74, Mah75, RS75, SSL76]. calculations [Blo70, BS79, Tho73]. call [Bru78]. canonical [Nen73, RP78]. capillarity [Row79, vdW79]. Carlo [BS72a, Bor74, Fli74, MKB73, SLM^{+}76]. Cascade [KS78a]. case [Bax77, Ver76]. catenary [RO73]. causality [Sch77]. Cell [Vis78, HMP74]. certain [WD75]. chain [BB78, CJSJT73, Cha72, GC73, Hyn74, HKW75, LMpY73, Nor73, vdBCWW78, vdBCGW79]. chains [HD74, SV78, Vil78]. chance [Man77, MD79]. Chang [TR73]. change [Ale77]. changes [Lan73, Lan74]. chaos [YY79]. Chaotic [TK79, YY79]. Chapman [DB71, Kih75]. characteristic [Ell77, Ell78]. characterization [ST74, Tan74]. characterizing [MD74]. charge [FRR73]. charged [Blu78a, HS78]. chemical [Aon75, CGMW77, Gar76, GMWM76, GC77, GF75, Gil77, JD79, Lov72, MWG75, MW74, MNM78, PK76, PM78, Sta71, Wid78]. chemically [DO79]. circuits [Lan75]. class [BTZ79, Fei78, RO73, Ray79, RW77]. Classical [Tav76, BS71, BN73, BG75, Coh70, De 75, DLP79, ED75, GP76a, GPE79, HH76, Kug73, ML79, ML72, ML74, Mye72b, Mye72a, Nis79, Per76a, RP78, Rei70e, Ric74, Ros79, RFB72, SM76, See72, SZ76, Wil76]. Closure [AM79, Ado71, BH78, FSV74, HB77]. Cluster [LP77a, BK72, HKM78, HKK79, KBS76, KS78b, Mer79, Pet79]. clustering [MM79]. Clusters [KLPS78, Del79, HBQ78, KM75, Kik71, PLM^{+}78, Rei70e]. coarse [HH76]. coarse-graining [HH76]. code [Coh75, HR71a]. coefficient [Har70, Hil74, RFB72, SZ76, Sut71]. coefficients [BFC74, FB73, GCFB73, Rai78, RS75, YK76]. Coexistence [Leb77a]. coherent [TS76b, Yos70]. coil [De 75]. Collective [Bau78, FBB73, Kug73]. Collision [IK79, ABO74, vWdB78]. collisional [RS75]. collisions [Maz75]. combinatorial [CH79]. Commensurability [SWS79]. Comment [GH73, Shi73, Bat74, Gar76, NR74]. Comments [Chu72, TM72, TR73, MNM78]. Comparison [Syk73, Bat74, HC73, MD73, WLS78]. complete [MR75a, MR76]. component [BW77, Ber79b, GPE79, LMpY73, MK75, SM76]. Computer [Ale72, MTH73, Yug77]. Computer-oriented [Ale72]. Concerning [Gil77]. condensate [Roe78]. condensation [GIB74, Sut71]. condition [Hyn74, MMSG77, VH74]. conditions [BSO71, BLP79, Ken78, MR75a, MR76, Mer79, MMS77]. conductances [Sni76]. conduction [Ali79, Dav78, FB78]. conductivity [Tun76a]. conjectured [Bax76]. Connection [McL78, HH76]. conservation [TS70]. conservative [McL78]. conserving [Vis78]. consistent [AZ73b, Zan74]. Consolidated [MC70]. constant [ML72, PW79, SJ70, Wei73]. constant-magnetization [ML72]. constants [GF75]. constrained [LT79]. constraints [Mou74]. construction [Var79]. Continuity [RP78]. continuous [AZ73a, CO77, FRR73, Jon78, KMS73, Row79, Shl74, Syl76, Tun74, Tun76b, vdW79, MS73]. continuous-spin [Syl76]. continuous-time [KMS73, Shl74, Tun74, Tun76b, MS73]. contracted [Ber79a]. contraction [Pet79]. contribution [Ler71, Nen73]. contributions [Ano72g, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. controlled [LK79]. convergent [BF73, BFC74, FB73]. convex [KM75]. Conway [SS78]. cooling [Ale75a]. cooperative [SK76]. Core [YK76, CWPS79, Huc77, RH72, RL76]. Corner [Bax76, Bax77, Tsa77, BE79]. correction [Kug75]. corrections [Sti76, Wil76]. Correlated [SV78]. Correlation [GT75, MMS77, MP79, Szà78, ALM78, AA78, AS72, AS73, Blu78a, CG78, Dun77, EHvL76a, EHvL76b, Jan77, Niw78, NM78, RS75, See72, Tur74]. Correlations [GMWM76, BBO75, BSO70, BSO71, BO75, PC77, PT77, Sas75, SD72, SO74]. Coulombic [NM78]. coupled [Ber79b, HR71b]. Coupling [NH78, Aon75, EHvL76b, GC79, JD79]. criteria [Pfa77]. Critical [Bar74, BP78, DO79, Fis78, NM78, Chu72, DLP79, Fox79, Kat70, Kaw75, KBS76, MR79, MA77, YOK77]. Critique [Gro73a]. crystal [vH78]. crystalline [FL79]. crystals [HL79, Iva73b]. cubic [SLM^{+}76]. Curie [EN78]. current [Tun76c]. curvature [MM71, Sut71]. curved [MR75b, VR73a, VR73b]. cycle [TK79]. cycles [GC73]. cylinders [Leu70].

D [Row79]. damping [Kug73]. Davies [Ali79]. death [Mat77]. Decay [BSO70, BSO71, SD72, SO74]. Decimation [Jon78]. decomposition [Bor74]. decrease [Ell77, Ell78]. definition [Ale77, Gru76, Pre75]. degenerate [Bax73a]. delocalized [Ken78]. Delta [Rai78]. Delta-function [Rai78]. demagnetization [Ale75b]. dense [BF73, BFC74, Dav73, FB73, Jan77, JD79, PC77]. density [BFC74, Bra74a, CK74, DMM77, FB73, Gro79, Mry74, Per76b, RL78, RW79, TS75, vdW79, Row79]. dependence [Sut71]. dependent [CK74, DS78, Fel72, Lev73, NPW70, TM76, Tur74]. Derivation [Spo77, GC79, Kan77, NZ75, RP78, TD71, Wil76, Yas77, dBvW78, vWdB78]. deriving [GCdR77, MMT78, dRGC78]. description [AZ73a, MMN75, Ste71]. descriptions [Ber79a, Tav76]. detailed [Ale77, GLS79]. detectabillity [Yos70]. Determination [Vis79]. developed [KS78a]. diagram [Gud73]. diagrammatic [WLS78]. diamagnetism [Bru70]. diameters [ALM78]. diatomic [GC73]. Dielectric [SG73b]. different [ALM78, Leb77b]. differential [AM79, Bud76, Bud77]. diffusely [SSM70]. Diffusion [CCR79, LK79, Tun76a, Aon75, GP76b, Kih75, Lut76, MD73, MD74, SS71, Tun74, Wei73, vK77]. Diffusion-controlled [LK79]. diffusivity [Sch75b]. dilute [BO75, Bla75, Cha70, FG76, MR75a, MR76, Maz75, Yen72]. dimension [Man77, RS77, MD79]. dimensional [ALM78, Bau78, BSO70, Blo70, Bra74b, CKS71, Fer73, Fer74, FF75a, FF75b, Fis78, Fli74, Fri79, FL79, GP76b, Gla75, GIB74, HS71, Hil77, HY75, Lak70, Lev73, LP74, Lin74b, LP77b, MS78, Mer79, Niw78, Per77, Pol72, SWS79, SM76, Sch74, Tro77b, Tro77a, Wie75, Yug77]. Dimensionality [DO78, GT75]. dimensions [AdFHK79, AZ74, ECD^{+}78, HKW78, Roe78]. dipolar [IF79]. dipoles [Blu78b]. Dirac [Fer74, Ada75, Fer73]. Direct [Lee72]. Direct-interaction [Lee72]. directed [Tan74]. Discrete [BV79, Vis79, MW75, RD71, Wei72]. Discretized [GVZ79]. discrimination [Ale75a, Ale75b]. disorder [Fis78]. disordered [Egg74, Kow77, RW77]. dispersion [ED75]. displaced [ALM78, AS72]. displacement [AA78]. displacements [Iva73b]. dissimilar [BGCP73]. dissipation [Ber79a, GKS72, Leu70]. dissipative [Kos75, McL78]. distortion [vdBCWW78]. distributed [Lin74b]. distribution [BSO70, BS76, BF73, BFGC73, Del79, FRR73, HS78, HKM78, HKK79, KS78b, Pol72, RD74, RP78, Ric74, TM76]. distributions [AA78, Lev73, PM78, Tom74]. Distributive [CH79]. disturbances [Bud77]. divergence [Tan74]. double [Mat78]. double-well-potential [Mat78]. draining [CJSJT73]. dressed [Gro74]. driven [MC70]. droplet [SKW71]. droplets [BS72a]. droplike [Kik71, Rei70e]. Duality [Run76]. Duffing [BLS76, Ued79]. Dynamic [MKB73, Sch75a, Bau78]. Dynamical [GP76a, Har70, RW77, SK76, SS78, Vst76]. Dynamics [BB78, BM78, BM79, CJM79, CH79, Coo79, Fri79, GLS79, Gra78, HS71, Kno79, KBS76, Lüc76, Mar77, NH78, Tat70, YOK77].

ed [May78]. edge [Abr78a, KF78]. editors [Ano73k, Ano73l, Ano74b, Ano71a, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l]. effect [Coo79, GZW75, HD74, Kei76, MM71, MT73, PC77, Sni76]. effective [Yug77]. effective-medium [Yug77]. Effects [FBB73, AA78, GP71, Hyn74]. Eigentheory [HM77]. eigenvalues [Bax73a]. eight [Bax73a, Bax76, Bax77]. eight-vertex [Bax73a, Bax76, Bax77]. Einstein [GIB74]. electric [IF79, PW79, Sch74]. electrical [Lan75]. electrolytes [Sni76]. electromagnetic [SJ70]. electromagnetic-wave [SJ70]. electron [Bau78, Gro77, Kug75, PW79, Sch74]. electrons [Spo77]. elementary [Len78]. Elliott [ML77]. energy [Car74, GPE79, KS79, Kow77, MA76, SG73b, vdBCWW78, vdBCGW79]. Ensemble [Hos76, Ale72, ML72]. ensembles [Kei70a, Kei70b, ML79]. Enskog [BGCP73, CP75, DB71, Kih75, PC76, Rés78]. Entropie [Hir75]. entropies [Lib74]. Entropy [Ale75a, Ale75b, Ale76, Lan73, Lan74, Rou73, Ale77, Ali79, Duf77, FS71, GH73, Lan75, MW77, Run73, Ska75, TM72]. equation [ABO74, Akc77, BBO75, BMP71, BGV74, BS76, BH78, CGZ79, CLLS73, CWPS79, DO79, Fer73, Fri77, GVZ79, Grm71, Gro73b, Gro73c, HM77, HSS77, HMMH77, HB77, KBR72, Kei72, Kol76, Ler71, MMN75, MM78, MPY76, MW75, Mry74, Onu78a, Piñ74, Pin75, Rés78, Sch75b, STH77, Spo77, Spo79, TS70, TS75, Tun75, Ued79, Var79, Wei72, Yas77, dBvW78, Fer74]. Equations [CL75, Ado71, AM79, ASJS71, AS72, Aon72, Bud76, Bud77, Cha70, CSW72, DB71, DO78, Fox77, Fox79, FT79, FSV74, GCdR77, GC79, Gar76, GC77, Kan77, KMS73, Ken78, MMT78, Nis76, Ray79, STH77, Syk73, TD71, Vis79, WD75, Wil76, Zwa73, dBvW78, dRGC78, vWdB78]. Equilibrium [Gir78, Per76a, Sch73, CM74, EHvL76b, HR71b, Kei76, MC74, RL75, Rés75, RL77, Spo74, Var79, YOK77]. Equivalence [Fer73, Fer74, BLP79, Leb77b, ML79]. Ergodic [Tro77b, Tro77a, GC73]. Erratum [MPPT80]. error [AM79]. Essential [KS78b]. Estimates [DLP79, Lav78]. estimating [MC70]. Euler [AdFHK79]. Evaluation [Rid77]. evidence [Sch74]. evolution [HTGT78, LLL77, Pfa77]. Exact [HS78, Hil77, HKW78, HR76, Kim79, LP74, RL78, Blu78a, DD72a, FF73, ML74, NZ75, Pli74, vdBCWW78, GCdR77, HMP74]. Exactly [BIG73, BTZ79, Ver76, Wie75]. example [Coo79, Gla75]. examples [Kim79]. excess [Run73]. exchange [Hil74]. excitable [AK76]. excited [Bud77]. exciton [HKM78]. excluded [HD74]. Existence [DB71, KS79, Spo79, Suz79, Red75]. expansion [HMMH77, Pli74, Rai78, Sla79, Tom74, WD75, WLS78]. expansions [Bax76, BE79, DO78, Gud73, MC70, Pet79, vWdB78]. experiment [CJM79, Sch77]. Explicit [Pol72]. exponential [GKPR78]. exponentially [Lin74b]. exponents [BP78, Chu72]. expressions [Duf77]. extension [BBO75]. extensivity [Suz76]. external [Per76a, Per77]. extremum [LSFL75].

factor [Bau78]. far [Kei76, YOK77]. Fermi [Ada75, FRR73]. ferromagnet [Abr78b, Bar77, MMS77]. ferromagnetic [Abr78a, Szà78]. Ferromagnetism [Pli74]. ferromagnets [Fis78, Leb77a, Leb77b, Mer79, MS74, New76, Syl76]. Field [AZ73a, Ano74b, Fer73, YOK77, BB78, DO78, FBB73, Gat76, IF79, Kug75, MT73, Per76a, Per77, PW79, Sch75b, Ver76, Fer74]. Field-theoretic [YOK77]. Fields [Ano71a, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Bra74b, RS73, Rou73, Sch74]. finite [CSW72, LS72, Mye72b, Mye72a, SWS79, Tro77a]. finite-dimensional [Tro77a]. finite-strength [CSW72]. First [LSFL75, Bra74a, FB73, KLPS78, MWG75, PLM^{+}78]. first-order [KLPS78, MWG75, PLM^{+}78]. Fisher [SKW71]. five [FT79]. five-mode [FT79]. flow [AdFHK79, ECD^{+}78, HY75, Hos76]. Fluctuating [HML73, VH74]. Fluctuation [KMK73, Suz76, AK76, Ber79a, BNW78, DO79, GKS72, Leu70, MA76, PK76, SNZ74, Suz77a]. fluctuation-dissipation [Ber79a, GKS72, Leu70]. Fluctuations [Akc77, FtHW70, Nic72, TS76a, DO78, Gra78, Kei76, Lan73, Lan74, Leu70, Lov72, MMN75, MM78, Mry74, NR74, Onu78b, MC74]. fluid [Bau78, Chu72, CWPS79, FF75a, FF75b, JD79, Kaw75, MK75, Man76, NN74, NM78, PC77, Per76a, Per76b, SLD76, TD71]. fluids [Aon72, BN73, BG75, Dav73, EGGM75, ED75, Eva79, GP76a, Ler71, Pol72, RL75, Rés75, RL77, SP77, VR73a]. flux [AK76]. Fokker [CLLS73, GLS79, HM77, Sch75b, WD75]. folding [Nor73]. force [AK76, Sch75b]. force-flux [AK76]. forced [HY75, Hos76]. forces [Bra74a, NPW70, YK76]. Form [MD79, GPE79, Man77, RO73]. Formal [Gro76, RL75]. formalism [HH76, HR71a, Kre75, MPS79, RH71]. formalisms [MD73]. formation [BK72]. formula [Ada75, Gud73, Len78, SKW71]. formulas [Blu78a, HS78]. formulation [Gat76, KBR72, Pat72, SG73b, TM76]. Foundations [Spo74]. Fourier [LS78]. Fractals [Man77, MD79]. frameworks [HN79]. Francis [Bru70]. free [Bar74, CJSJT73, Fer73, Fer74, KS79, Kow77, MA76, Nen73].

Frequency [Lov72, Ale76]. Friction [Kos75]. fully [KS78a]. function [AS73, BG75, BS76, BF73, CJSJT73, Dun79, FF73, Gro74, Gro77, HK75, Jan77, Kug73, LS72, Mah75, NM78, Oon72, Rai78, Rés75, SG73a, See72, SLS79, SLD76, TM76, Tur74]. Functional [Kol76, Mah75, RS73, Cho75]. functionals [Blo73]. functions [ALM78, AS72, BSO70, Blu78a, BFGC73, Cha70, CG78, Dun77, Ell77, Ell78, EHvL76a, EHvL76b, GT75, HS78, Ler71, MMS77, Nis79, Niw78, Pol72, RO73, RS75]. Further [VR73b]. Future [Ano72g, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l].

Gage [Fri73]. game [SS78]. games [Sch73]. gas [Bis73, Bla75, CP75, FG74a, FG74b, FG76, Gir78, Gro73a, Hal75, HH79, Hub76, Kug75, Lak70, LS78, Maz75, Mon73, PW79, RL76, See72, SS71, Wie73]. gas-liquid [Bis73, Gro73a]. gases [BO75, BF73, BFC74, Cha70, FBB73, FB73, GVZ79, GIB74, HS71, Huc77, Kih75, KM75, Nen73, RS75, RH72, Sti76, YK76]. gauge [Pat72]. Gauss [Dut72]. Gaussian [Bri77, Bud77, Dun79, Fox77, Rou73]. General [ABO74, CGWE73, Lav78, New76, MPS79, TM76]. Generalization [Yen72, vH78]. generalizations [Tan74]. Generalized [DO78, KMS73, Ken78, PC76, TS76b, BB78, Fox77, KBR72, NZ75, NW74, Rou73, STH77, Zwa73]. Generating [Blo73, SLS79]. generators [IK79]. Gennes [Kle74]. Geometry [KM75]. Gibbs [Len78, Lib74, ML79, Mou74, RP78, Hir75]. Gibbsian [Yen72]. Ginzburg [Bra74b]. GKS [MS74]. Glauber [BB78]. Global [CGZ79, San72]. governed [Ued79]. graining [HH76]. grand [Nen73]. grand-canonical [Nen73]. graph [BS72b]. graphs [Lav78]. Green [Ler71, SG73a]. ground [FtHW70, GR79, HR79, Sla79, Tho73, vdBCGW79]. ground-state [vdBCGW79]. Group [HD74, BS79, GP79, KHY76a, KHY76b]. Growth [Kaw75, PLM^{+}78].

half [Abr78a]. half-plane [Abr78a]. Hall [Har70]. Hamiltonians [DS78, KS79]. hard [ALM78, Blu78a, Blu78b, FA75, HP73, HMP74, HS78, Hil74, Hub76, Huc77, IF79, Per76a, PM78, RS77, RL75, Rés75, RL77, RH72, RL76, Syk73, VR73a, vBE79]. hard-core [Huc77, RH72, RL76]. hard-sphere [Hub76, RL75, Rés75, RL77, VR73a]. harmonic [Bud77, HR71b, Hyn74, HKW75, RW77, Tho73, vH78]. having [Huc77].

Heat [FB78, Ali79, Bru78, Dav78]. Heisenberg [ML72, ML74, SG73a]. held [ML77]. helical [Rei71]. helix [De 75]. helix-coil [De 75]. Henderson [Blu78a]. Hénon [Sim79]. Hénon-Pomeau [Sim79]. Hermitian [ASJS71]. Hestenes [Fri73]. heteropolymers [De 75, SV78, Vil78]. hexagons [HP73]. hierarchy [Ado71, AM79, MPY76, MMT78, RS77, TS75, Var79]. High [Hil74, RL78, FA75, GT75, Pli74, Sch74]. High-density [RL78]. High-temperature [Hil74, Pli74]. Hobson [TR73]. homogeneous [CP75, Pet79, PC76]. Hubbard [Pli74]. hydrodynamic [MM78, Rés70, TD71]. hydrodynamics [BV79, HML73, VH74, Vis78, Vis79]. hydrogen [AH72, AH75]. hydrogen-bonded [AH72, AH75]. hyperscaling [KvLH77]. hypothesis [FSV74, Row79, vdW79].

ideal [GZW75, Hal75]. II [AH75, Bax77, BLS76, CG78, EHvL76a, FG74b, FB73, GCFB73, Gro73b, HMP74, HKM78, HR71b, Kei70b, Lav78, Lin74b, MC74, MLS78, Mye72a, Pet79, Rés75, SLS79, SV78, VH74, dBvW78, dRGC78]. III [BSO70, BN73, BFC74, Bra74a, Bud77, EHvL76b, Gro73c, HKK79]. Illustration [Sch77]. Im [Ale77]. Important [Ano72h, Ano72i, Ano73y, Ano76o, Ano77n, Ano78n, Ano79m]. impurities [Spo77]. impurity [HKW75]. inaccuracy [Tan74, Tan76]. include [BBO75]. Incompatibility [Pat72]. independent [LSFL75]. individual [Ale76]. inducing [Son79]. inductive [Hob72, Shi73]. inelastic [EGGM75, Maz75]. Inequalities [Syl76, Dun77, Dun79, LP77a, MS74, MP79, Szà78]. inequality [Bri77, Kow77, San72, Sch74]. inference [MW77, Sýk74]. infinite [ALM78, FT79, Fri79, GR79, LLL77, MPPT78, MPPT80, Niw78, Tro77b, vH78]. infinite-dimensional [Tro77b]. infinite-particle [Fri79]. infinite-volume [GR79]. Influence [Mry74, MS73]. Information [Pfa77, Bat74, HC73, Kre75, Lin74a, RH71, ST74, Ska75, Tan76, TR73, Vis78]. information-conserving [Vis78]. Information-theoretic [Pfa77, ST74]. informational [BS73]. inhomogeneous [HM77, RD74]. Initial [Hyn74, Ken78, Ric74]. instabilities [NMMVK76, PM78]. instability [NN74, Suz77b]. integral [ABO74, Cho75, KBR72, KM76, Mah75, Pat72, Rei70e]. integrals [Ada75, Gro79, Hil74, Kol76, LP74, PD78, Rid77, Wie79]. integration [RS73]. interacting [Bra74a, CM74, MC74, Rid77]. interaction [Lee72, Lee73, LMpY73, vdBCGW79]. interactions [BTZ79, CSW72, DLS78, HKM78, KvLH77, LP77a, ML74, Run73, Tur74]. Interest [Ano73h, Ano75h, Ano71a, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73i, Ano73j, Ano73k, Ano73l, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74b, Ano74l, Ano74m, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75i, Ano75j, Ano75k, Ano75l]. interesting [BS72b]. interface [KPR77]. interfaces [Wid78]. interfacial [Bax73a]. intermediate [Bar77]. intermolecular [YK76]. internal [GPE79]. interparticle [SP77]. interpretation [Hob72, Shi73, Swe73]. Introduction [CFTA79]. invariance [Pat72]. Invariant [MM79]. Invariants [DMM77]. Inversion [RFB72, SZ76]. involving [ASJS71]. ionized [Lav78]. ions [Blu78b]. irreducible [Gud73, LP74]. irregularities [MW75]. Irreversible [Ale75b, CM74, MC74, Ale75a, Ale77, Coo79, MPS79, Pet79, Pfa77]. Ising [Fer74, Abr78a, Abr78b, BB78, Bax77, BE79, BSO70, Blo73, Blo70, Del79, Egg74, Fer73, Fis78, Fli74, Kim79, Leb77a, Leb77b, LMpY73, Mar77, MS78, MA77, Mer79, MMS77, MS74, New76, Per77, SK76, Syl76, Tsa77, Tsa79, Ver76]. isotherms [SP77]. isotopic [Kih75]. isotropic [DLS78]. IV [BSO71, BGV74, MS73, WLS78].

J. [Row79]. Jaynes [FS71, GH73, TM72]. jellium [LN75, LN76]. joint [Tan74]. Jones [GR79]. Journal [Ano72g, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l]. justification [Kei70a, Kei70b].

Kac [ML74, Tav76]. kernel [MD74]. Kihara [Kih75]. kind [Bru78]. Kinetic [AS72, Cha70, ECD^{+}78, Grm71, MT73, vBE79, CCSD71, CPC79, DB71, DD72a, Dav73, GCdR77, Gro76, Kim79, MMT78, Nis76, Piñ74, RL75, Rés75, RL77, Ros79, Sas75, SZ74, SSM70, Syk73, dBvW78, dRGC78, vWdB78, EHvL76a]. Kinetics [KLPS78, Gil77, HR76, Mat74, Mat75, PC77]. Kirkwood [Fri77, NN74]. Kirkwood-modified [Fri77]. Knefler [Wid79]. Kullback [Bat74, HC73, TR73].

L [Hir75]. Landau [Bra74b, Bru70, MM78]. Langevin [Akc77, Fox77, Gar76, KBR72, Onu78a, Yas77, Zwa73]. large [BP78, Coo79]. large-order [BP78]. laser [Cho75, HN79, Lov72, Wal76]. Lattice [AH72, AH75, GVZ79, HL79, Bax78, BE79, BSO70, BIG73, CGZ79, DLP79, FG74a, FG74b, Hal76, HP73, HH79, Huc77, Jon78, LP77a, MPPT78, MPPT80, ML79, MW75, NPV79, Neu75, NH78, Pol72, RL78, RW79, RH72, RL76, Run76, SSL76, Sla79, Spo79, Sta78, Sti76, SLM^{+}76, Tsa79, Tun76a]. lattices [Fel72, MA77, MS73, RW78, SWS79, SLS79]. Lauritzen [Nor73]. law [LS78]. layers [HKM78, MR75b, VR73a]. leading [McL76]. learning [GP71]. Lebowitz [Swe73]. Lee [RH72, KF78]. length [Lut76]. lengths [DLP79, HBQ78]. Lennard [GR79]. Lennard-Jones [GR79]. levels [Car74]. Liboff [Hir75]. Life [SS78]. Lifschitz [RW79]. Lifshitz [MM78]. Light [MK75, FBB73]. limit [CJSJT73, DL79, GVZ79, LN75, LN76, PT77, SM76, Suz79, TK79, Yos70, vdBCWW78]. limitations [Coh75]. limiting [Mah75]. line [CCSD71]. Linear [BMP71, KBS76, NPV79, AS79, DS78, DB71, Fel72, GC73, Hil77]. linearization [Bud76, WLS78]. linearized [ABO74, Pin75, Rés70]. Liouville [BGV74, Gro73b, Gro73c, STH77]. liquid [BS72a, Bis73, BS73, FL79, GF75, Gro73a, HL79, KPR77, MK75]. liquid-crystalline [FL79]. liquid-vapor [KPR77]. liquids [CL75, DB71, HMP74, Kug73, Nis79]. local [DO78, Kug75, MMN75, EHvL76b]. locality [Pat72]. locomotion [BNW78]. Long [AA78, BW77, BO75, PW79, Sas75, SLD76, CJSJT73, KS79, KvLH77, Lak70, ML74, Lav78]. long-chain [CJSJT73]. Long-range [BO75, Sas75, KS79, KvLH77, ML74, Lav78]. Long-time [AA78, BW77, PW79, SLD76]. Lorentz [LS78, PW79]. Lorenz [Kno79, Lüc76, YY79]. loss [RO73, SG73b]. Lotka [Kei76, Red75, vK74]. Lovett [MMSG77]. Low [Sla79, Ler71, Bax76]. Low-temperature [Sla79, Bax76]. lower [Gro79].

M [Son79]. macro [TS76a]. macro-spin [TS76a]. macroprocesses [Mat74, Mat75]. Macroscopic [SZ74, Wal76, Coo79, NPW70, See72, Vst76]. macrovariables [Akc77, KMK73, Suz76]. magnetic [BB78, BIG73, FBB73, Mon73, Ver76]. magnetization [ML72]. magnetoelastic [BIG73]. many [CKS71, Ros79, Sla79]. many-dimensional [CKS71]. many-particle [Ros79]. Markov [Ber79a, HTGT78, HD74, IK79, LSFL75, Mou74]. Markovian [STH77]. Master [TS75, BMP71, DO78, DO79, GC77, HSS77, HMMH77, Kei72, KMS73, Ken78, MMN75, STH77, Wei72]. Mathematical [GP79]. matrices [Bax76, Bax77, BE79, Car74, DMM77, Gro79, Pie72, Tsa77]. matrix [Bax73a, RW77]. matter [Lee73]. maximum [Bax73a, FS71, GH73, TM72]. Maxwell [ML73, PL71]. Maxwellian [GVZ79]. Mayer [Lav78, Rai78]. mean [AM79, Blu78a, Blu78b, DO78, Gat76, GPE79, HS78]. mean-field [Gat76]. means [Rei70e, vdBCWW78]. measure [Spo74, TR73]. measure-theoretic [Spo74]. Measurements [Lin74a]. measures [AdFHK79, HC73, ST74, Tan76, Bat74]. mechanical [BTZ79, BF73, BFC74, CM79, FB73, FA75, Fur74, Pre75, RD71, Sni76]. Mechanics [Ano76p, Ano77o, Ano78p, Ano79n, Leb76, Ano77p, Ano78o, Ano78q, Bax78, Bis72, BFGC73, Bra74a, Bra74b, CK74, DZ78, DD72b, DF72, Dut72, GCFB73, GCdR77, Gla75, Hos76, Kei70a, Kei70b, Kos75, LT79, LP77b, May78, ML77, MKB73, Mye72b, Mye72a, Ogu72, RP78, Ric74, SS78, dRGC78, vK78]. media [Cho75]. medium [FB78, Yug77]. Meeting [Ano76p, Ano77o, Ano78p, Ano79n, Leb76, Ano77p, Ano78o, Ano78q]. melting [SV78, Vil78]. membranes [AK76]. Memoriam [Wid79]. Memory [Gro74, AS72, BG75, MD74, Nis79]. memoryless [STH77]. metabolizing [MTH73]. Metastability [MS78, CO77, KLPS78]. Metastable [ML73, YY79, MLS78, PL71]. method [AM79, Bud76, Cho75, Hil74, Kik71, MM78, Mat78, MA77, MMT78, MKB73, TS76b]. methods [ASJS71, AS79, Fox77, ML77, WLS78, vK78]. metric [Fei79]. Microscopic [CP75, BW77, Coo79, GC79, Kle74, Mat74, Mat75, MA76, TS75]. Mixture [BN73, ALM78, BGCP73, BH78, BFGC73, GCFB73, Piñ74]. Mixture-model [BN73]. mixtures [MR75a, MR76]. Mobility [Sch75b]. modal [Lee72]. modal-interaction [Lee72]. mode [EHvL76b, FT79, GC79, JD79, Mat77, Tom74]. mode-coupling [EHvL76b]. mode-mode [GC79, JD79]. Model [Per76b, Ale75a, Ale75b, Ali79, AZ74, Bar74, Bar77, Bax73a, Bax73b, Bax76, Bax77, BE79, BN73, Blo73, Blo70, BS73, BM78, BM79, CO77, Cha72, CGWE73, CFTA79, CPC79, Dav78, Del79, DZ78, Dun79, Egg74, Fer73, Fer74, FF73, FF75a, FF75b, FL79, GP76a, HK75, Har70, HMP74, HKW78, HKM78, Iva73a, Kei76, KS78a, Kim79, Kno79, Kol76, LS72, Lüc76, MS78, ML72, ML74, MTH73, MW75, Neu75, Nor73, NH78, Per77, Pli74, RD74, Red75, Run76, SK76, Tav76, Tsa77, Tsa79, Tun74, Tun76c, Ver76, Wie75, Wu78, YY79, vK74, vK77]. Models [DF72, AH72, AH75, Bax78, BTZ79, EN78, HL79, HH79, KS79, KF78, Lov78, Mar77, MWG75, MA77, MP79, Sch75a]. modes [Bau78, Kug73, Rés70]. Modified [BG75, Fri77, Piñ74, Rés78]. modulation [Coh75, HR71a, Lov72]. Molecular [HS71, TD71, CJM79, EGGM75]. molecule [KM75]. molecules [CJSJT73, Huc77, Rei71]. moment [MMSG77]. Moments [Tun76b, Mah75]. momentum [HKW75, TS70]. monomode [Wal76]. Monotone [Ell77, Ell78]. Monte [BS72a, Bor74, Fli74, MKB73, SLM^{+}76]. Montroll [Tun75, Tun76b]. mössbauer [GZW75]. motion [Ber79b, BZ71, Bru78, CJSJT73, CL75, De 75, GZW75, HML73, Iva73a, Mat78, Maz75, NZ75, VH74, Vis79, Wil76]. Motroni [Fri73]. moving [Spo77]. multi [Onu78a]. multi-Poissonian [Onu78a]. multicomponent [Aon72, Bri77, RL76]. multiphase [RD74]. Multiple [RW78, MLS78]. multiplicative [Fox77].

narrowing [Kle74]. nature [Lee73]. Navier [FT79, Hos76]. near [Kaw75, KBS76, MR79, RL75, Rés75, RL77, SKW71, Suz77b]. nearest [LMpY73, vdBCGW79]. nearest-neighbor [LMpY73]. Necessary [BSO71]. neighbor [LMpY73, Lin74b, vdBCGW79]. network [Oon72]. neural [Oon72]. neutron [EGGM75]. News [Ano70a, Ano70b, Ano70c, Ano70d, Ano71b, Ano71c, Ano72j, Ano72k, Ano72l, Ano72m, Fut71]. next [vdBCGW79]. next-nearest-neighbor [vdBCGW79]. noise [Bud77, But72, Onu78a, Sch74, Tun76c]. noise-temperature [Sch74]. non [ASJS71, Fox77, GVZ79, HTGT78, Szà78, STH77]. non-Gaussian [Fox77]. non-Hermitian [ASJS71]. non-Markov [HTGT78]. non-Maxwellian [GVZ79]. non-purely-ferromagnetic [Szà78]. nonadditive [Bra74a].

Nonanalytic [ED75]. nonatomic [Sch73]. Nonclassical [Fox79, Wie73]. nonelectrolytes [BN73, Sni76]. Nonequilibrium [Bra74a, Duf77, HR76, MW74, Mye72b, Mye72a, BF73, BFGC73, BFC74, DMM77, DF72, FB73, GCFB73, GCdR77, Lov78, MM78, MWG75, MPY76, MPS79, MNM78, Nic72, NMMVK76, NR74, Pet79, Ros79, Suz76, Suz79, dRGC78]. Nonexistence [NN74]. nonisotropic [DLS78]. Nonlinear [DL79, GKS72, Gra78, HKW75, Zwa73, Akc77, BZ71, Bud76, BLS76, Bud77, CKS71, DO78, DZ78, ECD^{+}78, Fei78, Fei79, Fur74, HMMH77, KS75, KBS76, Mat77, Mat78, MLS78, McL78, MC70, Nic72, Ray79, Rés78, SZ74, SNZ74, WLS78, AK76]. Nonlinearity [SK76]. nonlocal [BW77, dBvW78]. nonnearest [Lin74b]. nonnearest-neighbor [Lin74b]. nonpolynomial [BTZ79]. nonrelativistic [Pat72]. nonstationary [Fox77]. nonsymmetric [Eva79]. nonuniform [SS71]. Note [HTGT78, Hir75, Ada75, PM78, Wei79, Yas77, VH74]. notions [Ska75]. November [ML77]. Nucleation [MR79, BS71, BK72, Coh70, HK75, KLPS78, Kat70, Kik71, PC77, Rei70e]. nuclei [Kaw75]. nucleotide [Kre75]. Number [BNW78, SSL76]. numbers [Rid77]. Numerical [BV79].

oard [Ano73k, Ano73l]. obeying [BMP71]. Oboukhov [Son79]. observables [See72]. observation [Kei76, Sch75a]. occupation [Wei79]. occurring [Lav78]. One [FL79, Lev73, Per77, ALM78, BW77, BSO70, BG75, Ber79b, Fer73, Fer74, FF75a, FF75b, GP76b, Gla75, GPE79, GIB74, Hil77, HY75, LP74, Lin74b, Niw78, Pol72, RS77, SWS79, SM76, Sch74]. one- [GIB74]. one-body [BG75]. one-component [BW77, Ber79b, GPE79, SM76]. One-dimensional [FL79, Lev73, Per77, ALM78, BSO70, Fer73, Fer74, FF75a, FF75b, GP76b, Gla75, Hil77, HY75, LP74, Lin74b, Niw78, Pol72, SWS79, Sch74]. onset [NMMVK76]. onto [IF79]. Open [DS78, BM78, BM79, Nic72]. operator [HH76, Mat78, MD73, WLS78]. operators [ASJS71, Gro73c, IK79, MM79]. optic [MT73]. optic-acoustic [MT73]. optical [Rou73, Yos70]. Optimal [BS79]. order [BFC74, Bra74a, BP78, FG74b, FB73, KLPS78, Leb77b, MWG75, PLM^{+}78, SG73a, Zan74]. ordered [WR76]. organizing [KS75]. orientations [Huc77]. oriented [Ale72]. Origin [MM78]. Ornstein [BS76, BH78, CWPS79, HB77]. Orthogonal [CLLS73]. Oscar [Wid79]. oscillating [CKS71]. oscillator [BZ71, BLS76, HR71b, Mat78, MC70, Sch76, Tho73]. oscillators [Mat78, McL78]. oscillatory [DO79]. other [Duf77, Syk73]. overarousal [GP71].

Pair [Jan77, Tur74, BBO75, BO75, Blu78a, TM76]. Paper [Hir75, Bat74, GH73, Gro73a, TM72]. parallel [HMP74]. paramagnet [SG73a]. parameter [Zan74]. parameters [Leb77b]. partially [LK79]. particle [AF74, Ber79b, BS76, CM79, De 75, Fer73, Fer74, FA75, Fri79, Hyn74, Lee73, MR75a, MR76, Maz75, Ros79, SG73b, VR73a, VR73b]. particles [Ale76, DD72b, Gro74, Nis76, SSM70]. Partition [Gro77, Dun77, Dun79, HK75, Mah75, Oon72]. passage [LSFL75]. Path [Gro79, Hil74, KM76, Pat72, PD78, Wie79]. Pearson [NW74]. Peierls [Egg74, vdBCWW78]. Percolation [HKM78, HKK79, Wu78, HBQ78, KS78b, LP77a, RL78, Sta78, SLM^{+}76, WP78, Yug77]. Percus [AF74, Var79]. periodic [AdFHK79, Man76, MW75, Pie72, RW78, SLS79]. perturbation [BP78, GP76a, HR79, Lee72, MPPT78, MPPT80, WD75]. Phase [DLS78, HP73, HH79, Huc77, AS73, AZ73a, AZ73b, BIG73, Bis73, Cho75, Fli74, FL79, Grm71, Gro73a, Gru76, Hal76, Hil77, KLPS78, ML79, MWG75, MW74, MNM78, PLM^{+}78, Rei70e, Rid77, TS76b, Zan74]. phase-integral [Cho75]. phases [Leb77a]. phenomena [Kos75, Lov78, SLM^{+}76, Suz77b, Ued79]. Phenomenological [Aon72, SZ74]. phenomenon [Bla75]. photon [Lee73]. photon’-matter [Lee73]. physics [Car74, Rés70, PD78, Ano72g, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i]. Physics [Ano79j, Ano79k, Ano79l]. picture [Lee73, SKW71]. placed [Tun76a]. Planck [CLLS73, GLS79, HM77, Sch75b, WD75]. plane [Abr78a, Dun79]. plasma [BW77, Ber79b, GPE79, Jan77, SM76, SG73b]. plus [SJ70]. Poincaré [GC73]. point [Kaw75, Suz77b]. points [Fox79, KBS76, Lav78, Sch73]. Poisson [CG78, GC77]. Poissonian [Onu78a]. polarization [Bax73b]. polyatomic [Eva79, FBB73, KM75, YK76]. polymer [Cha72, Nor73]. polymers [HD74]. polynomial [CLLS73]. Pomeau [Sim79]. pools [Kre75]. population [FtHW70]. Position [KvLH77, GP79]. Position-space [KvLH77, GP79]. positive [FRR73]. Possible [SKW71]. potential [CCR79, De 75, EHvL76a, GR79, Gro77, Mat78, MA76, Wid78, YK76, vK77]. potentially [GZW75]. potentials [HS71]. Potts [HKW78, Wu78]. predicted [Kat70]. prescription [FS71, GH73, TM72]. presence [IF79]. pressure [Eva79, Nen73, Pre75]. principle [Dut72, Rei72]. principles [Gro73b, Gro73c]. Probabilistic [AS79, Mat74, Mat75, Tho73, WP78, RD74]. probabilities [GKPR78, Hob72, Shi73]. probability [CK74, FS71, GH73, TM72]. probe [GZW75]. problem [FRR73, FG74a, FG74b, HKK79, MW75, SM76, Sch74, Yug77]. problems [AS79, CM79, KS78b, RH71, Sýk74]. procedure [Mah75, vdBCWW78]. procedures [BS79]. proceedings [ML77]. process [Ale75b, Ale77, BNW78]. processes [Ale75a, Ber79a, Bud76, BLS76, Bud77, CH79, Fox77, HTGT78, IK79, KM76, KS75, LK79, LSFL75, Mat77, MLS78, MPS79, Pet79, Pie72, SZ74, WLS78]. production [Ali79, Lan75, Nis76]. Program [Ano76p, Ano77o, Ano77p, Ano78o, Ano78p, Ano78q, Ano79n, Leb76]. projection [Gro73c, HH76, Mat78, MD73, WLS78]. proof [Len78]. propagation [Cho75]. properties [Bax76, BS73, Fei79, FBB73, GP79, Hal76, LS78, LSFL75, Lov78, Man76, Mon73, MKB73, Rei71, SLS79, Tro77b, Tro77a, Tun75, Bra74a]. property [Mer79]. proposal [TR73]. Prospectus [Rei70a, Rei70b, Rei70c, Rei70d]. protein [Rei71]. pulse [Coh75, HR71a]. purely [Szà78]. pycnonuclear [Jan77].

quadratures [MW75]. Quantal [HSS77]. Quantitative [Fei78]. quantities [Duf77, Lin74a]. quantization [GLS79]. Quantum [FA75, MPY76, Sti76, Sýk74, Bis72, CM79, DS78, DF72, DLP79, DLS78, Kei70a, Kei70b, Kos75, Len78, MPS79, NPV79, Nen73, Ogu72, PD78, Pet79, RD71, Spo74, Ste71, Tav76, Wal76, dBvW78, vWdB78]. quasi [Sta71]. quasi-chemical [Sta71]. quasicrystal [Iva73a]. quasideterministic [Tro77a]. quasihydrodynamic [Rés75]. Quasinormal [Tom74]. quasiparticle [CFTA79]. quenched [Fis78]. Questions [BG71, Sha72, She70, Swe74, Tyk71, Whe72].

radial [Pol72]. radiation [MT73, Yos70]. Random [GKPR78, KBR72, Lin74b, MS73, SLS79, BNW78, Car74, Cho75, FB78, Gro77, Hos76, KMS73, KS79, Mou74, NW74, RW78, RW79, SJ70, Shl74, SSL76, Spo77, Tun74, Tun76b, Tun76c, WR76, Wei76, Wei79, WP78, SLS79, Vil78]. random-forced [Hos76]. Randomly [Ued79, HY75, MC70, Tun76a]. range [BO75, HKM78, KS79, KvLH77, Lav78, ML74, Sas75]. rapid [BSO71]. rate [CSW72, GF75, Kaw75, MLS78]. Rayleigh [vH78]. reacting [MK75]. reaction [Aon75, CGMW77, JD79, Lov72]. reactions [FG76, GMWM76, GF75, Jan77, MWG75, MW74]. reactive [NH78]. real [Neu75]. realistic [HS71]. reconsideration [BS71]. rectangular [Abr78b]. reducing [MM78]. Reduction [MW75]. reflecting [SSM70]. reformulating [Ogu72]. Reiss [Kik71]. related [Gud73]. relation [AF74, AK76, Leu70, Run76, Yen72]. relations [ED75, Pol72, SZ74]. Relaxation [Mat77, HKW75, KBS76, KMK73, MT73, NPV79, Sch74, SO74, Suz76, Suz77a, BSO70, MLS78]. Remark [vK74]. remarks [Sch77]. renormalization [BS79, GP79, Hil77, KHY76a, KHY76b, KvLH77, SNZ74, vdBCWW78]. renormalization-group [BS79, GP79]. Renormalized [Ros79, TM76]. Replies [Fri73]. representation [CG78, GC77, KM76, TS76b]. representations [Mah75, MD74]. repulsions [SP77]. resonant [McL78]. Response [Bru78, May78, BMP71, DS78, MC70, NPV79, TK79]. restricted [HMP74]. results [BV79, DD72a, GCdR77, GKPR78, HKW78, New76, dRGC78]. reversed [Sch77]. reversibility [Mat74, Mat75]. review [Bru78, Gri70, May78, MD79, Sat70, Sch72, Wie79, vK78]. reviews [FTM70, TD78]. revisited [BLS76, SM76]. Rice [Wid79]. Richard [Hir75]. rigid [BGCP73, DD72b, Piñ74]. Rigorous [New76, PL71, CO77, Gat76]. ring [Tav76]. Rochester [ML77]. rods [ALM78, Lak70, Per76a]. root [Lav78]. Rotational [Iva73b, CL75, FL79, Iva73a]. rotator [Dun79]. rotators [Bri77]. rough [MM71, SLD76]. Ruelle [RH72]. rule [ML79].

scale [HSS77]. Scaled [MR75a, MR76, AF74, FA75, VR73a, VR73b]. Scaling [Hal75, Lov78, Sta78, Suz77b, MMT78]. Scattering [Lut76, MW75, EGGM75, FBB73, Kug73, MK75, MM71, SJ70]. Schrödinger [MW75, Yas77]. screening [Bra74b]. Second [SG73a, BFC74, MMSG77, RFB72, SZ76, FG74b]. second-moment [MMSG77]. Second-order [SG73a, FG74b]. Self [AZ73b, Zan74, KS75, MD73, MD74]. Self-consistent [AZ73b, Zan74]. self-diffusion [MD73, MD74]. self-organizing [KS75]. Semiclassical [HN79]. semiinfinite [GP76b]. separation [Fli74]. Sequences [FT79]. serial [GP71]. Series [BE79]. set [MR75a, MR76]. Shannon [Bat74, Hir75, HC73, Lib74]. shapes [CCSD71]. shear [BW77]. Short [Syk73]. Short-time [Syk73]. signals [Yos70]. Simple [CPC79, Kow77, Har70, Kim79, Mee73, Nis79, NH78, SSL76, SLM^{+}76]. simple-source [Mee73]. simulation [MTH73, Vis79]. simulations [Bor74]. Single [Ber79b, Pli74]. single-band [Pli74]. Single-particle [Ber79b]. singular [Gla75, WD75]. singularity [KS78b, KF78, RW79]. sinks [LK79]. site [HKK79, Yug77]. site-bond [HKK79]. sites [Tun76a]. size [Blu78b, KS78b]. slip [VH74]. Smith [Blu78a]. Smoluchowski [BNW78, Wil76]. smooth [IF79]. soft [BV79, CWPS79]. soft-core [CWPS79]. soft-sphere [BV79]. solid [PD78]. soluble [BIG73, BTZ79, DO79, Ver76, vK77]. Solution [BS76, Blu78b, BH78, CWPS79, FRR73, HB77, MR75a, MR76, AM79, Cha72, GF75, Har70, HMP74, ML74, Nor73, Pin75, RL78, RH71, WD75]. solutions [BN73, BGV74, Bud76, Bud77, CGZ79, CLLS73, DB71, Gro73b, Gro73c, HR76, Kei72, Pol72, Ray79, Sch75b, Shl74, Spo79, Sta71, Tun74, Tun75, Wei72]. solvable [GP76a, Wie75]. solvents [AH72, AH75]. solving [ASJS71]. Some [BS73, DD72a, MNM78, Sch74, Tun75, Wei72]. Sommerfeld [Ada75]. sound [Mee73]. source [Mee73]. space [GP79, KvLH77, Onu78b, TS76b]. spaces [MW75]. spacing [Pol72]. Spanning [HBQ78]. spans [WR76]. spatial [CGMW77, GT75, PC77]. spatially [RD74]. Special [Sch75b]. species [Rid77]. spectra [IK79, RD71]. Spectral [MD74, CCSD71]. spectrum [FBB73, RW77]. sphere [BV79, Hub76, RL75, Rés75, RL77, VR73a]. spheres [BGCP73, Blu78a, FA75, HS78, Hil74, IF79, Piñ74, RS77, SLD76, Syk73, vBE79]. spherical [Bar74, Bar77, Blu78a, Blu78b, FRR73, GPE79, HS78, KF78, PT77]. Spiking [GP71]. spin [BSO70, DLS78, GT75, MS74, MP79, SD72, Syl76, TS76a, TS76b]. Spinodal [Bor74]. spins [Abr78a, Jon78]. Spontaneous [Bax73b]. spring [Cha72]. Square [Tsa79, Bax78, BE79]. squares [HMP74]. Stability [Lan75, PK76, Pfa77]. staggered [Bax73b]. stars [Jan77]. starting [Tun76b]. state [BS73, Fox79, FtHW70, Fur74, GR79, Lan73, Lan74, Ler71, MLS78, Mry74, PD78, PC77, Per76a, Red75, TS76b, Tho73, Wei73, Wei76, vdBCGW79]. states [Ale76, HR79, Kei72, MLS78, ML73, Nic72, NR74, PL71, RW79, Sla79]. Stationary [AdFHK79, AS79, LK79, MLS78, PM78, Gro73b]. Statistical [Ano72g, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76p, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77o, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78p, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i]. Statistical [Ano79j, Ano79k, Ano79l, Ano79n, Bra74b, Car74, DZ78, Dut72, Fur74, Leb76, Lüc76, LP77b, MK75, MW77, May78, RD71, RH71, SS78, Sch76, Sni76, Tat70, AK76, Ano77p, Ano78o, Ano78q, Bax78, Bis72, BTZ79, BF73, BFGC73, BFC74, Bra74a, Bud76, CK74, DD72b, DF72, FB73, GCFB73, GCdR77, Gla75, HR71a, Kei70a, Kei70b, Kno79, Kol76, Kre75, LT79, Mee73, MPY76, ML77, MKB73, Mye72b, Mye72a, PD78, RP78, Rés70, Ric74, Ste71, VR73b, WLS78, dRGC78, vK78, ML77, vK78]. Statistically [Pie72]. Statistics [SJ70, EN78, Fel72, Lak70, Spo74, BNW78]. steady [Fur74, Kei72, Lan73, Lan74, NR74, PC77, Red75]. steady-state [Lan73, Lan74, PC77]. steps [Lin74b, SSL76]. Stillinger [MMSG77]. Stochastic [CGMW77, FF75a, FF75b, Fri79, GLS79, GF75, MWG75, MLS78, Pin75, AM79, Ale75a, Ale75b, Ale77, Bud76, BLS76, Bud77, CSW72, DZ78, FG76, Fox77, GMWM76, Gil77, HN79, KM76, KS75, Mar77, McL76, MA77, MTH73, Ogu72, Red75, STH77, Tro77b, Tro77a, WLS78]. Stochasticity [CKS71]. Stokes [FT79, Hos76]. strange [Son79]. Strategies [SNZ74]. strength [CSW72]. strong [SP77]. strongly [Ber79b]. Structure [KPR77, Wid78, Bau78, Gro76, MR75a, MR76]. structures [CGMW77, Sch75b]. Studies [Bud76, BLS76, Bud77, WLS78, SLM^{+}76]. study [BS72a, CO77, KS75, MD73, Yug77]. subspaces [MM79]. substance [Neu75]. Successive [McL76]. sufficient [BSO71]. sums [Hal76]. superconducting [Leu70]. superfluidity [CS76]. Superposition [Rei72]. superradiance [Gir78]. superradiant [TS76a]. supersaturations [Kat70]. surface [Bar74, BS72a, MR75a, MR76, Nen73, Per76b, Run73, Sut71]. surfaces [HK75, MM71, VR73b]. susceptibilities [Kim79]. Susceptibility [Abr78b]. sustained [YY79]. symmetrization [SKW71]. symmetry [CPC79, HR79]. Symposium [ML77]. system [CGWE73, Coo79, FF73, Fli74, FSV74, GP76b, Gla75, Hil77, LS72, Niw78, SS78, Son79, TS76a, TS75, Tur74, Ued79, Zan74]. systematic [CCR79, NZ75]. Systems [FtHW70, Akc77, BMP71, Bra74a, CKS71, Chu72, CM74, CSW72, DS78, DO79, DLP79, DLS78, Fri79, GT75, HR71b, Jon78, KvLH77, Kow77, LLL77, Lav78, LP77a, Len78, Lin74b, LT79, MC74, MPPT78, ML79, MM78, MTH73, MPY76, MPS79, MNM78, Mou74, Mye72b, Mye72a, NPV79, Nic72, NMMVK76, Nis76, PC77, Pet79, PC76, PK76, RD71, RD74, Rid77, RW77, RW79, Ros79, SK76, SD72, Sla79, Spo74, Ste71, Suz76, Suz77a, Suz79, Szà78, TS76b, Tro77b, Tro77a, Yen72, MPPT80]. systems-dynamical [SK76]. systems-I [Mye72b]. systems-spin [TS76b].

tailing [RW79]. Tait [Fri77]. technique [GC77, Lee72]. techniques [SLS79, Mah75]. temperature [Bax76, Fis78, FA75, Hil74, MR79, NPW70, Pli74, RP78, SWS79, Sch74, Sla79]. temperature-dependent [NPW70]. temperatures [DLP79, Ler71]. temporal [CGMW77]. tension [Bax73a]. tensor [Eva79]. terms [EHvL76a]. test [SG73b]. their [DS78, IK79, PD78, Tan74]. theorem [Ber79a, GKS72, Hub76, Rés78, Swe73, vH78]. theoretic [Pfa77, ST74, Spo74, YOK77]. theories [CGWE73, GMWM76, Syk73]. Theory [EGGM75, Kug75, MPS79, Nis76, Pet79, AF74, AK76, AS79, Ale72, AS73, AZ73a, BGCP73, BS71, BMP71, BN73, BK72, BF73, BFC74, BP78, But72, Car74, CCSD71, CP75, Coh70, DB71, DD72a, Dav73, DO78, EHvL76b, ECD^{+}78, Fer73, Fer74, FG76, FB73, FS71, FA75, Fur74, GH73, Gat76, GF75, Gro76, HN79, HD74, JD79, Kat70, Kik71, Kol76, Lav78, Lin74b, MK75, MR75a, MR76, MLS78, Maz75, MT73, Mee73, ML73, ML77, NPW70, NZ75, Ogu72, PL71, PC76, RD71, Rei70e, RH71, RL75, Rés75, RL77, Ros79, Row79, RH72, Sas75, SG73a, SZ74, SV78, SSM70, Sni76, Sta71, Sta78, SNZ74, Suz77b, Sýk74, TM76, TM72, TR73, Tun76c, VR73a, VR73b, Vil78, Vis78, Vst76, WR76, Wil76, Yug77, vBE79, vK78, vdW79]. thermal [Kih75, Yos70]. Thermodynamic [Aon75, Duf77, HN79, HR71a, Kre75, Ler71, Leu70, LN75, LN76, Lin74a, Pre75, RH71, Row79, San72, SM76, Suz79, vdW79]. Thermodynamical [Len78]. Thermodynamics [MR75b, BK72, Pfa77, Sch76, VR73a, VR73b]. Thermoelastic [Rei71]. thermohydrodynamic [Kan77]. thin [Lak70]. third [Hil74, YK76]. Thomas [FRR73]. Thorne [BGCP73, Piñ74]. Thought [CJM79]. Thought-experiment [CJM79]. Three [Yug77, ECD^{+}78, KM75, Lak70]. Three-dimensional [Yug77, Lak70]. three-molecule [KM75]. threshold [GP71]. Time [ALM78, CK74, Fel72, LLL77, Niw78, Oon72, See72, AA78, AS72, BW77, CG78, DS78, EHvL76a, EHvL76b, HTGT78, HSS77, HR71a, KMS73, Lev73, LSFL75, MLS78, MS73, PW79, RS75, Shl74, SLD76, Syk73, TM76, Tun74, Tun76b].

Time-correlation [See72]. Time-dependent [CK74, Fel72, DS78, Lev73, TM76]. Time-displaced [ALM78, AS72]. times [Wei79]. Toeplitz [Pie72]. tracer [RD74]. transfer [Bax73a, Bax76, Bax77, BE79, Tsa77]. transformation [RD71]. transformations [Fei78, Fei79, GP79, Jon78, KHY76a, KHY76b]. transient [Cha72, Suz77b]. transition [AS73, AZ73b, BIG73, Bis73, Gru76, GKPR78, Hal76, HP73, Hil77, PLM^{+}78, YY79, Zan74, Gro73a]. transitional [Ued79]. transitions [AS73, AZ73a, De 75, DLS78, FL79, Grm71, HH79, Huc77, KLPS78, Lin74b, MWG75, MW74, MNM78]. Translation [Row79]. translational [NH78]. Transmission [HR71a, Kre75]. Transport [LS78, ABO74, AS79, Bla75, CM79, CGWE73, DL79, GCFB73, Gra78, MK75, Mon73, RD74, Sch74, Sch75b, Spo77, SNZ74, YK76, BFC74, FB73, RS75]. trapping [RW78, SSL76]. Treatment [Kik71, Rei70e, AZ73b, CCR79, PL71, Zan74]. triangular [HP73, Run76, Tsa77, Ver76]. Tribus [Fri73]. tribute [ML77]. tricritical [Fox79]. truncation [BS79, FT79]. Tunneling [CM79, Lee73]. Turbulence [HY75, FT79, KS78a, Kol76, Mee73]. Two [LMpY73, AdFHK79, AZ74, Bau78, BGV74, Blo70, BF73, DB71, ECD^{+}78, Fer73, Fer74, Fis78, Fli74, Fri79, GIB74, HS71, HKW78, Huc77, Kim79, Lav78, LP77b, MK75, MS78, Mer79, Wei73, Wei76, Wie75, CG78]. two-body [BGV74, BF73]. Two-component [LMpY73, MK75]. two-dimensional [Bau78, Blo70, Fer73, Fer74, Fis78, Fli74, Fri79, GIB74, HS71, LP77b, MS78, Mer79, Wie75]. two-state [Wei73, Wei76]. Two-time [CG78].

Uehling [dBvW78]. Uhlenbeck [dBvW78]. unified [GCdR77, dRGC78]. uniform [NN74, PW79, Sch75b]. uniformity [Wid78]. Uniqueness [RS77, Spo79]. universal [Fei79]. universality [Chu72, DO79, Fei78]. University [ML77]. unrestricted [WR76]. unstable [Suz77a]. upper [vdBCGW79]. use [ASJS71, MA76]. using [Fox77, Kik71].

valid [HSS77]. validity [CSW72, Gil77, San72, Sch74]. vapor [KPR77]. variable [HKM78]. variable-range [HKM78]. variables [Bis72, KS75]. variation [MM71, Per76b, Rei72, Row79, vdW79]. Variational [Bax78, KHY76a, KHY76b, ASJS71, CS76, Gro73b, Tsa79, Gro73c]. vector [AZ74, MP79, PT77, SJ70, Zan74]. Velocity [CJSJT73, MPPT78, FF73, LS72, Rés75, SLD76, Tur74, MPPT80]. velocity-dependent [Tur74]. versus [STH77]. vertex [Bax73a, Bax76, Bax77, Run76]. via [Dun77, Hil74, SLS79]. vibrational [MT73, SO74]. VIII [Maz75]. virial [AF74, Hil74, Rai78, RFB72, SZ76, YK76]. viscosity [BW77, BV79]. viscous [ECD^{+}78]. Volterra [Kei76, Red75, Rid77, vK74]. volume [GR79, HD74]. vortices [LP77b]. vs [Hir75, Lib74].

W [ML77]. Waal [Row79]. Waals [ML73, PL71, NPW70, Wie73]. Walk [GKPR78, NW74, SLS79, Tun74, Tun76b, Tun76c, Wei76]. walkers [RW78]. walks [BS72b, KMS73, Lin74b, MS73, SLS79, Shl74, SSL76, WR76, Wei79]. wall [BS76, IF79]. wall-particle [BS76]. water [FG74a, FG74b, HH79]. wave [SJ70]. waves [MW75]. Weak [Hal76, KvLH77]. Weiss [EN78, Tun75, Tun76b]. well [Mat78]. Wilson [AS73]. without [AS72, Mah75]. WKB [CCR79].

Yang [KF78, RH72]. Yevick [AF74, Var79]. York [ML77]. Yukawa [BH78, CWPS79, HB77].

Zernike [BS76, BH78, CWPS79, HB77]. zero [Bra74b, MA76]. zero-dimensional [Bra74b]. Zeros [Dun77, Dun79]. Zwanzig [HH76, Nor73].

References

Alder:1978:LTC

[AA78] B. J. Alder and W. E. Alley. Long-time correlation effects on displacement distributions. Journal of Statistical Physics, 19 (4):341–347, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011753.

Aisbett:1974:GCC

[ABO74] Janet Aisbett, John M. Blatt, and Alex H. Opie. General calculation of the collision integral for the linearized Boltzmann transport equation. Journal of Statistical Physics, 11(6):441–456, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008888.

Abraham:1978:BSE

[Abr78a] D. B. Abraham. Block spins in the edge of an Ising ferromagnetic half-plane. Journal of Statistical Physics, 19(5):553–556, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011699.

Abraham:1978:SRI

[Abr78b] D. B. Abraham. Susceptibility of the rectangular Ising ferromagnet. Journal of Statistical Physics, 19(4):349–358, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011754.

Adawi:1975:NSF

[Ada75] I. Adawi. A note on Sommerfeld’s formula for Fermi–Dirac integrals. Journal of Statistical Physics, 12(3):263–265, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011970.

Albeverio:1979:SMP

[AdFHK79] S. Albeverio, M. Ribeiro de Faria, and R. Høegh-Krohn. Stationary measures for the periodic Euler flow in two dimensions. Journal of Statistical Physics, 20(6):585–595, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009512.

Adomian:1971:CAH

[Ado71] G. Adomian. The closure approximation in the hierarchy equations. Journal of Statistical Physics, 3(2):127–133, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019846.

Ahn:1974:PYV

[AF74] Saeyoung Ahn and H. L. Frisch. Percus–Yevick virial relation and scaled particle theory. Journal of Statistical Physics, 11 (4):283–289, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009789.

Abraham:1972:LMH

[AH72] Douglas B. Abraham and Ole J. Heilmann. Lattice models for hydrogen-bonded solvents. Journal of Statistical Physics, 4 (1):15–41, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008469.

Abraham:1975:LMH

[AH75] D. B. Abraham and O. J. Heilmann. Lattice models of hydrogenbonded solvents. II. Journal of Statistical Physics, 13(6):461– 472, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013145.

Aizawa:1976:STE

[AK76] Yoji Aizawa and Yonosuke Kobatake. A statistical theory of excitable membranes. Nonlinear force-flux relation and fluctuation. Journal of Statistical Physics, 15(2):129–139, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012032.

Akcasu:1977:FMN

[Akc77] Ziya A. Akcasu. Fluctuations of macrovariables in nonlinear systems: Langevin equation approach. Journal of Statistical Physics, 16(1):33–58, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014604.

Alexandrowicz:1972:COA

[Ale72] Z. Alexandrowicz. Computer-oriented approach to the ensemble theory. Journal of Statistical Physics, 5(1–2):19–34, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008369.

Alexandrowicz:1975:EIC

[Ale75a] Z. Alexandrowicz. Entropy of irreversible cooling and the ‘discrimination’ of model stochastic processes. Journal of Statistical Physics, 12(4):271–289, April 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012065.

Alexandrowicz:1975:IAD

[Ale75b] Z. Alexandrowicz. Irreversible adiabatic demagnetization: Entropy and discrimination of a model stochastic process. Journal of Statistical Physics, 13(3):231–244, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012840.

Alexandrowicz:1976:ECF

[Ale76] Z. Alexandrowicz. Entropy calculated from the frequency of states of individual particles. Journal of Statistical Physics, 14(1):1–9, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020131.

Alexandrowicz:1977:DIE

[Ale77] Z. Alexandrowicz. A definition of irreversible entropy change by detailed (im)balance of a stochastic process. Journal of Statistical Physics, 16(2):139–148, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01418748.

Alicki:1979:EPD

[Ali79] Robert Alicki. On the entropy production for the Davies model of heat conduction. Journal of Statistical Physics, 20(6):671–677, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009517.

Aizenman:1978:TDC

[ALM78] Michael Aizenman, Joel Lebowitz, and Joaquin Marro. Timedisplaced correlation functions in an infinite one-dimensional mixture of hard rods with different diameters. Journal of Statistical Physics, 18(2):179–190, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014309.

Adomian:1979:CAE

[AM79] G. Adomian and K. Malakian. Closure approximation error in the mean solution of stochastic differential equations by the hierarchy method. Journal of Statistical Physics, 21(2):181–189, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008697.

Anonymous:1970:Na

[Ano70a] Anonymous. News. Journal of Statistical Physics, 2(1):107–111, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009714.

Anonymous:1970:Nb

[Ano70b] Anonymous. News. Journal of Statistical Physics, 2(2):203–204, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009741.

Anonymous:1970:Nc

[Ano70c] Anonymous. News. Journal of Statistical Physics, 2(3):291, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030748.

Anonymous:1970:Nd

[Ano70d] Anonymous. News. Journal of Statistical Physics, 2(4):389–390, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020446.

Anonymous:1971:FIB

[Ano71a] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 3(4):423–424, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008282.

Anonymous:1971:Na

[Ano71b] Anonymous. News. Journal of Statistical Physics, 3(1):93, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012191.

Anonymous:1971:Nb

[Ano71c] Anonymous. News. Journal of Statistical Physics, 3(2):243, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019853.

Anonymous:1972:FIBa

[Ano72a] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 4(1):81–82, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008474.

Anonymous:1972:FIBb

[Ano72b] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 4(2–3):229–230, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008553.

Anonymous:1972:FIBc

[Ano72c] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 5(1–2):143–144, May 1972. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01008379;   http://link.springer.com/article/
10.1007/BF01008379.

Anonymous:1972:FIBd

[Ano72d] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 5(3):237–238, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01023745;   http://link.springer.com/article/
10.1007/BF01023745.

Anonymous:1972:FIBe

[Ano72e] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 6(1):63–64, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01060203.

Anonymous:1972:FIBf

[Ano72f] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 6(2–3):227–228, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023691.

Anonymous:1972:FCB

[Ano72g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 6(2–3):229, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01023692.

Anonymous:1972:IAa

[Ano72h] Anonymous. Important announcement. Journal of Statistical Physics, 4(2–3):i, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008544.

Anonymous:1972:IAb

[Ano72i] Anonymous. Important announcement. Journal of Statistical Physics, 6(2–3):65, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023678.

Anonymous:1972:Na

[Ano72j] Anonymous. News. Journal of Statistical Physics, 4(1):79, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008473.

Anonymous:1972:Nb

[Ano72k] Anonymous. News. Journal of Statistical Physics, 5(1–2): 141, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01008378;   http://link.
springer.com/article/10.1007/BF01008378.

Anonymous:1972:Nc

[Ano72l] Anonymous. News. Journal of Statistical Physics, 5(3): 235, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01023744;   http://link.
springer.com/article/10.1007/BF01023744.

Anonymous:1972:Nd

[Ano72m] Anonymous. News. Journal of Statistical Physics, 6(1): 61, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060202.

Anonymous:1973:FIBa

[Ano73a] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 7(1):91–92, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01406134.

Anonymous:1973:FIBb

[Ano73b] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 7(4):373–374, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014911.

Anonymous:1973:FIBc

[Ano73c] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 8(1):103–104, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008444.

Anonymous:1973:FIBd

[Ano73d] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 8(2):213–214, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008542.

Anonymous:1973:FIBe

[Ano73e] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 8(3):295–296, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012574.

Anonymous:1973:FIBf

[Ano73f] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 8(4):373–374, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020276.

Anonymous:1973:FIBg

[Ano73g] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 9(1):97–98, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016798.

Anonymous:1973:FIBh

[Ano73h] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 9(2):193–194, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016848.

Anonymous:1973:FIBi

[Ano73i] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 9(3):271–272, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008733.

Anonymous:1973:FIBj

[Ano73j] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 9(4):373–374, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012168.

Anonymous:1973:FIOa

[Ano73k] Anonymous. Fields of interest of oard of editors. Journal of Statistical Physics, 7(2):185–186, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01024214.

Anonymous:1973:FIOb

[Ano73l] Anonymous. Fields of interest of oard of editors. Journal of Statistical Physics, 7(3):283–284, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01030308.

Anonymous:1973:FCBa

[Ano73m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 7(1):93, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01406135.

Anonymous:1973:FCBb

[Ano73n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 7(2):187, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024215.

Anonymous:1973:FCBc

[Ano73o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 7(3):285, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030309.

Anonymous:1973:FCBd

[Ano73p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 7(4):375–376, April 1973.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014912.

Anonymous:1973:FCBe

[Ano73q] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 8(1):105, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008445.

Anonymous:1973:FCBf

[Ano73r] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 8(2):215, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008543.

Anonymous:1973:FCBg

[Ano73s] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 8(3):297–298, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012575.

Anonymous:1973:FCBh

[Ano73t] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 8(4):375, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020277.

Anonymous:1973:FCBi

[Ano73u] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 9(1):99, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016799.

Anonymous:1973:FCBj

[Ano73v] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 9(2):195, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016849.

Anonymous:1973:FCBk

[Ano73w] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 9(3):273, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008734.

Anonymous:1973:FCBl

[Ano73x] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 9(4):375–376, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012169.

Anonymous:1973:IA

[Ano73y] Anonymous. Important announcement. Journal of Statistical Physics, 7(1):1, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01406129.

Anonymous:1974:A

[Ano74a] Anonymous. Announcement. Journal of Statistical Physics, 11 (3):277, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010221.

Anonymous:1974:FIBj

[Ano74b] Anonymous. Field of interest of bord of editors. Journal of Statistical Physics, 11(4):359–360, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009794.

Anonymous:1974:FIBa

[Ano74c] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 10(1):107–108, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011721.

Anonymous:1974:FIBb

[Ano74d] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 10(2):179–180, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009720.

Anonymous:1974:FIBc

[Ano74e] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 10(3):255–256, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016180.

Anonymous:1974:FIBd

[Ano74f] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 10(4):347–348, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012256.

Anonymous:1974:FIBe

[Ano74g] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 10(5):435–436, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008804.

Anonymous:1974:FIBf

[Ano74h] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 10(6):505–506, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020404.

Anonymous:1974:FIBg

[Ano74i] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 11(1):83–84, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019479.

Anonymous:1974:FIBh

[Ano74j] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 11(2):177–178, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012519.

Anonymous:1974:FIBi

[Ano74k] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 11(3):279–280, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010222.

Anonymous:1974:FIBk

[Ano74l] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 11(5):437–438, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026735.

Anonymous:1974:FIBl

[Ano74m] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 11(6):527–528, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008895.

Anonymous:1974:FCBa

[Ano74n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(1):109, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011722.

Anonymous:1974:FCBb

[Ano74o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(2):181, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009721.

Anonymous:1974:FCBc

[Ano74p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(3):257, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016181.

Anonymous:1974:FCBd

[Ano74q] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(4):349, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012257.

Anonymous:1974:FCBe

[Ano74r] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(5):437, May 1974.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008805.

Anonymous:1974:FCBf

[Ano74s] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 10(6):507, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020405.

Anonymous:1974:FCBg

[Ano74t] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(1):85, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019480.

Anonymous:1974:FCBh

[Ano74u] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(2):179, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012520.

Anonymous:1974:FCBi

[Ano74v] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(3):281, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01010223.

Anonymous:1974:FCBj

[Ano74w] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(4):361, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009795.

Anonymous:1974:FCBk

[Ano74x] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(5):439, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01026736.

Anonymous:1974:FCBl

[Ano74y] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 11(6):529, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008896.

Anonymous:1975:FIBa

[Ano75a] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(1):89–90, January 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01024184.

Anonymous:1975:FIBb

[Ano75b] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(2):163–164, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010031.

Anonymous:1975:FIBc

[Ano75c] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(3):267–268, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011971.

Anonymous:1975:FIBd

[Ano75d] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(4):361–362, April 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012068.

Anonymous:1975:FIBe

[Ano75e] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(5):445–446, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012888.

Anonymous:1975:FIBf

[Ano75f] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 12(6):517–518, June 1975. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025836.

Anonymous:1975:FIBg

[Ano75g] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 13(1):103–104, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012603.

Anonymous:1975:FIBh

[Ano75h] Anonymous. Fields of interest of board of Editors. Journal of Statistical Physics, 13(2):185–186, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01221378.

Anonymous:1975:FIBi

[Ano75i] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 13(3):279–280, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012844.

Anonymous:1975:FIBj

[Ano75j] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 13(4):375–376, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012015.

Anonymous:1975:FIBk

[Ano75k] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 13(5):457–458, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013120.

Anonymous:1975:FIBl

[Ano75l] Anonymous. Fields of interest of Board of Editors. Journal of Statistical Physics, 13(6):531–532, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013149.

Anonymous:1975:FCBa

[Ano75m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(1):91, January 1975.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024185.

Anonymous:1975:FCBb

[Ano75n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(2):165, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01010032.

Anonymous:1975:FCBc

[Ano75o] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(3):269, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011972.

Anonymous:1975:FCBd

[Ano75p] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(4):363, April 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012069.

Anonymous:1975:FCBe

[Ano75q] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(5):447, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012889.

Anonymous:1975:FCBf

[Ano75r] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 12(6):519, June 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025837.

Anonymous:1975:FCBg

[Ano75s] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(1):105, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012604.

Anonymous:1975:FCBh

[Ano75t] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(2):187, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01221379.

Anonymous:1975:FCBi

[Ano75u] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(3):281, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012845.

Anonymous:1975:FCBj

[Ano75v] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(4):377, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012016.

Anonymous:1975:FCBk

[Ano75w] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(5):459, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01013121.

Anonymous:1975:FCBl

[Ano75x] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 13(6):533, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01013150.

Anonymous:1976:Aa

[Ano76a] Anonymous. Announcement. Journal of Statistical Physics, 14(4):395, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030202.

Anonymous:1976:Ac

[Ano76b] Anonymous. Announcement. Journal of Statistical Physics, 15 (5):438, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020346.

Anonymous:1976:Ab

[Ano76c] Anonymous. Announcements. Journal of Statistical Physics, 15(2):180, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012038.

Anonymous:1976:FCBa

[Ano76d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 14(2):205, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011766.

Anonymous:1976:FCBb

[Ano76e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 14(3):305, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023631.

Anonymous:1976:FCBc

[Ano76f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 14(4):397, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030203.

Anonymous:1976:FCBd

[Ano76g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 14(5):467, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01040706.

Anonymous:1976:FCBe

[Ano76h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 14(6):541, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012854.

Anonymous:1976:FCBf

[Ano76i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(1):85–86, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012811.

Anonymous:1976:FCBg

[Ano76j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(2):179, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012037.

Anonymous:1976:FCBh

[Ano76k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(3):265, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012882.

Anonymous:1976:FCBi

[Ano76l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(4):343, October 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023058.

Anonymous:1976:FCBj

[Ano76m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(5):437, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01020345.

Anonymous:1976:FCBk

[Ano76n] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 15(6):519, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01020806.

Anonymous:1976:IA

[Ano76o] Anonymous. Important announcement. Journal of Statistical Physics, 15(6):i, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020797.

Anonymous:1976:PSM

[Ano76p] Anonymous. Program of the 34th Statistical Mechanics Meeting. Journal of Statistical Physics, 14(1):87–89, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020137.

Anonymous:1977:A

[Ano77a] Anonymous. Announcement. Journal of Statistical Physics, 16(1):120, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014611.

Anonymous:1977:FCBa

[Ano77b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(1):119, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014610.

Anonymous:1977:FCBb

[Ano77c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(2):233, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01418754.

Anonymous:1977:FCBc

[Ano77d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(3):319, March 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020386.

Anonymous:1977:FCBd

[Ano77e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(4):397, April 1977.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020431.

Anonymous:1977:FCBe

[Ano77f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(5):451, May 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013186.

Anonymous:1977:FCBf

[Ano77g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 16(6):505, June 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01152286.

Anonymous:1977:FCBg

[Ano77h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(1):45, July 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01089376.

Anonymous:1977:FCBh

[Ano77i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(2):95, August 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01268921.

Anonymous:1977:FCBi

[Ano77j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(3):153, September 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01106914.

Anonymous:1977:FCBj

[Ano77k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(4):263, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01040106.

Anonymous:1977:FCBk

[Ano77l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(5):383, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014406.

Anonymous:1977:FCBl

[Ano77m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 17(6):529, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014353.

Anonymous:1977:IA

[Ano77n] Anonymous. Important announcement. Journal of Statistical Physics, 16(1):1, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014601.

Anonymous:1977:PSMa

[Ano77o] Anonymous. Program of the 36th Statistical Mechanics Meeting. Journal of Statistical Physics, 16(1):121–127, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014612.

Anonymous:1977:PSMb

[Ano77p] Anonymous. Program of the 37th statistical mechanics meeting. Journal of Statistical Physics, 16(6):507–510, June 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01152287.

Anonymous:1978:A

[Ano78a] Anonymous. Announcement. Journal of Statistical Physics, 19(1):108, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020337.

Anonymous:1978:FCBa

[Ano78b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(1):109, January 1978.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014674.

Anonymous:1978:FCBb

[Ano78c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(2):235, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014313.

Anonymous:1978:FCBc

[Ano78d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(3):333, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018098.

Anonymous:1978:FCBd

[Ano78e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(4):421, April 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01106732.

Anonymous:1978:FCBe

[Ano78f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(5):533, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014523.

Anonymous:1978:FCBf

[Ano78g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 18(6):637–638, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014483.

Anonymous:1978:FCBg

[Ano78h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(1):107, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020336.

Anonymous:1978:FCBh

[Ano78i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(2):203–204, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012513.

Anonymous:1978:FCBi

[Ano78j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(3):315, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011728.

Anonymous:1978:FCBj

[Ano78k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(4):427, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011760.

Anonymous:1978:FCBk

[Ano78l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(5):559–560, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011701.

Anonymous:1978:FCBl

[Ano78m] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 19(6):655, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011775.

Anonymous:1978:IA

[Ano78n] Anonymous. Important announcement. Journal of Statistical Physics, 19(6):561, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011767.

Anonymous:1978:PSMa

[Ano78o] Anonymous. Program of the 38th statistical mechanics meeting. Journal of Statistical Physics, 18(1):111–116, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014675.

Anonymous:1978:PSMb

[Ano78p] Anonymous. Program of the 39th Statistical Mechanics Meeting. Journal of Statistical Physics, 18(6):639–643, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014484.

Anonymous:1978:PSMc

[Ano78q] Anonymous. Program of the 40th statistical mechanics meeting. Journal of Statistical Physics, 19(6):657–662, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011776.

Anonymous:1979:FCBa

[Ano79a] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(1):121, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013750.

Anonymous:1979:FCBb

[Ano79b] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(2):245, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011515.

Anonymous:1979:FCBc

[Ano79c] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(3):349, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011943.

Anonymous:1979:FCBd

[Ano79d] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(4):471, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011783.

Anonymous:1979:FCBe

[Ano79e] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(5):571, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012900.

Anonymous:1979:FCBf

[Ano79f] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 20(6):731, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009523.

Anonymous:1979:FCBg

[Ano79g] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(1):101–102, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011484.

Anonymous:1979:FCBh

[Ano79h] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(2):229, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008700.

Anonymous:1979:FCBi

[Ano79i] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(3):345, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011476.

Anonymous:1979:FCBj

[Ano79j] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(4):495, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009613.

Anonymous:1979:FCBk

[Ano79k] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(5):613, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011173.

Anonymous:1979:FCBl

[Ano79l] Anonymous. Future contributions to Journal of Statistical Physics. Journal of Statistical Physics, 21(6):747, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01107913.

Anonymous:1979:IA

[Ano79m] Anonymous. Important announcement. Journal of Statistical Physics, 20(1):1, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013742.

Anonymous:1979:PSM

[Ano79n] Anonymous. Program of the 41st Statistical Mechanics Meeting. Journal of Statistical Physics, 20(6):733–737, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009524.

Aono:1972:PEM

[Aon72] Osamu Aono. Phenomenological equations for multicomponent fluids. Journal of Statistical Physics, 5(1–2):113–125, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008374.

Aono:1975:TCD

[Aon75] Osamu Aono. Thermodynamic coupling of diffusion with chemical reaction. Journal of Statistical Physics, 13(4):331–335, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012011.

Altenberger:1972:KEM

[AS72] A. R. Altenberger and J. Stecki. Kinetic equations without ‘memory’ for the time-displaced correlation functions. Journal of Statistical Physics, 5(1–2):83–97, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008372.

Amit:1973:CFW

[AS73] Daniel J. Amit and A. Shcherbakov. The correlation function in Wilson’s theory of phase transitions: Above and below the transition. Journal of Statistical Physics, 9(4):317–329, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012164.

Aizenman:1979:PMS

[AS79] Michael Aizenman and Herbert Spohn. Probabilistic methods for stationary problems of linear transport theory. Journal of Statistical Physics, 21(1):23–32, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011478.

Ah-Sam:1971:UVM

[ASJS71] L. E. G. Ah-Sam, H. Højgaard Jensen, and H. Smith. On the use of variational methods for solving Boltzmann equations involving non-Hermitian operators. Journal of Statistical Physics, 3(1):17–34, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012184.

Amit:1973:FTD

[AZ73a] Daniel J. Amit and Marco Zannetti. Field theory description of continuous phase transitions. Journal of Statistical Physics, 7(1):31–63, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01406131.

Amit:1973:SCT

[AZ73b] Daniel J. Amit and Marco Zannetti. Self-consistent treatment of a phase transition. Journal of Statistical Physics, 9 (1):1–21, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016794.

Amit:1974:ATD

[AZ74] Daniel J. Amit and Marco Zannetti. Approach to two dimensions in the n-vector model. Journal of Statistical Physics, 11 (2):133–141, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012515.

Barber:1974:CBS

[Bar74] Michael N. Barber. Critical behavior of a spherical model with a free surface. Journal of Statistical Physics, 10(1):59–88, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011717.

Barrett:1977:ISM

[Bar77] Wayne W. Barrett. An intermediate spherical model of a ferromagnet. Journal of Statistical Physics, 16(4):371–394, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01020429.

Batty:1974:CPC

[Bat74] Michael Batty. A comment on the paper ‘A comparison of the Shannon and Kullback information measures’. Journal of Statistical Physics, 11(6):523–524, December 1974. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01008893. See [HC73].

Baus:1978:CMD

[Bau78] Marc Baus. Collective modes and dynamic structure factor of a two-dimensional electron fluid. Journal of Statistical Physics, 19 (2):163–176, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012509.

Baxter:1973:ADM

[Bax73a] Rodney Baxter. Asymptotically degenerate maximum eigenvalues of the eight-vertex model transfer matrix and interfacial tension. Journal of Statistical Physics, 8(1):25–55, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008441.

Baxter:1973:SSP

[Bax73b] Rodney J. Baxter. Spontaneous staggered polarization of the F -model. Journal of Statistical Physics, 9(2):145–182, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016845.

Baxter:1976:CTM

[Bax76] R. J. Baxter. Corner transfer matrices of the eight-vertex model. I. Low-temperature expansions and conjectured properties. Journal of Statistical Physics, 15(6):485–503, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020802.

Baxter:1977:CTM

[Bax77] R. J. Baxter. Corner transfer matrices of the eight-vertex model. II. The Ising model case. Journal of Statistical Physics, 17(1):1– 14, July 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01089373.

Baxter:1978:VAS

[Bax78] R. J. Baxter. Variational approximations for square lattice models in statistical mechanics. Journal of Statistical Physics, 19(5): 461–478, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011693.

Baumgartner:1978:DGG

[BB78] A. Baumgärtner and K. Binder. Dynamics of the generalized Glauber–Ising chain in a magnetic field. Journal of Statistical Physics, 18(5):423–444, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014516.

Barber:1975:EBE

[BBO75] Michael N. Barber, John M. Blatt, and Alex H. Opie. On the extension of the Boltzmann equation to include pair correlations.

Journal of Statistical Physics, 13(5):379–384, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013114.

Baxter:1979:SEC

[BE79] R. J. Baxter and I. G. Enting. Series expansions from corner transfer matrices: The square lattice Ising model. Journal of Statistical Physics, 21(2):103–123, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008694.

Berman:1979:FDT

[Ber79a] D. H. Berman. The fluctuation-dissipation theorem for contracted descriptions of Markov processes. Journal of Statistical Physics, 20(1):57–81, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013746.

Bernu:1979:SPM

[Ber79b] B. Bernu. Single-particle motion in the strongly coupled one-component plasma. Journal of Statistical Physics, 21(4): 447–464, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009611.

Braun:1973:CNS

[BF73] E. Braun and A. Flores. A convergent nonequilibrium statistical mechanical theory for dense gases. I. The two-body distribution function. Journal of Statistical Physics, 8(2):155–165, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008536.

Braun:1974:CNS

[BFC74] E. Braun, A. Flores, and G. Coutiño. A convergent nonequilibrium statistical mechanical theory for dense gases. III. Transport coefficients to second order in the density. Journal of Statistical Physics, 10(1):49–58, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011716.

Braun:1973:NSM

[BFGC73] E. Braun, A. Flores, and L. S. García-Colín. On the nonequilibrium statistical mechanics of a binary mixture. I. The distribution functions. Journal of Statistical Physics, 8(3):239–248, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012569.

Basano:1971:QA

[BG71] L. Basano and A. Gamba. Questions and answers. Journal of Statistical Physics, 3(2):237–241, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019852.

Bergeron:1975:MOB

[BG75] Kenneth D. Bergeron and Eugene P. Gross. Modified onebody memory function for classical fluids. Journal of Statistical Physics, 13(1):85–100, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012601.

Barajas:1973:ETT

[BGCP73] L. Barajas, L. S. García-Colín, and E. Piña. On the Enskog– Thorne theory for a binary mixture of dissimilar rigid spheres. Journal of Statistical Physics, 7(2):161–183, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024213.

Bergeron:1974:ASL

[BGV74] Kenneth D. Bergeron, Eugene P. Gross, and Rodney L. Varley. Approximate solutions of the Liouville equation. IV. The twobody additive approximation. Journal of Statistical Physics, 10 (2):111–138, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009716.

Blum:1978:SOZ

[BH78] L. Blum and J. S. Høye. Solution of the Ornstein–Zernike equation with Yukawa closure for a mixture. Journal of Statistical Physics, 19(4):317–324, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011750.

Bergman:1973:ESM

[BIG73] David J. Bergman, Yoseph Imry, and Leon Gunther. Exactly soluble magnetoelastic lattice with a magnetic phase transition. Journal of Statistical Physics, 7(4):337–360, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014909.

Biswas:1972:AAV

[Bis72] A. C. Biswas. Action-angle variables in quantum statistical mechanics. Journal of Statistical Physics, 5(3):173–183, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023740.

Biswas:1973:GLP

[Bis73] A. C. Biswas. On the gas-liquid phase transition. Journal of Statistical Physics, 7(2):131–142, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01024211. See cri-

tique [Gro73a].

Blander:1972:TCF

[BK72] Milton Blander and Joseph L. Katz. The thermodynamics of cluster formation in nucleation theory. Journal of Statistical Physics, 4(1):55–59, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008471.

Blatt:1975:NDG

[Bla75] John M. Blatt. A new dilute gas transport phenomenon. Journal of Statistical Physics, 13(5):389–392, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013116.

Blood:1970:ACT

[Blo70] Frank A. Blood, Jr. Approximate calculations for the twodimensional Ising model. Journal of Statistical Physics, 2(4): 301–304, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020439.

Blomberg:1973:GFI

[Blo73] Clas Blomberg. Generating functionals and the Ising model. Journal of Statistical Physics, 7(3):197–212, March 1973. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030303.

Bricmont:1979:EBC

[BLP79] Jean Bricmont, Joel L. Lebowitz, and Charles E. Pfister. On the equivalence of boundary conditions. Journal of Statistical Physics, 21(5):573–582, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011169.

Budgor:1976:SNSb

[BLS76] A. B. Budgor, K. Lindenberg, and K. E. Shuler. Studies in nonlinear stochastic processes. II. The Duffing oscillator revisited. Journal of Statistical Physics, 15(5):375–391, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020340.

Blum:1978:AHS

[Blu78a] L. Blum. Addendum to Henderson and Smith’s exact formulas for the pair correlation functions of charged hard spheres in the mean spherical approximation. Journal of Statistical Physics, 19 (2):201–202, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012512.

Blum:1978:SMS

[Blu78b] L. Blum. Solution of the mean spherical approximation for hard ions and dipoles of arbitrary size. Journal of Statistical Physics, 18(5):451–474, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014518.

Buffet:1978:DOB

[BM78] Emmanuel Buffet and Philippe A. Martin. Dynamics of the open BCS model. Journal of Statistical Physics, 18(6):585–632, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014481.

Buffet:1979:DOB

[BM79] Emmanuel Buffet and Philippe A. Martin. Dynamics of the open BCS model. Journal of Statistical Physics, 20(3):347, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011942.

Bedeaux:1971:LRT

[BMP71] Dick Bedeaux, Sava Milosevic, and Gerald Paul. Linear response theory for systems obeying the master equation. Journal of Statistical Physics, 3(1):39–46, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012186.

Ben-Naim:1973:MMA

[BN73] A. Ben-Naim. Mixture-model approach to the theory of classical fluids. III. Application to aqueous solutions of nonelectrolytes. Journal of Statistical Physics, 7(1):3–30, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01406130.

Brenner:1978:NFA

[BNW78] Stephen L. Brenner, Ralph J. Nossal, and George H. Weiss. Number fluctuation analysis of random locomotion. Statistics of a Smoluchowski process. Journal of Statistical Physics, 18 (1):1–18, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014667.

Blatt:1975:LRP

[BO75] John M. Blatt and Alex H. Opie. Long-range pair correlations in dilute gases. Journal of Statistical Physics, 13(5): 385–388, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013115.

Bortz:1974:SDM

[Bor74] Alfred B. Bortz. Spinodal decomposition in Monte Carlo simulations of a binary alloy. Journal of Statistical Physics, 11(3): 181–193, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010216.

Brezin:1978:CEL

[BP78] E. Brézin and G. Parisi. Critical exponents and large-order behavior of perturbation theory. Journal of Statistical Physics, 19(3):269–292, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011726.

Braun:1974:NSM

[Bra74a] E. Braun. Nonequilibrium statistical mechanics of systems interacting with nonadditive forces. III. Properties to first order in the density. Journal of Statistical Physics, 10(1):35–47, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011715.

Bray:1974:SMZ

[Bra74b] Alan J. Bray. Statistical mechanics of zero-dimensional Ginzburg–Landau fields: Accuracy of the screening approximation. Journal of Statistical Physics, 11(1):29–50, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019476.

Bricmont:1977:GIM

[Bri77] J. Bricmont. The Gaussian inequality for multicomponent rotators. Journal of Statistical Physics, 17(5):289–300, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014399.

Brush:1970:FBL

[Bru70] Stephen G. Brush. Francis Bitter and ‘Landau diamagnetism’. Journal of Statistical Physics, 2(2):195–197, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009738.

Brush:1978:RBR

[Bru78] S. G. Brush. Response to book review: The kind of motion we call heat. Journal of Statistical Physics, 18(5):523–528, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014521.

Bashkirov:1971:RCN

[BS71] A. G. Bashkirov and A. G. Sutugin. On a reconsideration of classical nucleation theory. Journal of Statistical Physics, 3(1):87–89, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012189.

Binder:1972:MCS

[BS72a] K. Binder and D. Stauffer. Monte Carlo study of the surface area of liquid droplets. Journal of Statistical Physics, 6(1): 49–59, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01060201.

Brostow:1972:IWG

[BS72b] Witold Brostow and Andrzej Schinzel. On interesting walks in a graph. Journal of Statistical Physics, 4(2–3):103–110, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008546.

Brostow:1973:SPI

[BS73] Witold Brostow and Yvon Sicotte. Some properties of the informational model of the liquid state. Journal of Statistical Physics, 9(4):339–349, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012166.

Blum:1976:SOZ

[BS76] L. Blum and G. Stell. Solution of Ornstein–Zernike equation for wall-particle distribution function. Journal of Statistical Physics, 15(6):439–449, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020798.

Burkhardt:1979:OTP

[BS79] Theodore W. Burkhardt and B. W. Southern. Optimal truncation procedures in renormalization-group calculations. Journal of Statistical Physics, 20(3):247–257, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011937.

Bedeaux:1970:DCI

[BSO70] Dick Bedeaux, Kurt E. Shuler, and Irwin Oppenheim. Decay of correlations. III. Relaxation of spin correlations and distribution functions in the one-dimensional Ising lattice. Journal of Statistical Physics, 2(1):1–19, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009708.

Bedeaux:1971:DCI

[BSO71] Dick Bedeaux, Kurt E. Shuler, and Irwin Oppenheim. Decay of correlations. IV. Necessary and sufficient conditions for a rapid decay of correlations. Journal of Statistical Physics, 3 (4):365–379, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008274.

Brankov:1979:CES

[BTZ79] J. G. Brankov, N. S. Tonchev, and V. A. Zagrebnov. On a class of exactly soluble statistical mechanical models with nonpolynomial interactions. Journal of Statistical Physics, 20(3):317–330, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011940.

Budgor:1976:SNSa

[Bud76] Aaron B. Budgor. Studies in nonlinear stochastic processes. I. Approximate solutions of nonlinear stochastic differential equations by the method of statistical linearization. Journal of Statistical Physics, 15(5):355–374, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020339.

Budgor:1977:SNS

[Bud77] Aaron B. Budgor. Studies in nonlinear stochastic processes. III. Approximate solutions of nonlinear stochastic differential equations excited by Gaussian noise and harmonic disturbances. Journal of Statistical Physics, 17(1):21–44, July 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01089375.

Butz:1972:TN

[But72] Arthur R. Butz. A theory of 1/f noise. Journal of Statistical Physics, 4(2–3):199–216, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008550.

Begum:1979:DHN

[BV79] Salma Begum and P. B. Visscher. Discrete hydrodynamics: Numerical results for the soft-sphere viscosity. Journal of Statistical Physics, 20(6):641–655, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009515.

Baus:1977:LTB

[BW77] M. Baus and J. Wallenborn. Long-time behavior of the nonlocal shear viscosity of a one-component plasma: A microscopic approach. Journal of Statistical Physics, 16(1):91–111, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014607.

Bixon:1971:BMN

[BZ71] Mordechai Bixon and Robert Zwanzig. Brownian motion of a nonlinear oscillator. Journal of Statistical Physics, 3(3):245–260, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011383.

Carmeli:1974:STE

[Car74] M. Carmeli. Statistical theory of energy levels and random matrices in physics. Journal of Statistical Physics, 10(4):259–297, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012252.

Caroli:1979:DBP

[CCR79] B. Caroli, C. Caroli, and B. Roulet. Diffusion in a bistable potential: A systematic WKB treatment. Journal of Statistical Physics, 21(4):415–437, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009609.

Chappell:1971:KTS

[CCSD71] W. R. Chappell, J. Cooper, E. W. Smith, and T. Dillon. A kinetic theory of spectral line shapes. Journal of Statistical Physics, 3(4):401–410, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008278.

Combis:1979:IBQ

[CFTA79] P. Combis, J. Fronteau, and A. Tellez-Arenas. Introduction to a Brownian quasiparticle model. Journal of Statistical Physics, 21 (4):439–446, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009610.

Chaturvedi:1978:PRI

[CG78] S. Chaturvedi and C. W. Gardiner. The Poisson representation. II. Two-time correlation functions. Journal of Statistical Physics, 18(5):501–522, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014520.

Chaturvedi:1977:SAC

[CGMW77] S. Chaturvedi, C. W. Gardiner, I. S. Matheson, and D. F. Walls. Stochastic analysis of a chemical reaction with spatial and temporal structures. Journal of Statistical Physics, 17(6): 469–489, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014350.

Clark:1973:AGT

[CGWE73] P. M. Clark, P. Gluck, and O. Wohlman-Entin. Application of general transport theories to model system. Journal of Statistical Physics, 7(4):361–371, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014910.

Cercignani:1979:GSB

[CGZ79] Carlo Cercignani, William Greenberg, and P. F. Zweifel. Global solutions of the Boltzmann equation on a lattice. Journal of Statistical Physics, 20(4):449–462, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011781.

Cooper:1979:DPC

[CH79] R. D. Cooper and M. R. Hoare. Distributive processes and combinatorial dynamics. Journal of Statistical Physics, 20(6):597– 628, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009513.

Chappell:1970:KEA

[Cha70] W. R. Chappell. Kinetic equations for autocorrelation functions in dilute gases. Journal of Statistical Physics, 2(3):267–277, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030746.

Chay:1972:BSM

[Cha72] Teresa Ree Chay. Bead-spring model of a polymer chain: A transient solution. Journal of Statistical Physics, 6(2–3): 73–86, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023680.

Chow:1975:FPI

[Cho75] Pao-Liu Chow. A functional phase-integral method and applications to the laser beam propagation in random media. Journal of Statistical Physics, 12(2):93–109, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01010026.

Chu:1972:CUC

[Chu72] B. Chu. Comments on universality of critical exponents for fluid systems. Journal of Statistical Physics, 6(2–3):173–177, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023685.

Ciccotti:1979:TEM

[CJM79] G. Ciccotti, G. Jacucci, and I. R. McDonald. ‘Thoughtexperiments’ by molecular dynamics. Journal of Statistical Physics, 21(1):1–22, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011477.

Caroli:1973:VAF

[CJSJT73] B. Caroli, G. Jannink, D. Saint-James, and D. Taupin. Velocity autocorrelation function for the motion of long-chain molecules in the free draining limit. Journal of Statistical Physics, 8(4): 367–371, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020275.

Carbonell:1974:TDP

[CK74] R. G. Carbonell and M. D. Kostin. Time-dependent probability density of statistical mechanics. Journal of Statistical Physics, 11 (5):379–390, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026729.

Chirikov:1971:SMD

[CKS71] B. V. Chirikov, E. Keil, and A. M. Sessler. Stochasticity in manydimensional nonlinear oscillating systems. Journal of Statistical Physics, 3(3):307–321, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011386.

Cukier:1975:EMR

[CL75] Robert I. Cukier and Katja Lindenberg. Equations of motion for rotational motion in liquids. Journal of Statistical Physics, 13 (5):451–455, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013119.

Cukier:1973:OPS

[CLLS73] R. I. Cukier, K. Lakatos-Lindenberg, and K. E. Shuler. Orthogonal polynomial solutions of the Fokker–Planck equation. Journal of Statistical Physics, 9(2):137–144, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016844.

Coopersmith:1974:IBI

[CM74] Michael Coopersmith and George Mandeville. Irreversible behavior of interacting systems. I. The approach to equilibrium. Journal of Statistical Physics, 10(5):391–403, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008801.

Carbonell:1979:TTP

[CM79] R. G. Carbonell and B. J. McCoy. Tunneling and transport problems for a quantum mechanical Brownian particle. Journal of Statistical Physics, 21(3):301–311, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011472.

Cassandro:1977:RSM

[CO77] M. Cassandro and E. Olivieri. A rigorous study of metastability in a continuous model. Journal of Statistical Physics, 17 (4):229–244, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01040104.

Cohen:1970:AAC

[Coh70] E. Richard Cohen. The accuracy of the approximations in classical nucleation theory. Journal of Statistical Physics, 2(2):147– 152, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009736.

Cohen:1975:BLP

[Coh75] E. Richard Cohen. Bandwidth limitations in pulse code modulation. Journal of Statistical Physics, 13(1):65–84, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012600.

Coopersmith:1979:EEI

[Coo79] Michael Coopersmith. An example of the effect of irreversible microscopic dynamics on the macroscopic behavior of a large system. Journal of Statistical Physics, 20(2):175–179, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011511.

Cichocki:1975:MAE

[CP75] B. Cichocki and J. Piasecki. Microscopic approach to the Enskog theory of a homogeneous gas. Journal of Statistical Physics, 12 (2):111–129, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010027.

Coste:1979:SKM

[CPC79] J. Coste, J. Peyraud, and P. Coullet. Simple kinetic model of symmetry breaking. Journal of Statistical Physics, 21(1):33–50, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011479.

Critchley:1976:VAS

[CS76] R. H. Critchley and A. I. Solomon. A variational approach to superfluidity. Journal of Statistical Physics, 14(4):381–393, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030201.

Cukier:1972:VSR

[CSW72] R. I. Cukier, K. E. Shuler, and J. D. Weeks. On the validity of stochastic rate equations in finite systems with finite-strength interactions. Journal of Statistical Physics, 5(1–2):99–112, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008373.

Cummings:1979:SOZ

[CWPS79] P. T. Cummings, C. C. Wright, J. W. Perram, and E. R. Smith. Solution of the Ornstein–Zernike equation for a soft-core Yukawa fluid. Journal of Statistical Physics, 21(6):659–667, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01107908.

Davis:1973:KTD

[Dav73] H. Ted Davis. A kinetic theory of dense fluids. Journal of Statistical Physics, 7(3):225–241, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01030305.

Davies:1978:MHC

[Dav78] E. B. Davies. A model of heat conduction. Journal of Statistical Physics, 18(2):161–170, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014307.

Davis:1971:ETL

[DB71] H. Ted Davis and Marc O. Baleiko. Existence theory of the linear equations appearing in the Chapman–Enskog solutions to two kinetic equations for liquids. Journal of Statistical Physics, 3(1):47–68, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012187.

deBoer:1978:DQK

[dBvW78] W. P. H. de Boer and Ch. G. van Weert. On the derivation of quantum kinetic equations. II. Nonlocal Uehling–Uhlenbeck equation. Journal of Statistical Physics, 18(3):281–292, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01018094.

Davis:1972:SER

[DD72a] H. Ted Davis and Gary R. Dowling. Some exact results in kinetic theory. Journal of Statistical Physics, 4(2–3):193–197, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008549.

Dowling:1972:SMR

[DD72b] Gary R. Dowling and H. Ted Davis. On the statistical mechanics of rigid particles. Journal of Statistical Physics, 4(1):1–13, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008468.

DeGennes:1975:BMC

[De 75] P. G. De Gennes. Brownian motion of a classical particle through potential barriers. Application to the helix-coil transitions of heteropolymers. Journal of Statistical Physics, 12(6):463–481, June 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01025834.

Delyon:1979:DCI

[Del79] F. Delyon. The distribution of clusters for the Ising model. Journal of Statistical Physics, 21(6):727–738, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01107911.

Dresden:1972:MNQ

[DF72] Max Dresden and Frank Feiock. Models in nonequilibrium quantum statistical mechanics. Journal of Statistical Physics, 4(2– 3):111–173, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008547.

Dufty:1979:NTB

[DL79] James W. Dufty and Michael J. Lindenfeld. Nonlinear transport in the Boltzmann limit. Journal of Statistical Physics, 20 (3):259–301, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011938.

Driessler:1979:ECL

[DLP79] W. Driessler, L. Landau, and J. Fernando Perez. Estimates of critical lengths and critical temperatures for classical and quantum lattice systems. Journal of Statistical Physics, 20(2): 123–162, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011509.

Dyson:1978:PTQ

[DLS78] Freeman J. Dyson, Elliott H. Lieb, and Barry Simon. Phase transitions in quantum spin systems with isotropic and nonisotropic interactions. Journal of Statistical Physics, 18(4):335–383, April 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01106729.

Dodonov:1977:IND

[DMM77] V. V. Dodonov, I. A. Malkin, and V. I. Man’ko. Invariants and nonequilibrium density matrices. Journal of Statistical Physics, 16(4):357–370, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020428.

DelleDonne:1978:GNM

[DO78] M. DelleDonne and P. Ortoleva. Generalized nonlinear master equations for local fluctuations: Dimensionality expansions and augmented mean field theory. Journal of Statistical Physics, 18 (3):319–331, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018097.

DelleDonne:1979:CFU

[DO79] M. DelleDonne and P. Ortoleva. Critical fluctuation universality in chemically oscillatory systems: A soluble master equation. Journal of Statistical Physics, 20(5):473–486, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012895.

delRio:1978:UAD

[dRGC78] J. L. del Rio and L. S. García-Colín. A unified approach for deriving kinetic equations in nonequilibrium statistical mechanics. II. Approximate results. Journal of Statistical Physics, 19 (2):109–127, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012506.

Davies:1978:OQS

[DS78] E. B. Davies and H. Spohn. Open quantum systems with timedependent Hamiltonians and their linear response. Journal of Statistical Physics, 19(5):511–523, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011696.

Dufour:1977:NEE

[Duf77] C. G. Dufour. Nonequilibrium expressions for entropy and other thermodynamic quantities. Journal of Statistical Physics, 17 (2):61–70, August 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01268918.

Dunlop:1977:ZPF

[Dun77] François Dunlop. Zeros of partition functions via correlation inequalities. Journal of Statistical Physics, 17(4):215–228, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01040103.

Dunlop:1979:ZPF

[Dun79] François Dunlop. Zeros of the partition function and Gaussian inequalities for the plane rotator model. Journal of Statistical Physics, 21(5):561–572, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011168.

Dutta:1972:SMG

[Dut72] M. Dutta. Statistical mechanics and the Gauss principle. Journal of Statistical Physics, 5(1–2):127–131, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008375.

Desai:1978:SMN

[DZ78] Rashmi C. Desai and Robert Zwanzig. Statistical mechanics of a nonlinear stochastic model. Journal of Statistical Physics, 19(1):1–24, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020331.

Ernst:1978:KTN

[ECD^{+}78] M. H. Ernst, B. Cichocki, J. R. Dorfman, J. Sharma, and H. van Beijeren. Kinetic theory of nonlinear viscous flow in two and three dimensions. Journal of Statistical Physics, 18(3):237–270, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018092.

Ernst:1975:NDR

[ED75] M. H. Ernst and J. R. Dorfman. Nonanalytic dispersion relations for classical fluids. Journal of Statistical Physics, 12(4):311–359, April 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012067.

Eggarter:1974:BPA

[Egg74] T. P. Eggarter. Bethe–Peierls approximation for the disordered Ising model. Journal of Statistical Physics, 11(5):363– 377, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026728.

Egelstaff:1975:TIN

[EGGM75] P. A. Egelstaff, C. G. Gray, K. E. Gubbins, and K. C. Mo. Theory of inelastic neutron scattering from molecular fluids. Journal of Statistical Physics, 13(4):315–330, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012010.

Ernst:1976:ATBa

[EHvL76a] M. H. Ernst, E. H. Hauge, and J. M. J. van Leeuwen. Asymptotic time behavior of correlation functions. II. Kinetic and potential terms. Journal of Statistical Physics, 15(1):7–22, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012807.

Ernst:1976:ATBb

[EHvL76b] M. H. Ernst, E. H. Hauge, and J. M. J. van Leeuwen. Asymptotic time behavior of correlation functions. III. Local equilibrium and mode-coupling theory. Journal of Statistical Physics, 15(1):23– 58, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012808.

Ellis:1977:MDC

[Ell77] Richard S. Ellis. Monotone decrease of characteristic functions. Journal of Statistical Physics, 16(1):117–118, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01014609.

Ellis:1978:MDC

[Ell78] Richard S. Ellis. Monotone decrease of characteristic functions. Journal of Statistical Physics, 18(1):107, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014673.

Ellis:1978:SCW

[EN78] Richard S. Ellis and Charles M. Newman. The statistics of Curie–Weiss models. Journal of Statistical Physics, 19(2): 149–161, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012508.

Evans:1979:NPT

[Eva79] Denis J. Evans. The nonsymmetric pressure tensor in polyatomic fluids. Journal of Statistical Physics, 20(5):547–555, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012898.

Frisch:1975:QMS

[FA75] Harry L. Frisch and Saeyoung Ahn. Quantum mechanical scaled particle theory of hard spheres at high temperature. Journal of Statistical Physics, 13(2):129–136, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01221373.

Flores:1973:CNS

[FB73] A. Flores and E. Braun. A convergent nonequilibrium statistical mechanical theory for dense gases. II. Transport coefficients to first order in the density. Journal of Statistical Physics, 8(2):167– 176, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008537.

Fox:1978:HCR

[FB78] Ronald Forrest Fox and Richard Barakat. Heat conduction in a random medium. Journal of Statistical Physics, 18(2):171–178, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014308.

Fleming:1973:CPP

[FBB73] Paul D. Fleming, Bruce J. Berne, and Marvin Bishop. Collective properties of polyatomic gases in a magnetic field: Effects on light scattering spectrum. Journal of Statistical Physics, 7 (4):311–328, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014907.

Feigenbaum:1978:QUC

[Fei78] Mitchell J. Feigenbaum. Quantitative universality for a class of nonlinear transformations. Journal of Statistical Physics, 19(1):25–52, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020332.

Feigenbaum:1979:UMP

[Fei79] Mitchell J. Feigenbaum. The universal metric properties of nonlinear transformations. Journal of Statistical Physics, 21(6): 669–706, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01107909.

Felderhof:1972:TDS

[Fel72] B. U. Felderhof. Time-dependent statistics of binary linear lattices. Journal of Statistical Physics, 6(1):21–38, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060199.

Ferrell:1973:FTT

[Fer73] Richard A. Ferrell. Field theory of the two-dimensional Ising model: Equivalence to the free particle one-dimensional Dirac equation. Journal of Statistical Physics, 8(3):265–278, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01012571. See addendum [Fer74].

Ferrell:1974:AFT

[Fer74] Richard A. Ferrell. Addendum to ‘Field theory of the twodimensional Ising model: Equivalence to the free particle onedimensional Dirac equation’. Journal of Statistical Physics, 10 (1):99–101, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011719. See [Fer73].

Fesciyan:1973:EVA

[FF73] S. Fesciyan and H. L. Frisch. The exact velocity autocorrelation function of a model system. Journal of Statistical Physics, 8 (4):353–358, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020273.

Fesciyan:1975:SMOa

[FF75a] Sezar Fesciyan and H. L. Frisch. Stochastic model of a onedimensional fluid. Journal of Statistical Physics, 12(2):153–162, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010030.

Fesciyan:1975:SMOb

[FF75b] Sezar Fesciyan and H. L. Frisch. Stochastic model of a onedimensional fluid. Journal of Statistical Physics, 13(3):277– 278, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012843.

Fleming:1974:ALGa

[FG74a] Paul D. Fleming III and Julian H. Gibbs. An adaptation of the lattice gas to the water problem. Journal of Statistical Physics, 10(2):157–173, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009718.

Fleming:1974:ALGb

[FG74b] Paul D. Fleming III and Julian H. Gibbs. An adaptation of the lattice gas to the water problem. II. Second-order approximation. Journal of Statistical Physics, 10(5):351–378, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008799.

Fleming:1976:AST

[FG76] Paul D. Fleming III and Julian H. Gibbs. Application of stochastic theory to dilute gas reactions. Journal of Statistical Physics, 14(3):271–289, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023629.

Fisch:1978:CTT

[Fis78] R. Fisch. Critical temperature for two-dimensional Ising ferromagnets with quenched bond disorder. Journal of Statistical Physics, 18(2):111–114, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014302.

Fulinski:1979:ODM

[FL79] A. Fuliński and L. Longa. One-dimensional model with rotational and liquid-crystalline phase transitions. Journal of Statistical Physics, 21(6):635–647, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01107906.

Flinn:1974:MCC

[Fli74] P. A. Flinn. Monte Carlo calculation of phase separation in a two-dimensional Ising system. Journal of Statistical Physics, 10 (1):89–97, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011718.

Fox:1977:ANG

[Fox77] Ronald Forrest Fox. Analysis of nonstationary, Gaussian and non-Gaussian, generalized Langevin equations using methods of multiplicative stochastic processes. Journal of Statistical Physics, 16(3):259–279, March 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020382.

Fox:1979:NES

[Fox79] Jeffrey R. Fox. Nonclassical equations of state for critical and tricritical points. Journal of Statistical Physics, 21(3): 243–261, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011468.

Friedman:1973:RTM

[Fri73] Kenneth Friedman. Replies to tribus and motroni and to gage and Hestenes. Journal of Statistical Physics, 9(3):265– 269, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008732.

Frisch:1977:KMT

[Fri77] H. L. Frisch. On the Kirkwood-modified tait equation. Journal of Statistical Physics, 17(2):89–94, August 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01268920.

Fritz:1979:SDT

[Fri79] J. Fritz. Stochastic dynamics of two-dimensional infinite-particle systems. Journal of Statistical Physics, 20(4):351–369, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011777.

Fenn:1973:STF

[FRR73] John B. Fenn, Jr., Philip K. Rawlings, and Howard Reiss. Solution of the Thomas–Fermi problem for continuous, spherical, positive charge distribution. Journal of Statistical Physics, 8 (4):335–339, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020271.

Friedman:1971:JME

[FS71] Kenneth Friedman and Abner Shimony. Jaynes’s maximum entropy prescription and probability theory. Journal of Statistical Physics, 3(4):381–384, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008275.

Frey:1974:NCH

[FSV74] J. Frey, J. Salmon, and M. Valton. A new closure hypothesis for the BBGKY system of equations. Journal of Statistical Physics, 11(6):457–474, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008889.

Franceschini:1979:SIB

[FT79] Valter Franceschini and Claudio Tebaldi. Sequences of infinite bifurcations and turbulence in a five-mode truncation of the Navier–Stokes equations. Journal of Statistical Physics, 21(6): 707–726, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01107910.

Fujiwara:1970:FPG

[FtHW70] I. Fujiwara, D. ter Haar, and H. Wergeland. Fluctuations in the population of the ground state of Bose systems. Journal of Statistical Physics, 2(4):329–346, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020442.

Frisch:1970:BR

[FTM70] H. L. Frisch, A. J. Tether, and C. F. Martin. Book reviews. Journal of Statistical Physics, 2(2):199–200, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009739.

Furukawa:1974:SMT

[Fur74] Hiroshi Furukawa. Statistical mechanical theory of the nonlinear steady state. Journal of Statistical Physics, 10(2):139–156, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009717.

Futrelle:1971:N

[Fut71] R. P. Futrelle. News. Journal of Statistical Physics, 3(4):419–421, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008281.

Gardiner:1976:CCL

[Gar76] C. W. Gardiner. A comment on chemical Langevin equations. Journal of Statistical Physics, 15(6):451–454, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020799.

Gates:1976:AFR

[Gat76] D. J. Gates. An alternative formulation of rigorous mean-field theory. Journal of Statistical Physics, 15(6):513–516, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020804.

Garcia-Colin:1973:PCE

[GC73] L. S. García-Colín. Poincaré cycles and ergodic behavior of a linear diatomic chain. Journal of Statistical Physics, 7(3):243–258, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030306.

Gardiner:1977:PRN

[GC77] C. W. Gardiner and S. Chaturvedi. The Poisson representation. I. A new technique for chemical master equations. Journal of Statistical Physics, 17(6):429–468, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014349.

Garcia-Colin:1979:MDM

[GC79] L. S. Garcia-Colin. On the microscopic derivation of mode-mode coupling equations. Journal of Statistical Physics, 20(1):19–24, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013744.

Garcia-Colin:1977:UAD

[GCdR77] Leopoldo S. García-Colín and Jose Luis del Rio. A unified approach for deriving kinetic equations in nonequilibrium statistical mechanics. I. Exact results. Journal of Statistical Physics, 16(3):235–258, March 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020381.

Garcia-Colin:1973:NSM

[GCFB73] L. S. García-Colín, A. Flores, and E. Braun. On the nonequilibrium statistical mechanics of a binary mixture. II. The transport coefficients. Journal of Statistical Physics, 8(3):249–264, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012570.

Gibbs:1975:STR

[GF75] Julian H. Gibbs and Paul D. Fleming III. Stochastic theory for rate constants of chemical reactions in liquid solution. Journal of Statistical Physics, 12(5):375–401, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012884.

Gage:1973:CPJ

[GH73] Douglas W. Gage and David Hestenes. Comment on the paper ‘Jaynes’s maximum entropy prescription and probability theory’. Journal of Statistical Physics, 7(1):89–90, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01406133.

Gunther:1974:BEC

[GIB74] L. Gunther, Yoseph Imry, and David J. Bergman. Bose–Einstein condensation in one- and two-dimensional gases. Journal of Statistical Physics, 10(4):299–319, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012253.

Gillespie:1977:CVS

[Gil77] Daniel T. Gillespie. Concerning the validity of the stochastic approach to chemical kinetics. Journal of Statistical Physics, 16 (3):311–318, March 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020385.

Girardeau:1978:ESB

[Gir78] M. D. Girardeau. Equilibrium superradiance in a Bose gas. Journal of Statistical Physics, 18(2):207–215, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014311.

Gutkowicz-Krusin:1978:ARA

[GKPR78] Dina Gutkowicz-Krusin, Itamar Procaccia, and John Ross. Analytic results for asymmetric random walk with exponential transition probabilities. Journal of Statistical Physics, 19(5): 525–541, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011697.

Golden:1972:NFD

[GKS72] K. I. Golden, G. Kalman, and M. B. Silevitch. Nonlinear fluctuation-dissipation theorem. Journal of Statistical Physics, 6(2–3):87–118, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023681.

Glasser:1975:SES

[Gla75] M. L. Glasser. A singular example in the statistical mechanics of a one-dimensional system. Journal of Statistical Physics, 13 (4):373–374, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012014.

Garrido:1979:SQD

[GLS79] L. Garrido, D. Lurié, and M. San Miguel. Stochastic quantization and detailed balance in Fokker–Planck dynamics. Journal of Statistical Physics, 21(3):313–335, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011473.

Gardiner:1976:CST

[GMWM76] C. W. Gardiner, K. J. McNeil, D. F. Walls, and I. S. Matheson. Correlations in stochastic theories of chemical reactions. Journal of Statistical Physics, 14(4):307–331, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01030197.

Grossberg:1971:STO

[GP71] Stephen Grossberg and James Pepe. Spiking threshold and overarousal effects in serial learning. Journal of Statistical Physics, 3(2):95–125, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01019845.

Gervois:1976:DPC

[GP76a] A. Gervois and Y. Pomeau. Dynamical perturbation for classical fluids: A solvable model. Journal of Statistical Physics, 14 (3):207–248, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023626.

Gervois:1976:ADS

[GP76b] A. Gervois and Y. Pomeau. On the absence of diffusion in a semiinfinite one-dimensional system. Journal of Statistical Physics, 14(6):483–500, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012847.

Griffiths:1979:MPP

[GP79] Robert B. Griffiths and Paul A. Pearce. Mathematical properties of position-space renormalization-group transformations. Journal of Statistical Physics, 20(5):499–545, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012897.

Gould:1979:AFM

[GPE79] Harvey Gould, R. G. Palmer, and Gentil A. Estévez. Asymptotic form of the mean spherical approximation for the internal energy of the classical one-component plasma. Journal of Statistical Physics, 21(1):55–64, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011481.

Gardner:1979:IVG

[GR79] Clifford S. Gardner and Charles Radin. The infinite-volume ground state of the Lennard-Jones potential. Journal of Statistical Physics, 20(6):719–724, June 1979. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009521.

Grabert:1978:NTD

[Gra78] Hermann Grabert. Nonlinear transport and dynamics of fluctuations. Journal of Statistical Physics, 19(5):479–497, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011694.

Griffiths:1970:BR

[Gri70] Robert B. Griffiths. Book review. Journal of Statistical Physics, 2(4):387, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020445.

Grmela:1971:KEA

[Grm71] M. Grmela. Kinetic equation approach to phase transitions. Journal of Statistical Physics, 3(3):347–364, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011389.

Groeneveld:1973:CPG

[Gro73a] J. Groeneveld. Critique of the paper ‘On the gas-liquid phase transition’. Journal of Statistical Physics, 9(2):183–185, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016846. See [Bis73].

Gross:1973:ASLa

[Gro73b] Eugene P. Gross. Approximate solutions of the Liouville equation. II. Stationary variational principles. Journal of Statistical Physics, 9(4):275–295, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012162.

Gross:1973:ASLb

[Gro73c] Eugene P. Gross. Approximate solutions of the Liouville equation. III. Variational principles and projection operators. Journal of Statistical Physics, 9(4):297–315, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012163.

Gross:1974:MFD

[Gro74] Eugene P. Gross. Memory function for dressed particles. Journal of Statistical Physics, 11(6):503–521, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008892.

Gross:1976:FSK

[Gro76] Eugene P. Gross. Formal structure of kinetic theory. Journal of Statistical Physics, 15(3):181–214, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012876.

Gross:1977:PFE

[Gro77] Eugene P. Gross. Partition function for an electron in a random potential. Journal of Statistical Physics, 17(5):265–287, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014398.

Gross:1979:PIL

[Gro79] Eugene P. Gross. Path integrals and lower bounds for density matrices. Journal of Statistical Physics, 21(2):215–227, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008699.

Gruber:1976:DPT

[Gru76] C. Gruber. On the definition of phase transition. Journal of Statistical Physics, 14(1):81–86, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01020136.

Gates:1975:CFS

[GT75] D. J. Gates and C. J. Thompson. Correlation functions for spin systems of high spatial dimensionality. Journal of Statistical Physics, 13(3):219–230, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012839.

Guderley:1973:FRI

[Gud73] K. G. Guderley. A formula related to irreducible diagram expansions. Journal of Statistical Physics, 8(3):231–237, July 1973.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012568.

Greenberg:1979:DBE

[GVZ79] William Greenberg, Jürgen Voigt, and P. F. Zweifel. Discretized Boltzmann equation: Lattice limit and non-Maxwellian gases. Journal of Statistical Physics, 21(6):649–657, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01107907.

Gunther:1975:MEP

[GZW75] L. Gunther and J. Zitkova-Wilcox. The mössbauer effect: A potentially ideal probe into Brownian motion. Journal of Statistical Physics, 12(3):205–215, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011967.

Hall:1975:SIB

[Hal75] Carol K. Hall. Scaling in the ideal Bose gas. Journal of Statistical Physics, 13(2):157–172, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01221376.

Hall:1976:WPT

[Hal76] George L. Hall. Weak ‘phase transitions’ in asymptotic properties of lattice sums. Journal of Statistical Physics, 14(6):521–524, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012850.

Harris:1970:HCS

[Har70] S. Harris. Hall coefficient in solution for a simple dynamical model. Journal of Statistical Physics, 2(4):379–385, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020444.

Hoye:1977:SYC

[HB77] J. S. Høye and L. Blum. Solution of the Yukawa closure of the Ornstein–Zernike equation. Journal of Statistical Physics, 16(5):399–413, May 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01013184.

Harrison:1978:SLP

[HBQ78] Ralph J. Harrison, George H. Bishop, Jr., and George D. Quinn. Spanning lengths of percolation clusters. Journal of Statistical Physics, 19(1):53–64, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020333.

Hobson:1973:CSK

[HC73] Arthur Hobson and Bin-Kang Cheng. A comparison of the Shannon and Kullback information measures. Journal of Statistical Physics, 7(4):301–310, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014906. See comments

[TR73, Bat74].

Hioe:1974:GTM

[HD74] F. T. Hioe and C. Domb. Group theory, Markov chains, and excluded volume effect in polymers. Journal of Statistical Physics, 10(3):183–203, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016176.

Hoyningen-Huene:1976:CBZ

[HH76] Paul Hoyningen-Huene. On the connection between Zwanzig’s classical projection operator formalism and coarse-graining. Journal of Statistical Physics, 15(6):517–518, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020805.

Heilmann:1979:PTL

[HH79] Ole J. Heilmann and Dale A. Huckaby. Phase transitions in lattice gas models for water. Journal of Statistical Physics, 20 (4):371–383, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011778.

Hill:1974:HTE

[Hil74] Robert Nyden Hill. High-temperature exchange third virial coefficient for hard spheres via an asymptotic method for path integrals. Journal of Statistical Physics, 11(3):207–230, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010218.

Hilhorst:1977:ELR

[Hil77] H. J. Hilhorst. Exact linear renormalization of a one-dimensional system with phase transition. Journal of Statistical Physics, 17 (6):413–427, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014348.

Hirsch:1975:NPG

[Hir75] Helmut Hirsch. Note on the paper ‘Gibbs vs. Shannon Entropies’ by Richard L. Liboff. Journal of Statistical Physics, 13(3): 273–275, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012842. See [Lib74].

Hale:1975:PFM

[HK75] Barbara N. Hale and Jerry Kiefer. A partition function model for nucleation on surfaces. Journal of Statistical Physics, 12(5):437– 444, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012887.

Hoshen:1979:PCD

[HKK79] J. Hoshen, P. Klymko, and R. Kopelman. Percolation and cluster distribution. III. Algorithms for the site-bond problem. Journal of Statistical Physics, 21(5):583–600, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011170.

Hoshen:1978:PCD

[HKM78] J. Hoshen, R. Kopelman, and E. M. Monberg. Percolation and cluster distribution. II. layers, variable-range interactions, and exciton cluster model. Journal of Statistical Physics, 19(3): 219–242, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011724.

Hynes:1975:NMR

[HKW75] James T. Hynes, Raymond Kapral, and Michael Weinberg. Nonlinear momentum relaxation of an impurity in a harmonic chain.

Journal of Statistical Physics, 13(5):427–449, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013118.

Hintermann:1978:ERP

[HKW78] A. Hintermann, H. Kunz, and F. Y. Wu. Exact results for the Potts model in two dimensions. Journal of Statistical Physics, 19 (6):623–632, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011773.

Heilmann:1979:LML

[HL79] Ole J. Heilmann and Elliott H. Lieb. Lattice models for liquid crystals. Journal of Statistical Physics, 20(6):679–693, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01009518.

Harris:1977:EIF

[HM77] S. Harris and J. L. Monroe. Eigentheory of the inhomogeneous Fokker–Planck equation. Journal of Statistical Physics, 17(5): 377–381, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014405.

Hauge:1973:FHB

[HML73] E. H. Hauge and A. Martin-Löf. Fluctuating hydrodynamics and Brownian motion. Journal of Statistical Physics, 7(3):259–281, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030307.

Horsthemke:1977:AEN

[HMMH77] W. Horsthemke, M. Malek-Mansour, and B. Hayez. An asymptotic expansion of the nonlinear master equation. Journal of Statistical Physics, 16(2):201–215, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01418752.

Heilmann:1974:CML

[HMP74] O. J. Heilmann, R. Moss, and E. L. Praestgaard. A cell model for liquids. II. Exact solution for the restricted cell model for hard, parallel squares. Journal of Statistical Physics, 10(6):449–470, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020401.

Hasegawa:1979:SLT

[HN79] H. Hasegawa and T. Nakagomi. Semiclassical laser theory in the stochastic and thermodynamic frameworks. Journal of Statistical Physics, 21(2):191–214, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008698.

Hobson:1972:IIP

[Hob72] Arthur Hobson. The interpretation of inductive probabilities. Journal of Statistical Physics, 6(2–3):189–193, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023687.

Hosokawa:1976:EMR

[Hos76] Iwao Hosokawa. Ensemble mechanics for the random-forced Navier–Stokes flow. Journal of Statistical Physics, 15(2):87–104, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012029.

Heilmann:1973:PTH

[HP73] Ole J. Heilmann and Eigil Praestgaard. Phase transition of hard hexagons on a triangular lattice. Journal of Statistical Physics, 9(1):23–44, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016795.

Huang:1971:TTB

[HR71a] C. Huang and H. Reiss. Transmission time and bandwidth in pulse code modulation (application of the statistical thermodynamic formalism). Journal of Statistical Physics, 3(3):261–289, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011384.

Huerta:1971:AEC

[HR71b] Manuel A. Huerta and Harry S. Robertson. Approach to equilibrium of coupled harmonic oscillator systems. II. Journal of Statistical Physics, 3(2):171–189, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019849.

Hudson:1976:NKE

[HR76] Suzanne Hudson and John Ross. Nonequilibrium kinetics: Exact and approximate solutions. Journal of Statistical Physics, 14(6):469–481, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012846.

Hamrick:1979:SGS

[HR79] Gary C. Hamrick and Charles Radin. The symmetry of ground states under perturbation. Journal of Statistical Physics, 21(5): 601–607, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011171.

Harrison:1971:MDT

[HS71] Herman W. Harrison and William C. Schieve. Molecular dynamics of two-dimensional gases with realistic potentials. Journal of Statistical Physics, 3(1):35–38, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012185.

Henderson:1978:EAF

[HS78] Douglas Henderson and William R. Smith. Exact analytical formulas for the distribution functions of charged hard spheres in the mean spherical approximation. Journal of Statistical Physics, 19(2):191–200, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012511.

Hashitsumae:1977:QME

[HSS77] Natsuki Hashitsumae, Fumiaki Shibata, and Mayumi Shingū. Quantal master equation valid for any time scale. Journal of Statistical Physics, 17(4):155–169, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01040099.

Hanggi:1978:NTE

[HTGT78] P. Hanggi, H. Thomas, H. Grabert, and P. Talkner. Note on time evolution of non-Markov processes. Journal of Statistical Physics, 18(2):155–159, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014306.

Hubert:1976:THS

[Hub76] D. Hubert. H-theorem for the hard-sphere gas. Journal of Statistical Physics, 15(1):73–83, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012810.

Huckaby:1977:PTL

[Huc77] Dale A. Huckaby. Phase transitions in lattice gases of hardcore molecules having two orientations. Journal of Statistical Physics, 17(5):371–375, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014404.

Hosokawa:1975:TRF

[HY75] I. Hosokawa and K. Yamamoto. Turbulence in the randomly forced, one-dimensional Burgers flow. Journal of Statistical Physics, 13(3):245–272, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01012841.

Hynes:1974:ICE

[Hyn74] James T. Hynes. Initial condition effects for a Brownian particle in a harmonic chain. Journal of Statistical Physics, 11(3): 257–275, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010220.

Isbister:1979:ADH

[IF79] D. J. Isbister and B. C. Freasier. Adsorption of dipolar hard spheres onto a smooth, hard wall in the presence of an electric field. Journal of Statistical Physics, 20(3):331–345, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011941.

Illner:1979:COG

[IK79] Reinhard Illner and Ivan Kuscer. Collision operators as generators of Markov processes and their spectra. Journal of Statistical Physics, 20(3):303–316, March 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011939.

Ivanov:1973:QMR

[Iva73a] E. N. Ivanov. The quasicrystal model of rotational Brownian motion. Journal of Statistical Physics, 8(2):183–188, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008539.

Ivanov:1973:RDC

[Iva73b] E. N. Ivanov. Rotational displacements in crystals. Journal of Statistical Physics, 8(2):177–181, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008538.

Jancovici:1977:PCF

[Jan77] B. Jancovici. Pair correlation function in a dense plasma and pycnonuclear reactions in stars. Journal of Statistical Physics, 17 (5):357–370, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014403.

Jhon:1979:MMC

[JD79] Myung S. Jhon and John S. Dahler. A mode-mode coupling theory of chemical reaction in a dense fluid. Journal of Statistical Physics, 20(1):3–18, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013743.

Jones:1978:DTL

[Jon78] Gerald L. Jones. Decimation transformations in lattice systems of continuous spins. Journal of Statistical Physics, 19(4): 417–425, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011759.

Kaniel:1977:DTE

[Kan77] Shmuel Kaniel. On the derivation of thermohydrodynamic equations from the Boltzmann equations. Journal of Statistical Physics, 16(5):415–450, May 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013185.

Katz:1970:CSP

[Kat70] Joseph L. Katz. The critical supersaturations predicted by nucleation theory. Journal of Statistical Physics, 2(2):137–146, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009735.

Kawasaki:1975:GRC

[Kaw75] Kyozi Kawasaki. Growth rate of critical nuclei near the critical point of a fluid. Journal of Statistical Physics, 12(5):365–374, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012883.

Kannman:1972:RIE

[KBR72] D. Kannman and A. T. Bharucha-Reid. Random integral equation formulation of a generalized Langevin equation. Journal of Statistical Physics, 5(3):209–233, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023743.

Kretschmer:1976:LNR

[KBS76] R. Kretschmer, K. Binder, and D. Stauffer. Linear and nonlinear relaxation and cluster dynamics near critical points. Journal of Statistical Physics, 15(4):267–297, October 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023054.

Keizer:1970:NAJa

[Kei70a] Joel E. Keizer. A new approach for the justification of ensembles in quantum statistical mechanics — I. Journal of Statistical Physics, 2(3):213–232, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01030743.

Keizer:1970:NAJb

[Kei70b] Joel E. Keizer. A new approach for the justification of ensembles in quantum statistical mechanics — II. Journal of Statistical Physics, 2(3):233–249, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01030744.

Keizer:1972:SSS

[Kei72] Joel Keizer. On the solutions and the steady states of a master equation. Journal of Statistical Physics, 6(2–3):67–72, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023679.

Keizer:1976:OEF

[Kei76] Joel Keizer. An observation on the effect of fluctuations far from equilibrium: The Lotka–Volterra model. Journal of Statistical Physics, 15(6):477–483, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020801.

Kenkre:1978:GME

[Ken78] V. M. Kenkre. Generalized master equations under delocalized initial conditions. Journal of Statistical Physics, 19(4):333–340, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011752.

Kurtze:1978:YLE

[KF78] Douglas A. Kurtze and Michael E. Fisher. The Yang–Lee edge singularity in spherical models. Journal of Statistical Physics, 19(3):205–218, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011723.

Kadanoff:1976:VARa

[KHY76a] Leo P. Kadanoff, Anthony Houghton, and Mehmet C. Yalabik. Variational approximations for renormalization group transformations. Journal of Statistical Physics, 14(2):171–203, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011765.

Kadanoff:1976:VARb

[KHY76b] Leo P. Kadanoff, Anthony Houghton, and Mehmet C. Yalabik. Variational approximations for renormalization group transformations. Journal of Statistical Physics, 15(3):263, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012881.

Kihara:1975:CEK

[Kih75] Taro Kihara. The Chapman–Enskog and Kihara approximations for isotopic thermal diffusion in gases. Journal of Statistical Physics, 13(2):137–143, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01221374.

Kikuchi:1971:TDC

[Kik71] Ryoichi Kikuchi. Treatment of droplike clusters in nucleation theory using Reiss’s method. Journal of Statistical Physics, 3 (3):331–345, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011388.

Kimball:1979:KIM

[Kim79] J. C. Kimball. The kinetic Ising model: Exact susceptibilities of two simple examples. Journal of Statistical Physics, 21(3): 289–300, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011471.

Kleban:1974:TMB

[Kle74] Peter Kleban. Toward a microscopic basis for the de Gennes narrowing. Journal of Statistical Physics, 11(4):317–322, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009791.

Kalos:1978:CMN

[KLPS78] M. Kalos, Joel L. Lebowitz, O. Penrose, and A. Sur. Clusters, metastability, and nucleation: Kinetics of first-order phase transitions. Journal of Statistical Physics, 18(1):39–52, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014669.

Kihara:1975:GTC

[KM75] Taro Kihara and Kazunori Miyoshi. Geometry of three convex bodies applicable to three-molecule clusters in polyatomic gases. Journal of Statistical Physics, 13(4):337–345, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012012.

Kitahara:1976:PIR

[KM76] Kazuo Kitahara and Horia Metiu. On the path integral representation of stochastic processes. Journal of Statistical Physics, 15(2):141–147, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012033.

Kubo:1973:FRM

[KMK73] Ryogo Kubo, Kazuhiro Matsuo, and Kazuo Kitahara. Fluctuation and relaxation of macrovariables. Journal of Statistical Physics, 9(1):51–96, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01016797.

Kenkre:1973:GME

[KMS73] V. M. Kenkre, E. W. Montroll, and M. F. Shlesinger. Generalized master equations for continuous-time random walks. Journal of Statistical Physics, 9(1):45–50, September 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01016796.

Knobloch:1979:SDL

[Kno79] Edgar Knobloch. On the statistical dynamics of the Lorenz model. Journal of Statistical Physics, 20(6):695–709, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009519.

Kollmann:1976:FIS

[Kol76] Wolfgang Kollmann. Functional integrals in the statistical theory of turbulence and the Burgers model equation. Journal of Statistical Physics, 14(3):291–303, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023630.

Kostin:1975:FDP

[Kos75] M. D. Kostin. Friction and dissipative phenomena in quantum mechanics. Journal of Statistical Physics, 12(2):145–151, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010029.

Kowalski:1977:SIF

[Kow77] Jacek M. Kowalski. Simple inequality for the free energy of disordered systems. Journal of Statistical Physics, 16(2):197–200, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01418751.

Kalos:1977:SLV

[KPR77] M. H. Kalos, J. K. Percus, and M. Rao. Structure of a liquid-vapor interface. Journal of Statistical Physics, 17(3): 111–136, September 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01106912.

Kremen:1975:TIN

[Kre75] Alexandr Kremen. Transmission of information in nucleotide pools: Application of the statistical thermodynamic formalism. Journal of Statistical Physics, 12(3):253–262, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011969.

Kometani:1975:SSO

[KS75] K. Kometani and H. Shimizu. A study of self-organizing processes of nonlinear stochastic variables. Journal of Statistical Physics, 13(6):473–490, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01013146.

Kerr:1978:CMF

[KS78a] Robert M. Kerr and Eric D. Siggia. Cascade model of fully developed turbulence. Journal of Statistical Physics, 19(5): 543–552, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011698.

Kunz:1978:ESP

[KS78b] Hervé Kunz and Bernard Souillard. Essential singularity in percolation problems and asymptotic behavior of cluster size distribution. Journal of Statistical Physics, 19(1):77–106, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020335.

Khanin:1979:EFE

[KS79] K. M. Khanin and Ya. G. Sinai. Existence of free energy for models with long-range random Hamiltonians. Journal of Statistical Physics, 20(6):573–584, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009511.

Kugler:1973:CMD

[Kug73] Alfred A. Kugler. Collective modes, damping, and the scattering function in classical liquids. Journal of Statistical Physics, 8(2):107–153, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008535.

Kugler:1975:TLF

[Kug75] A. A. Kugler. Theory of the local field correction in an electron gas. Journal of Statistical Physics, 12(1):35–87, January 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024183.

Knops:1977:PSR

[KvLH77] H. J. F. Knops, J. M. J. van Leeuwen, and P. C. Hemmer. Position-space renormalization for systems with weak long-range interactions and the breakdown of hyperscaling. Journal of Statistical Physics, 17(4):197–214, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01040102.

Lakatos:1970:STD

[Lak70] Katja Lakatos. On the statistics of a three-dimensional gas of long thin rods. Journal of Statistical Physics, 2(2):121–136, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009734.

Landauer:1973:ECS

[Lan73] R. Landauer. Entropy changes for steady-state fluctuations. Journal of Statistical Physics, 9(4):351–371, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012167.

Landauer:1974:ECS

[Lan74] R. Landauer. Entropy changes for steady-state fluctuations. Journal of Statistical Physics, 11(6):525, December 1974. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008894.

Landauer:1975:SEP

[Lan75] R. Landauer. Stability and entropy production in electrical circuits. Journal of Statistical Physics, 13(1):1–16, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012596.

Lavaud:1978:EGM

[Lav78] Michel Lavaud. Estimates of general Mayer graphs. II. Longrange behavior of graphs with two root points occurring in the theory of ionized systems. Journal of Statistical Physics, 19(5): 429–451, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011691.

Lebowitz:1976:PSM

[Leb76] Joel L. Lebowitz. Program of the 35th Statistical Mechanics Meeting. Journal of Statistical Physics, 14(6):537–539, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012853.

Lebowitz:1977:CPIb

[Leb77a] Joel L. Lebowitz. Coexistence of phases in Ising ferromagnets. Journal of Statistical Physics, 16(6):463–476, June 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01152284.

Lebowitz:1977:EDO

[Leb77b] Joel L. Lebowitz. On the equivalence of different order parameters for Ising ferromagnets. Journal of Statistical Physics, 16(1):3–9, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014602.

Lee:1972:DIA

[Lee72] Jon Lee. Direct-interaction approximation by the modalinteraction perturbation technique. Journal of Statistical Physics, 4(2–3):175–191, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008548.

Lee:1973:PPT

[Lee73] V. J. Lee. A particle picture of ‘tunneling’ and the nature of ‘photon’-matter interaction. Journal of Statistical Physics, 8 (2):189–195, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008540.

Lenard:1978:TPG

[Len78] A. Lenard. Thermodynamical proof of the Gibbs formula for elementary quantum systems. Journal of Statistical Physics, 19 (6):575–586, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011769.

Leribaux:1971:CES

[Ler71] Henri R. Leribaux. A contribution to the equation of state of fluids at low temperatures based on thermodynamic Green’s functions. Journal of Statistical Physics, 3(1):1–16, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012183.

Leung:1970:FDR

[Leu70] M. C. Leung. A fluctuation-dissipation relation and its applications to thermodynamic fluctuations in superconducting cylinders. Journal of Statistical Physics, 2(4):305–322, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020440.

Levitt:1973:ODT

[Lev73] David G. Levitt. One-dimensional time-dependent distributions. Journal of Statistical Physics, 7(4):329–336, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014908.

Liboff:1974:GVS

[Lib74] Richard L. Liboff. Gibbs vs. Shannon entropies. Journal of Statistical Physics, 11(4):343–357, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009793. See note

[Hir75].

Lindblad:1974:MIT

[Lin74a] Göran Lindblad. Measurements and information for thermodynamic quantities. Journal of Statistical Physics, 11(3):231– 255, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010219.

Lindenberg:1974:RWN

[Lin74b] Katja Lindenberg. Random walks with nonnearest-neighbor transitions. II. Analytic one-dimensional theory for exponentially distributed steps in systems with boundaries. Journal of Statistical Physics, 10(6):485–503, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020403.

Lebenhaft:1979:DCP

[LK79] Julian R. Lebenhaft and Raymond Kapral. Diffusion-controlled processes among partially absorbing stationary sinks. Journal of Statistical Physics, 20(1):25–56, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013745.

Lanford:1977:TEI

[LLL77] Oscar E. Lanford III, Joel L. Lebowitz, and Elliott H. Lieb. Time evolution of infinite anharmonic systems. Journal of Statistical Physics, 16(6):453–461, June 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01152283.

Lee:1973:TCI

[LMpY73] F. T. Lee, E. W. Montroll, and Lee po Yu. Two-component Ising chain with nearest-neighbor interaction. Journal of Statistical Physics, 8(4):309–333, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020270.

Lieb:1975:TLJ

[LN75] Elliott H. Lieb and Heide Narnhofer. The thermodynamic limit for jellium. Journal of Statistical Physics, 12(4):291–310, April 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012066.

Lieb:1976:TLJ

[LN76] Elliott H. Lieb and Heide Narnhofer. The thermodynamic limit for jellium. Journal of Statistical Physics, 14(5):465, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01040705.

Lovett:1972:FML

[Lov72] Ronald Lovett. Frequency modulation of a laser by fluctuations in a chemical reaction. Journal of Statistical Physics, 4(1):43–54, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008470.

Lovesey:1978:SPM

[Lov78] S. W. Lovesey. Scaling properties of models of nonequilibrium phenomena. Journal of Statistical Physics, 19(4):367–390, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011756.

Liebman:1974:ECO

[LP74] Alan J. Liebman and Jean-Yves Parlange. Exact calculation of one-dimensional irreducible integrals. Journal of Statistical Physics, 11(1):71–81, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019478.

Lebowitz:1977:CPIa

[LP77a] J. L. Lebowitz and O. Penrose. Cluster and percolation inequalities for lattice systems with interactions. Journal of Statistical Physics, 16(4):321–337, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020425.

Lundgren:1977:SMT

[LP77b] T. S. Lundgren and Y. B. Pointin. Statistical mechanics of two-dimensional vortices. Journal of Statistical Physics, 17(5): 323–355, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014402.

Lebowitz:1972:VAF

[LS72] J. L. Lebowitz and J. Sykes. The velocity autocorrelation function of a finite model system. Journal of Statistical Physics, 6(2– 3):157–171, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023684.

Lebowitz:1978:TPL

[LS78] Joel L. Lebowitz and Herbert Spohn. Transport properties of the Lorentz gas: Fourier’s law. Journal of Statistical Physics, 19(6): 633–654, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011774.

Lindenberg:1975:FPT

[LSFL75] K. Lindenberg, K. E. Shuler, J. Freeman, and T. J. Lie. First passage time and extremum properties of Markov and independent processes. Journal of Statistical Physics, 12(3):217–251, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011968.

Lumsden:1979:SMC

[LT79] Charles J. Lumsden and L. E. H. Trainor. On the statistical mechanics of constrained biophysical systems. Journal of Statistical Physics, 20(6):657–669, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009516.

Lucke:1976:SDL

[Lüc76] M. Lücke. Statistical dynamics of the Lorenz model. Journal of Statistical Physics, 15(6):455–475, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020800.

Luttinger:1976:SLD

[Lut76] J. M. Luttinger. Scattering length and diffusion. Journal of Statistical Physics, 15(3):215–218, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012877.

Meirovitch:1976:ZFM

[MA76] Hagai Meirovitch and Z. Alexandrowicz. On the zero fluctuation of the ‘microscopic free energy’ and its potential use. Journal of Statistical Physics, 15(2):123–127, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012031.

Meirovitch:1977:SMM

[MA77] Hagai Meirovitch and Z. Alexandrowicz. The stochastic models method applied to the critical behavior of Ising lattices. Journal of Statistical Physics, 16(2):121–138, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01418747.

Maheshwari:1975:FIR

[Mah75] Amar Maheshwari. Functional integral representations of partition function without limiting procedure. Techniques of calculation of moments. Journal of Statistical Physics, 12(1):11–20, January 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01024181.

Mandell:1976:PPF

[Man76] M. J. Mandell. On the properties of a periodic fluid. Journal of Statistical Physics, 15(4):299–305, October 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023055.

Mandelbrot:1977:FFC

[Man77] Benoît B. Mandelbrot. Fractals: form, chance, and dimension. W. H. Freeman and Company, New York, NY, USA, revised edition, 1977. ISBN 0-7167-0473-0, 0-7167-0474-9 (paperback). xvi + 365 pp. LCCN QA447 .M3613 1977a. Translated from the French.

Martin:1977:SDI

[Mar77] Ph. A. Martin. On the stochastic dynamics of Ising models. Journal of Statistical Physics, 16(2):149–168, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01418749.

Matsuno:1974:PKM

[Mat74] Koichiro Matsuno. Probabilistic kinetics of macroprocesses in broken microscopic reversibility. Journal of Statistical Physics, 11(2):87–132, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012514.

Matsuno:1975:PKM

[Mat75] Koichiro Matsuno. Probabilistic kinetics of macroprocesses in broken microscopic reversibility. Journal of Statistical Physics, 13(1):101–102, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012602.

Matsuo:1977:RMA

[Mat77] Kazuhiro Matsuo. Relaxation mode analysis of nonlinear birth and death processes. Journal of Statistical Physics, 16(2): 169–195, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01418750.

Matsuo:1978:BMN

[Mat78] Kazuhiro Matsuo. Brownian motion of nonlinear oscillators: The projection operator method and its application to a double-well-potential oscillator. Journal of Statistical Physics, 18(6):535– 555, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014478.

Mayer:1978:RBR

[May78] J. E. Mayer. Response to book review: Statistical mechanics, 2nd ed. 1. Journal of Statistical Physics, 18(5):529–531, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014522.

Mazo:1975:TBM

[Maz75] Robert M. Mazo. On the theory of Brownian motion. VIII. A Brownian particle in a dilute gas with inelastic collisions. Journal of Statistical Physics, 12(5):427–436, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012886.

Morton:1970:CEE

[MC70] J. B. Morton and S. Corrsin. Consolidated expansions for estimating the response of a randomly driven nonlinear oscillator. Journal of Statistical Physics, 2(2):153–194, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01009737.

Mandeville:1974:IBI

[MC74] George Mandeville and Michael Coopersmith. Irreversible behavior of interacting systems. II. Fluctuations in equilibrium. Journal of Statistical Physics, 10(5):405–420, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008802.

McLaughlin:1976:SBL

[McL76] John McLaughlin. Successive bifurcations leading to stochastic behavior. Journal of Statistical Physics, 15(4):307–326, October 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023056.

McLaughlin:1978:CBD

[McL78] John McLaughlin. Connection between dissipative and resonant conservative nonlinear oscillators. Journal of Statistical Physics, 19(6):587–591, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011770.

Mosteller:1973:CPO

[MD73] R. D. Mosteller and J. J. Duderstadt. A comparison of projection operator formalisms for the study of self-diffusion. Journal of Statistical Physics, 9(3):197–213, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008728.

Mosteller:1974:SRM

[MD74] R. D. Mosteller and J. J. Duderstadt. Spectral representations for the memory kernel characterizing self-diffusion. Journal of Statistical Physics, 11(5):409–420, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026732.

McKean:1979:BRB

[MD79] H. P. McKean and James W. Dufty. Book review: Fractals: Form, chance, and dimension. Journal of Statistical Physics, 20(6):725–729, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009522.

Meecham:1973:SST

[Mee73] W. C. Meecham. On the simple-source theory of sound from statistical turbulence. Journal of Statistical Physics, 8(2):197– 212, June 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008541.

Merlini:1979:BCC

[Mer79] D. Merlini. Boundary conditions and cluster property in twodimensional Ising ferromagnets. Journal of Statistical Physics, 21 (6):739–745, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01107912.

Madden:1975:STT

[MK75] Paul Madden and Daniel Kivelson. Statistical transport theory of reacting two-component fluid: Light scattering from an A ⇀ ↽ B liquid. Journal of Statistical Physics, 12(3):167–204, March 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011966.

Muller-Krumbhaar:1973:DPM

[MKB73] H. Müller-Krumbhaar and K. Binder. Dynamic properties of the Monte Carlo method in statistical mechanics. Journal of Statistical Physics, 8(1):1–24, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008440.

Millard:1972:CME

[ML72] Kenneth Millard and Harvey S. Leff. A constant-magnetization ensemble for the classical anisotropic Heisenberg model. Journal of Statistical Physics, 6(2–3):133–155, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01023683.

Millard:1973:MSV

[ML73] Kenneth Millard and Louis H. Lund. Metastable states in the van der Waals–Maxwell theory. Journal of Statistical Physics, 8(3):225–230, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012567.

Millard:1974:ESC

[ML74] Kenneth Millard and Harvey S. Leff. An exact solution to the classical, anisotropic Heisenberg model with long-range Kac interactions. Journal of Statistical Physics, 10(3):205–235, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01016177.

Montroll:1977:SMS

[ML77] E. W. Montroll and Uzi Landman, editors. Statistical mechanics and statistical methods in theory and application: a tribute to Elliott W. Montroll: [proceedings of a Symposium on Statistical Mechanics and Statistical Methods in Theory and Application, held at the University of Rochester, Rochester, New York, November 4, 1976]. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-31077-5. LCCN QC174.7 .S94 1976.

Martin-Lof:1979:EEG

[ML79] Anders Martin-Löf. The equivalence of ensembles and the Gibbs phase rule for classical lattice systems. Journal of Statistical Physics, 20(5):557–569, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012899.

Matsuo:1978:STN

[MLS78] K. Matsuo, K. Lindenberg, and K. E. Shuler. Stochastic theory of nonlinear rate processes with multiple stationary states. II. Relaxation time from a metastable state. Journal of Statistical Physics, 19(1):65–75, July 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020334.

Maltz:1971:ECV

[MM71] F. H. Maltz and W. C. Meecham. The effect of curvature variation on the scattering from rough surfaces. Journal of Statistical Physics, 3(4):385–394, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008276.

Mashiyama:1978:OLL

[MM78] Kazuko T. Mashiyama and Hazime Mori. Origin of the Landau– Lifshitz hydrodynamic fluctuations in nonequilibrium systems and a new method for reducing the Boltzmann equation. Journal of Statistical Physics, 18(4):385–407, April 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01106730.

Malyshev:1979:ISC

[MM79] V. A. Malyshev and R. A. Minlos. Invariant subspaces of clustering operators. Journal of Statistical Physics, 21(3):231– 242, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011467.

Malek-Mansour:1975:MED

[MMN75] M. Malek-Mansour and G. Nicolis. A master equation description of local fluctuations. Journal of Statistical Physics, 13(3): 197–217, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012838.

Messager:1977:CFB

[MMS77] A. Messager and S. Miracle-Sole. Correlation functions and boundary conditions in the Ising ferromagnet. Journal of Statistical Physics, 17(4):245–262, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01040105.

Mitchell:1977:SMC

[MMSG77] D. J. Mitchell, D. A. McQuarrie, Attila Szabo, and J. Groeneveld. On the second-moment condition of stillinger and lovett. Journal of Statistical Physics, 17(1):15–20, July 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01089374.

Morita:1978:SMD

[MMT78] Terumitsu Morita, Hazime Mori, and Michio Tokuyama. A scaling method for deriving kinetic equations from the BBGKY hierarchy. Journal of Statistical Physics, 18(2):137–153, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014305.

Mou:1978:SCN

[MNM78] C. Y. Mou, G. Nicolis, and R. M. Mazo. Some comments on nonequilibrium phase transitions in chemical systems. Journal of Statistical Physics, 18(1):19–38, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014668.

Monchick:1973:AMB

[Mon73] L. Monchick. Asymptotic magnetic behavior of gas transport properties. Journal of Statistical Physics, 9(3):221–234, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008730.

Moussouris:1974:GMR

[Mou74] John Moussouris. Gibbs and Markov random systems with constraints. Journal of Statistical Physics, 10(1):11–33, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011714.

Monroe:1979:CIV

[MP79] James L. Monroe and Paul A. Pearce. Correlation inequalities for vector spin models. Journal of Statistical Physics, 21(6): 615–633, December 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01107905.

Marchioro:1978:VPI

[MPPT78] C. Marchioro, A. Pellegrinotti, M. Pulvirenti, and L. Triolo. Velocity of a perturbation in infinite lattice systems. Journal of Statistical Physics, 19(5):499–510, November 1978.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01011695. See erratum [MPPT80].

Marchioro:1980:EVP

[MPPT80] C. Marchioro, A. Pellegrinotti, M. Pulvirenti, and L. Triolo. Erratum: Velocity of a perturbation in infinite lattice systems. Journal of Statistical Physics, 22(6):743, June 1980. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01013939. See [MPPT78].

Mishima:1979:TIP

[MPS79] Nobuhiko Mishima, Tomio Yamakoshi Petrosky, and Ryoji Suzuki. Theory of irreversible processes in nonequilibrium quantum systems. I. General formalism. Journal of Statistical Physics, 21(4):347–392, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009607.

Mishima:1976:QSH

[MPY76] Nobuhiko Mishima, Tomio Yamakoshi Petrosky, and Miwae Yamazaki. Quantum statistical hierarchy equation in nonequilibrium systems. Journal of Statistical Physics, 14(4):359–380, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030200.

Mandell:1975:SPT

[MR75a] M. J. Mandell and H. Reiss. Scaled particle theory: Solution to the complete set of scaled particle theory conditions: Applications to surface structure and dilute mixtures. Journal of Statistical Physics, 13(2):113–128, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01221372.

Mandell:1975:TCB

[MR75b] M. J. Mandell and H. Reiss. Thermodynamics of curved boundary layers. Journal of Statistical Physics, 13(2):107–112, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01221371.

Mandell:1976:SPT

[MR76] M. J. Mandell and H. Reiss. Scaled particle theory: Solution to the complete set of scaled particle theory conditions: Applications to surface structure and dilute mixtures. Journal of Statistical Physics, 14(6):535–536, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012852.

McGraw:1979:NNC

[MR79] Robert McGraw and Howard Reiss. Nucleation near a critical temperature. Journal of Statistical Physics, 20(4):385–413, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011779.

Mrygon:1974:IDF

[Mry74] B. Mrygoń. Influence of density fluctuations on the equation of state. Journal of Statistical Physics, 10(6):471–484, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020402.

Montroll:1973:RWL

[MS73] Elliott W. Montroll and Harvey Scher. Random walks on lattices. IV. Continuous-time walks and influence of absorbing boundaries. Journal of Statistical Physics, 9(2):101–135, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01016843.

Monroe:1974:GIA

[MS74] J. L. Monroe and A. J. F. Siegert. GKS inequalities for arbitrary spin Ising ferromagnets. Journal of Statistical Physics, 10 (3):237–244, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016178.

McCraw:1978:MTD

[MS78] R. J. McCraw and L. S. Schulman. Metastability in the twodimensional Ising model. Journal of Statistical Physics, 18 (3):293–308, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01018095.

McCourt:1973:KTV

[MT73] F. R. McCourt and A. Tip. Kinetic theory of vibrational relaxation in a radiation field: The optic-acoustic effect. Journal of Statistical Physics, 8(1):57–78, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01008442.

Milton:1973:SMM

[MTH73] J. Susan Milton, Chris P. Tsokos, and S. T. Hardiman. A stochastic model for metabolizing systems with computer simulation. Journal of Statistical Physics, 8(1):79–101, May 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008443.

McNeil:1974:NPT

[MW74] K. J. McNeil and D. F. Walls. Nonequilibrium phase transitions in chemical reactions. Journal of Statistical Physics, 10(6):439– 448, June 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020400.

Montroll:1975:SWI

[MW75] Elliott W. Montroll and Bruce J. West. Scattering of waves by irregularities in periodic discrete lattice spaces. I. Reduction of problem to quadratures on a discrete model of the Schrödinger equation. Journal of Statistical Physics, 13(1):17–42, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012597.

Marchand:1977:SIE

[MW77] Jean-Paul Marchand and Walter Wyss. Statistical inference and entropy. Journal of Statistical Physics, 16(4):349–355, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01020427.

Matheson:1975:SMF

[MWG75] I. Matheson, D. F. Walls, and C. W. Gardiner. Stochastic models of first-order nonequilibrium phase transitions in chemical reactions. Journal of Statistical Physics, 12(1):21–34, January 1975.

CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024182.

Myerscough:1972:NSMb

[Mye72a] C. J. Myerscough. Nonequilibrium statistical mechanics of finite classical systems — II. Journal of Statistical Physics, 5(1–2):59– 82, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008371.

Myerscough:1972:NSMa

[Mye72b] C. J. Myerscough. Nonequilibrium statistical mechanics of finite classical systems-i. Journal of Statistical Physics, 5(1–2):35–58, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008370.

Nenciu:1973:SCG

[Nen73] G. Nenciu. On the surface contribution to the grand-canonical pressure of free quantum gases. Journal of Statistical Physics, 7 (2):119–130, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01024210.

Neumann:1975:LMR

[Neu75] H. P. Neumann. A lattice model for a real substance. Journal of Statistical Physics, 13(2):173–183, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01221377.

Newman:1976:RRG

[New76] Charles M. Newman. Rigorous results for general Ising ferromagnets. Journal of Statistical Physics, 15(5):399–406, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020342.

Northrup:1978:CTR

[NH78] Scott H. Northrup and James T. Hynes. Coupling of translational and reactive dynamics for a simple lattice model. Journal of Statistical Physics, 18(1):91–105, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014672.

Nicolis:1972:FAN

[Nic72] G. Nicolis. Fluctuations around nonequilibrium states in open nonlinear systems. Journal of Statistical Physics, 6(2–3):195– 222, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023688.

Nishigori:1976:TKE

[Nis76] Takeo Nishigori. Theory of kinetic equations for systems with production and absorption of particles. Journal of Statistical Physics, 14(3):249–262, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023627.

Nishigori:1979:MFS

[Nis79] T. Nishigori. On the memory functions in simple classical liquids. Journal of Statistical Physics, 20(1):83–93, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01013747.

Niwa:1978:TCF

[Niw78] Toshio Niwa. Time correlation functions of a one-dimensional infinite system. Journal of Statistical Physics, 18(3):309–317, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01018096.

Noyola:1978:CBC

[NM78] Magdaleno Medina Noyola and Donald A. McQuarrie. Critical behavior of the correlation function of a Coulombic fluid. Journal of Statistical Physics, 18(5):445–450, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014517.

Nicolis:1976:OIN

[NMMVK76] G. Nicolis, M. Malek-Mansour, A. Van Nypelseer, and K. Kitahara. The onset of instabilities in nonequilibrium systems. Journal of Statistical Physics, 14(5):417–432, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01040702.

Naitoh:1974:NKI

[NN74] Toyoaki Naitoh and Katsuhiko Nagai. Nonexistence of Kirkwood’s instability in a uniform fluid. Journal of Statistical Physics, 11(5):391–395, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01026730.

Nordholm:1973:NAS

[Nor73] K. Sture J. Nordholm. A new analytic solution for the Zwanzig– Lauritzen model of polymer chain folding. Journal of Statistical Physics, 9(3):235–264, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008731.

Naudts:1979:LRR

[NPV79] J. Naudts, J. V. Pulè, and A. Verbeure. Linear response and relaxation in quantum lattice systems. Journal of Statistical Physics, 21(3):279–288, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01011470.

Ninham:1970:MTT

[NPW70] Barry W. Ninham, V. Adrian Parsegian, and George H. Weiss. On the macroscopic theory of temperature-dependent van der Waals forces. Journal of Statistical Physics, 2(4):323–328, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01020441.

Nitzan:1974:CFA

[NR74] A. Nitzan and J. Ross. A comment on fluctuations around nonequilibrium steady states. Journal of Statistical Physics, 10(5):379–390, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008800.

Nossal:1974:GPR

[NW74] Ralph J. Nossal and George H. Weiss. A generalized Pearson random walk allowing for bias. Journal of Statistical Physics, 10(3):245–253, March 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01016179.

Nordholm:1975:SDE

[NZ75] Sture Nordholm and Robert Zwanzig. A systematic derivation of exact generalized Brownian motion theory. Journal of Statistical Physics, 13(4):347–371, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012013.

Ogunlana:1972:RQM

[Ogu72] J. Ogunlana. On reformulating quantum mechanics and stochastic theory. Journal of Statistical Physics, 4(2–3):217–225, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008551.

Onuki:1978:LEM

[Onu78a] Akira Onuki. Langevin equation with multi-poissonian noise. Journal of Statistical Physics, 19(4):325–332, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011751.

Onuki:1978:FS

[Onu78b] Akira Onuki. On fluctuations in μ-space. Journal of Statistical Physics, 18(5):475–499, May 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014519.

Oonuki:1972:TPF

[Oon72] Makoto Oonuki. Time partition function analysis of a neural network. Journal of Statistical Physics, 6(2–3):119–131, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023682.

Patera:1972:IGI

[Pat72] J. Patera. Incompatibility of gauge invariance and nonrelativistic locality in path integral formulation. Journal of Statistical Physics, 5(3):185–188, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01023741.

Piasecki:1976:GET

[PC76] J. Piasecki and B. Cichocki. Generalized Enskog theory for homogeneous systems. Journal of Statistical Physics, 14(5):433– 459, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01040703.

Peak:1977:ESC

[PC77] David Peak and James W. Corbett. The effect of spatial correlations on steady-state nucleation kinetics in dense fluid systems. Journal of Statistical Physics, 17(3):97–110, September 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01106911.

Papadopoulos:1978:PIT

[PD78] George J. Papadopoulos and J. T. (Jozef T.) Devreese, editors. Path integrals and their applications in quantum, statistical, and solid state physics, volume 34 of NATO advanced study institutes series: Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1978. ISBN 0-306-40017-0. LCCN QC174.17.P27 N18 1977.

Percus:1976:ESC

[Per76a] J. K. Percus. Equilibrium state of a classical fluid of hard rods in an external field. Journal of Statistical Physics, 15(6): 505–511, December 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020803.

Percus:1976:MDV

[Per76b] J. K. Percus. Model for density variation at a fluid surface. Journal of Statistical Physics, 15(5):423–435, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020344.

Percus:1977:ODI

[Per77] J. K. Percus. One-dimensional Ising model in arbitrary external field. Journal of Statistical Physics, 16(3):299–309, March 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020384.

Petrosky:1979:TIP

[Pet79] Tomio Yamakoshi Petrosky. Theory of irreversible processes in nonequilibrium quantum systems. II. Cluster expansions with contraction for homogeneous systems. Journal of Statistical Physics, 21(4):393–413, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009608.

Pfaffelhuber:1977:ITS

[Pfa77] E. Pfaffelhuber. Information-theoretic stability and evolution criteria in irreversible thermodynamics. Journal of Statistical Physics, 16(1):69–90, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014606.

Pierce:1972:SPP

[Pie72] William H. Pierce. Statistically periodic processes and Toeplitz_{(αbyβ)} matrices. Journal of Statistical Physics, 5(3):189– 207, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023742.

Pina:1974:MTK

[Piñ74] E. Piña. On a modified Thorne kinetic equation for a binary mixture of rigid spheres. Journal of Statistical Physics, 11(5): 433–436, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026734.

Pinsky:1975:SSL

[Pin75] Mark A. Pinsky. Stochastic solution of the linearized Boltzmann equation. Journal of Statistical Physics, 13(3):189–196, September 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012837.

Portnow:1976:SFC

[PK76] J. Portnow and K. Kitahara. Stability and fluctuation in chemical systems. Journal of Statistical Physics, 14(6):501–508, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012848.

Penrose:1971:RTM

[PL71] O. Penrose and J. L. Lebowitz. Rigorous treatment of metastable states in the van der Waals–Maxwell theory. Journal of Statistical Physics, 3(2):211–236, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019851.

Plischke:1974:FSB

[Pli74] Michael Plischke. Ferromagnetism in the single-band Hubbard model: An exact high-temperature expansion. Journal of Statistical Physics, 11(2):159–167, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012517.

Penrose:1978:GCF

[PLM^{+}78] O. Penrose, Joel L. Lebowitz, J. Marro, M. H. Kalos, and A. Sur. Growth of clusters in a first-order phase transition. Journal of Statistical Physics, 19(3):243–267, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011725.

Procaccia:1978:NBS

[PM78] Itamar Procaccia and Shaul Mukamel. A note on the bimodal stationary distributions in hard chemical instabilities. Journal of Statistical Physics, 18(6):633–635, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01014482.

Poland:1972:ERR

[Pol72] Douglas Poland. Explicit relations for the radial distribution functions for one-dimensional lattice (arbitrary spacing) fluids and solutions. Journal of Statistical Physics, 5(3):159–171, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01023739.

Presutti:1975:MDT

[Pre75] E. Presutti. A mechanical definition of the thermodynamic pressure. Journal of Statistical Physics, 13(4):301–314, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012009.

Pearce:1977:SLV

[PT77] Paul A. Pearce and Colin J. Thompson. The spherical limit for nvector correlations. Journal of Statistical Physics, 17(4):189–196, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01040101.

Piasecki:1979:LTB

[PW79] J. Piasecki and E. Wajnryb. Long-time behavior of the Lorentz electron gas in a constant, uniform electric field. Journal of Statistical Physics, 21(5):549–559, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011167.

Rainwater:1978:DFE

[Rai78] James C. Rainwater. Delta-function expansion of Mayer function with application to virial coefficients. Journal of Statistical Physics, 19(2):177–189, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012510.

Ray:1979:CSN

[Ray79] Dipankar Ray. On a class of solutions of nonlinear Boltzmann equations. Journal of Statistical Physics, 20(1):115–119, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013749.

Rae:1971:SMT

[RD71] J. Rae and R. Davidson. Statistical mechanical transformation theory for quantum systems with discrete spectra. Journal of Statistical Physics, 3(2):135–147, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01019847.

Rappaport:1974:PMT

[RD74] David Rappaport and Joshua Dayan. A probabilistic model for tracer distribution in multiphase spatially inhomogeneous transport systems. Journal of Statistical Physics, 11(6):481–501, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008891.

Reddy:1975:ESS

[Red75] V. T. N. Reddy. On the existence of the steady state in the stochastic Volterra–Lotka model. Journal of Statistical Physics, 13(1):61–64, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012599.

Reiss:1970:Pa

[Rei70a] H. Reiss. Prospectus. Journal of Statistical Physics, 2(1):113– 117, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009715.

Reiss:1970:Pb

[Rei70b] H. Reiss. Prospectus. Journal of Statistical Physics, 2(2):205– 209, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009742.

Reiss:1970:Pc

[Rei70c] H. Reiss. Prospectus. Journal of Statistical Physics, 2(3):293– 297, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01030749.

Reiss:1970:Pd

[Rei70d] H. Reiss. Prospectus. Journal of Statistical Physics, 2(4):391– 395, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020447.

Reiss:1970:TDC

[Rei70e] H. Reiss. Treatment of droplike clusters by means of the classical phase integral in nucleation theory. Journal of Statistical Physics, 2(1):83–104, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009712.

Reiss:1971:TPH

[Rei71] H. Reiss. Thermoelastic properties of helical protein molecules. Journal of Statistical Physics, 3(1):69–86, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012188.

Reiss:1972:SAV

[Rei72] H. Reiss. Superposition approximations from a variation principle. Journal of Statistical Physics, 6(1):39–47, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01060200.

Resibois:1970:LHM

[Rés70] P. Résibois. On linearized hydrodynamic modes in statistical physics. Journal of Statistical Physics, 2(1):21–51, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009709.

Resibois:1975:AKTb

[Rés75] P. Résibois. Approximate kinetic theory of hard-sphere fluids near equilibrium: II. A quasihydrodynamic approximation for the velocity autocorrelation function. Journal of Statistical Physics, 13(5):393–426, November 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01013117.

Resibois:1978:TMN

[Rés78] P. Résibois. H-theorem for the (modified) nonlinear Enskog equation. Journal of Statistical Physics, 19(6):593–609, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011771.

Rudd:1972:ICS

[RFB72] W. G. Rudd, H. L. Frisch, and Louis Brickman. Inversion of the classical second virial coefficient. Journal of Statistical Physics, 5(1–2):133–135, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008376.

Reiss:1971:STF

[RH71] H. Reiss and C. Huang. Statistical thermodynamic formalism in the solution of information theory problems. Journal of Statistical Physics, 3(2):191–210, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01019850.

Runnels:1972:AYL

[RH72] L. K. Runnels and J. B. Hubbard. Applications of the Yang–Lee– Ruelle theory to hard-core lattice gases. Journal of Statistical Physics, 6(1):1–20, September 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01060198.

Richardson:1974:IDC

[Ric74] J. M. Richardson. The initial distribution in classical statistical mechanics. Journal of Statistical Physics, 11(4):323–341, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009792.

Ridgeway:1977:EPI

[Rid77] Don Ridgeway. Evaluation of phase integrals for Volterra systems of arbitrary numbers of interacting species. Journal of Statistical Physics, 16(1):59–68, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014605.

Resibois:1975:AKTa

[RL75] P. Résibois and Joel L. Lebowitz. Approximate kinetic theory of hard-sphere fluids near equilibrium. I. Formal theory. Journal of Statistical Physics, 12(6):483–516, June 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01025835.

Runnels:1976:AHC

[RL76] L. K. Runnels and J. L. Lebowitz. Analyticity of a hard-core multicomponent lattice gas. Journal of Statistical Physics, 14 (6):525–533, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012851.

Resibois:1977:AKT

[RL77] P. Résibois and J. L. Lebowitz. Approximate kinetic theory of hard-sphere fluids near equilibrium. Journal of Statistical Physics, 16(4):395–396, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020430.

Reich:1978:HDP

[RL78] G. R. Reich and P. L. Leath. High-density percolation: Exact solution on a Bethe lattice. Journal of Statistical Physics, 19(6): 611–622, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011772.

Raeside:1973:CLF

[RO73] D. E. Raeside and R. J. Owen. A class of loss functions of catenary form. Journal of Statistical Physics, 7(3):189–195, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01030302.

Roepstorff:1978:BBC

[Roe78] G. Roepstorff. Bounds for a Bose condensate in dimensions v ≥ 3. Journal of Statistical Physics, 18(2):191–206, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014310.

Rose:1979:RKT

[Ros79] Harvey A. Rose. Renormalized kinetic theory of nonequilibrium many-particle classical systems. Journal of Statistical Physics, 20(4):415–447, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011780.

Rousseau:1973:EGG

[Rou73] Martine Rousseau. Entropy of generalized Gaussian optical fields. Journal of Statistical Physics, 8(4):341–352, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020272.

Rowlinson:1979:TJD

[Row79] J. S. Rowlinson. Translation of J. D. van der Waals’ The thermodynamic theory of capillarity under the hypothesis of a continuous variation of density. Journal of Statistical Physics, 20 (2):197–200, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011513.

Rechtman:1978:CTD

[RP78] R. Rechtman and O. Penrose. Continuity of the temperature and derivation of the Gibbs canonical distribution in classical statistical mechanics. Journal of Statistical Physics, 19(4):359–366, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011755.

Revzen:1973:FIB

[RS73] M. Revzen and L. S. Schulman. Functional integration for Bose fields. Journal of Statistical Physics, 8(3):217–219, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012565.

Reissner:1975:CAC

[RS75] John E. Reissner and William A. Steele. A collisional approach to the calculation of time correlation functions. Transport coefficients of gases. Journal of Statistical Physics, 13(4):283–300, October 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012008.

Raveche:1977:UBH

[RS77] Harold J. Raveché and Charles A. Stuart. Uniqueness for the BBGKY hierarchy for hard spheres in one dimension. Journal of Statistical Physics, 17(5):311–321, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014401.

Runnels:1973:AIS

[Run73] L. K. Runnels. Adsorbate interactions and surface excess entropy. Journal of Statistical Physics, 8(4):359–366, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020274.

Runnels:1976:DRV

[Run76] L. K. Runnels. Duality relation for 32-vertex model on the triangular lattice. Journal of Statistical Physics, 14(1):39–47, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020133.

Romerio:1977:SDM

[RW77] M. Romerio and W. F. Wreszinski. On the spectrum of the dynamical matrix for a class of disordered harmonic systems. Journal of Statistical Physics, 17(5):301–310, November 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014400.

Robledo:1978:MTR

[RW78] Alberto Robledo and Luis Woodhouse. Multiple trapping of random walkers on periodic lattices. Journal of Statistical Physics, 19(2):129–147, August 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012507.

Romerio:1979:LST

[RW79] M. Romerio and W. Wreszinski. On the Lifschitz singularity and the tailing in the density of states for random lattice systems. Journal of Statistical Physics, 21(2):169–179, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008696.

Sanchez:1972:VGT

[San72] Isaac C. Sanchez. On the validity of a global thermodynamic inequality. Journal of Statistical Physics, 6(2–3):225–226, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023690.

Sastri:1975:LRC

[Sas75] Chelluri C. A. Sastri. Long-range correlations in kinetic theory. Journal of Statistical Physics, 13(1):43–60, July 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012598.

Sato:1970:BR

[Sat70] Hiroshi Sato. Book review. Journal of Statistical Physics, 2 (1):105–106, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009713.

Scheuer:1972:BR

[Sch72] Ernest M. Scheuer. Book review. Journal of Statistical Physics, 5(1–2):137–139, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008377.

Schmeidler:1973:EPN

[Sch73] David Schmeidler. Equilibrium points of nonatomic games. Journal of Statistical Physics, 7(4):295–300, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01014905.

Schlup:1974:SEV

[Sch74] W. A. Schlup. Some evidence for the validity of the noisetemperature inequality θ ≥ T in the relaxation approximation of the one-dimensional electron transport problem in high electric fields. Journal of Statistical Physics, 11(5):421–431, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01026733.

Schilder:1975:DOM

[Sch75a] Michael Schilder. Dynamic observation models. Journal of Statistical Physics, 13(6):491–514, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01013147.

Schlup:1975:SSF

[Sch75b] W. A. Schlup. Special solutions of the Fokker–Planck transport equation for arbitrary band structures; mobility and diffusivity in a uniform field of force. Journal of Statistical Physics, 12(2): 131–143, February 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01010028.

Schwarz:1976:STA

[Sch76] Martin Schwarz, Jr. Statistical thermodynamics of an anharmonic oscillator. Journal of Statistical Physics, 15(3):255– 261, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012880.

Schulman:1977:IRC

[Sch77] L. S. Schulman. Illustration of reversed causality with remarks on experiment. Journal of Statistical Physics, 16(2):217–231, February 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01418753.

Scott:1972:DCS

[SD72] Joanna I. Scott and J. M. Deutch. Decay of correlations in spin systems. Journal of Statistical Physics, 5(3):145–157, July 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023738.

Seeberg:1972:TCF

[See72] Otto Seeberg. Time-correlation function for macroscopic observables in a classical gas. Journal of Statistical Physics, 4 (2–3):83–102, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01008545.

Scales:1973:SOG

[SG73a] Sam A. Scales and H. A. Gersch. Second-order Green’s function theory of a Heisenberg paramagnet. Journal of Statistical Physics, 7(2):95–118, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01024209.

Silevitch:1973:DFT

[SG73b] M. B. Silevitch and K. I. Golden. Dielectric formulation of test particle energy loss in a plasma. Journal of Statistical Physics, 7(1):65–87, January 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01406132.

Shankar:1972:QA

[Sha72] R. Shankar. Questions and answers. Journal of Statistical Physics, 4(1):61–77, January 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008472.

Sher:1970:QA

[She70] D. Sher. Questions and answers. Journal of Statistical Physics, 2(2):201, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009740.

Shimony:1973:CII

[Shi73] Abner Shimony. Comment on the interpretation of inductive probabilities. Journal of Statistical Physics, 9(2):187–191, October 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01016847.

Shlesinger:1974:ASC

[Shl74] Michael F. Shlesinger. Asymptotic solutions of continuous-time random walks. Journal of Statistical Physics, 10(5):421–434, May 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01008803.

Simo:1979:HPA

[Sim79] Carles Simó. On the Hénon-Pomeau attractor. Journal of Statistical Physics, 21(4):465–494, October 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009612.

Seifer:1970:SRP

[SJ70] Arnold D. Seifer and Melvin J. Jacobson. Statistics of a random plus constant vector with application to acoustic and electromagnetic-wave scattering. Journal of Statistical Physics, 2(3):279–289, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030747.

Saito:1976:NCS

[SK76] Yukio Saito and Ryogo Kubo. Nonlinearity in cooperative systems-dynamical Bethe–Ising model. Journal of Statistical Physics, 15(3):233–253, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01012879.

Skagerstam:1975:NEI

[Ska75] Bo-Sture K. Skagerstam. On the notions of entropy and information. Journal of Statistical Physics, 12(6):449–462, June 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01025833.

Stauffer:1971:PSF

[SKW71] D. Stauffer, C. S. Kiang, and G. H. Walker. Possible symmetrization of Fisher’s droplet picture near T_{c}: A bubble–droplet formula. Journal of Statistical Physics, 3(3):323–330, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011387.

Slawny:1979:LTE

[Sla79] J. Slawny. Low-temperature expansion for lattice systems with many ground states. Journal of Statistical Physics, 20(6):711– 717, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01009520.

Subramanian:1976:LTB

[SLD76] G. Subramanian, D. G. Levitt, and H. T. Davis. Long-time behavior of the angular velocity autocorrelation function for a fluid of rough spheres. Journal of Statistical Physics, 15(1):1–6, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012806.

Sur:1976:MCS

[SLM^{+}76] Amit Sur, Joel L. Lebowitz, J. Marro, M. H. Kalos, and S. Kirkpatrick. Monte Carlo studies of percolation phenomena for a simple cubic lattice. Journal of Statistical Physics, 15(5): 345–353, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020338.

Seshadri:1979:RWP

[SLS79] V. Seshadri, Katja Lindenberg, and Kurt E. Shuler. Random walks on periodic and random lattices. II. Random walk properties via generating function techniques. Journal of Statistical Physics, 21(5):517–548, November 1979. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011166.

Sari:1976:DOC

[SM76] R. R. Sari and D. Merlini. On the ν-dimensional one-component classical plasma: The thermodynamic limit problem revisited. Journal of Statistical Physics, 14(2):91–100, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011761.

Snipes:1976:SMT

[Sni76] Ray F. Snipes. Statistical mechanical theory of the effect of added nonelectrolytes on the conductances of electrolytes. Journal of Statistical Physics, 14(1):11–37, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01020132.

Sture:1974:SFR

[SNZ74] K. Sture, J. Nordholm, and Robert Zwanzig. Strategies for fluctuation renormalization in nonlinear transport theory. Journal of Statistical Physics, 11(2):143–158, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012516.

Scott:1974:DCV

[SO74] Joanna I. Scott and Irwin Oppenheim. Decay of correlations in vibrational relaxation. Journal of Statistical Physics, 11(1):51– 70, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019477.

Sonechkin:1979:BIS

[Son79] D. M. Sonechkin. On bifurcations inducing a strange attractor in the system of A. M. Oboukhov. Journal of Statistical Physics, 21(1):51–54, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011480.

Smith:1977:AIF

[SP77] E. R. Smith and J. W. Perram. Adsorption isotherms for fluids with strong interparticle repulsions. Journal of Statistical Physics, 17(2):47–60, August 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01268917.

Sposito:1974:FQS

[Spo74] Garrison Sposito. Foundations of the quantum statistics of systems in equilibrium: A measure-theoretic approach. Journal of Statistical Physics, 10(4):321–335, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012254.

Spohn:1977:DTE

[Spo77] Herbert Spohn. Derivation of the transport equation for electrons moving through random impurities. Journal of Statistical Physics, 17(6):385–412, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014347.

Spohn:1979:BEL

[Spo79] Herbert Spohn. Boltzmann equation on a lattice: Existence and uniqueness of solutions. Journal of Statistical Physics, 20 (4):463–470, April 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011782.

Slinn:1971:BDN

[SS71] W. G. N. Slinn and S. F. Shen. Brownian diffusion in a nonuniform gas. Journal of Statistical Physics, 3(3):291–305, Autumn 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011385.

Schulman:1978:SMD

[SS78] L. S. Schulman and P. E. Seiden. Statistical mechanics of a dynamical system based on Conway’s game of life. Journal of Statistical Physics, 19(3):293–314, September 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011727.

Shuler:1976:SCA

[SSL76] Kurt E. Shuler, Howard Silver, and Katja Lindenberg. A simple calculation for the average number of steps to trapping in lattice random walks. Journal of Statistical Physics, 15(5): 393–397, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020341.

Slinn:1970:KTD

[SSM70] W. G. N. Slinn, S. F. Shen, and R. M. Mazo. A kinetic theory of diffusely reflecting Brownian particles. Journal of Statistical Physics, 2(3):251–265, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01030745.

Sharma:1974:ACI

[ST74] Bhu Dev Sharma and Inder Jeet Taneja. On axiomatic characterization of information-theoretic measures. Journal of Statistical Physics, 10(4):337–346, April 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01012255.

Stark:1971:QCT

[Sta71] J. P. Stark. On the quasi-chemical theory of solutions. Journal of Statistical Physics, 3(4):395–399, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01008277.

Stauffer:1978:SAL

[Sta78] D. Stauffer. Scaling assumption for lattice animals in percolation theory. Journal of Statistical Physics, 18(2):125–136, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014304.

Stenholm:1971:SDQ

[Ste71] Stig Stenholm. On the statistical description of quantum systems. Journal of Statistical Physics, 3(4):411–416, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008279.

Shibata:1977:GSL

[STH77] Fumiaki Shibata, Yoshinori Takahashi, and Natsuki Hashitsume. A generalized stochastic Liouville equation. Non–Markovian versus memoryless master equations. Journal of Statistical Physics, 17(4):171–187, October 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01040100.

Stillinger:1976:QCL

[Sti76] Frank H. Stillinger. Quantum corrections for lattice gases. Journal of Statistical Physics, 15(5):407–421, November 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020343.

Sutugin:1971:DCC

[Sut71] A. G. Sutugin. On the dependence of the condensation coefficient on surface curvature. Journal of Statistical Physics, 3(1):91–92, Spring 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012190.

Suzuki:1976:FRE

[Suz76] Masuo Suzuki. Fluctuation, relaxation, and extensivity of macrovariables in nonequilibrium systems. Journal of Statistical Physics, 14(2):129–151, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011763.

Suzuki:1977:AFR

[Suz77a] Masuo Suzuki. Anomalous fluctuation and relaxation in unstable systems. Journal of Statistical Physics, 16(6):477–504, June 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01152285.

Suzuki:1977:STT

[Suz77b] Masuo Suzuki. Scaling theory of transient phenomena near the instability point. Journal of Statistical Physics, 16(1):11–32, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014603.

Suzuki:1979:ETL

[Suz79] Masuo Suzuki. Existence of the thermodynamic limit in nonequilibrium systems. Journal of Statistical Physics, 20(2):163–173, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011510.

Simon:1978:TMH

[SV78] I. Simon and A. Vilenkin. The theory of melting in heteropolymers. II. Correlated chains. Journal of Statistical Physics, 19 (4):405–415, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011758.

Swendsen:1973:ITL

[Swe73] Robert H. Swendsen. The interpretation of a theorem by Lebowitz. Journal of Statistical Physics, 8(3):293–294, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01012573.

Swendsen:1974:QA

[Swe74] Robert H. Swendsen. Questions and answers. Journal of Statistical Physics, 10(2):175–177, February 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01009719.

Sacco:1979:COD

[SWS79] J. E. Sacco, A. Widom, and J. B. Sokoloff. Commensurability in one-dimensional lattices at finite temperature. Journal of Statistical Physics, 21(5):497–516, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011165.

Sykes:1973:STK

[Syk73] J. Sykes. Short-time kinetic equations for hard spheres: Comparison with other theories. Journal of Statistical Physics, 8 (3):279–292, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012572.

Sykora:1974:QTB

[Sýk74] Stanislav Sýkora. Quantum theory and the Bayesian inference problems. Journal of Statistical Physics, 11(1):17–27, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01019475.

Sylvester:1976:ICS

[Syl76] Garrett S. Sylvester. Inequalities for continuous-spin Ising ferromagnets. Journal of Statistical Physics, 15(4):327–341, October 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01023057.

Shavit:1974:MPR

[SZ74] A. Shavit and Y. Zvirin. Macroscopic phenomenological relations for nonlinear processes in kinetic theory. Journal of Statistical Physics, 11(4):291–316, October 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01009790.

Singh:1976:ICS

[SZ76] S. R. Singh and M. C. Zerner. On the inversion of the classical second virial coefficient. Journal of Statistical Physics, 14 (4):351–357, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01030199.

Szasz:1978:CIN

[Szà78] Domokos Szàsz. Correlation inequalities for non-purely-ferromagnetic systems. Journal of Statistical Physics, 19(5): 453–459, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01011692.

Taneja:1974:JCD

[Tan74] Inder Jeet Taneja. A joint characterization of directed divergence, inaccuracy, and their generalizations. Journal of Statistical Physics, 11(2):169–176, August 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012518.

Taneja:1976:MII

[Tan76] Inder Jeet Taneja. On measures of information and inaccuracy. Journal of Statistical Physics, 14(3):263–270, March 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01023628.

Tate:1970:SD

[Tat70] A. Tate. Statistical dynamics. Journal of Statistical Physics, 2(1):53–59, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009710.

Tavernier:1976:CQD

[Tav76] Jean Tavernier. Classical and quantum descriptions of the Kac ring model. Journal of Statistical Physics, 14(2):101–128, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011762.

Trimble:1971:MDH

[TD71] R. H. Trimble and J. M. Deutch. Molecular derivation of the hydrodynamic equations for a binary fluid. Journal of Statistical Physics, 3(2):149–169, Summer 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01019848.

Tisza:1978:BR

[TD78] Laszlo Tisza and J. R. Dorfman. Book reviews. Journal of Statistical Physics, 18(4):409–420, April 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01106731.

Thomas:1973:PCG

[Tho73] L. C. Thomas. Probabilistic calculations on the ground state of the harmonic oscillator. Journal of Statistical Physics, 7 (4):287–294, April 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014904.

Tomita:1979:CRL

[TK79] Kazuhisa Tomita and Tohru Kai. Chaotic response of a limit cycle. Journal of Statistical Physics, 21(1):65–86, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011482.

Tribus:1972:CPJ

[TM72] Myron Tribus and Hector Motroni. Comments on the paper ‘Jaynes’s maximum entropy prescription and probability theory’. Journal of Statistical Physics, 4(2–3):227–228, March 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008552.

Tenti:1976:RTT

[TM76] G. Tenti and F. R. McCourt. Renormalized theory of the timedependent pair distribution function. I. General formulation. Journal of Statistical Physics, 15(1):59–71, July 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012809.

Tomozawa:1974:QDE

[Tom74] Yukio Tomozawa. Quasinormal distributions and expansion at the mode. Journal of Statistical Physics, 11(3):195–205, September 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01010217.

Tribus:1973:KIM

[TR73] Myron Tribus and Richard Rossi. On the Kullback information measure as a basis for information theory: Comments on a proposal by Hobson and Chang. Journal of Statistical Physics, 9(4): 331–338, December 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012165. See [HC73].

Tropper:1977:EQP

[Tro77a] M. M. Tropper. Ergodic and quasideterministic properties of finite-dimensional stochastic systems. Journal of Statistical Physics, 17(6):491–509, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014351.

Tropper:1977:EPI

[Tro77b] M. M. Tropper. Ergodic properties of infinite-dimensional stochastic systems. Journal of Statistical Physics, 17(6):511– 528, December 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014352.

Thomas:1970:BEA

[TS70] M. W. Thomas and R. F. Snider. Boltzmann equation and angular momentum conservation. Journal of Statistical Physics, 2(1):61–81, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009711.

Tsuge:1975:NHS

[TS75] Shunichi Tsug’e and Kenshi Sagara. A new hierarchy system on the basis of a ‘master’ Boltzmann equation for microscopic density. Journal of Statistical Physics, 12(5):403–425, May 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012885.

Takahashi:1976:FMS

[TS76a] Yoshinori Takahashi and Fumiaki Shibata. Fluctuations of a macro-spin in a superradiant system. Journal of Statistical Physics, 14(1):67–80, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020135.

Takahashi:1976:GPS

[TS76b] Yoshinori Takahashi and Fumiaki Shibata. Generalized phase space method in spin systems-spin coherent state representation. Journal of Statistical Physics, 14(1):49–65, January 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020134.

Tsang:1977:CTM

[Tsa77] S. K. Tsang. Corner transfer matrices of the triangular Ising model. Journal of Statistical Physics, 17(3):137–152, September 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01106913.

Tsang:1979:SLV

[Tsa79] S. K. Tsang. Square lattice variational approximations applied to the Ising model. Journal of Statistical Physics, 20(1):95–114, January 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013748.

Tunaley:1974:ASC

[Tun74] J. K. E. Tunaley. Asymptotic solutions of the continuous-time random walk model of diffusion. Journal of Statistical Physics, 11 (5):397–408, November 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01026731.

Tunaley:1975:SPA

[Tun75] J. K. E. Tunaley. Some properties of the asymptotic solutions of the Montroll–Weiss equation. Journal of Statistical Physics, 12(1):1–10, January 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01024180.

Tunaley:1976:DLR

[Tun76a] J. K. E. Tunaley. Diffusion in a lattice of randomly placed sites and application to ac conductivity. Journal of Statistical Physics, 15(2):167–177, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012036.

Tunaley:1976:MMW

[Tun76b] J. K. E. Tunaley. Moments of the Montroll–Weiss continuoustime random walk for arbitrary starting time. Journal of Statistical Physics, 14(5):461–463, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01040704.

Tunaley:1976:TCN

[Tun76c] J. K. E. Tunaley. A theory of 1/f current noise based on a random walk model. Journal of Statistical Physics, 15(2):149–156, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01012034.

Turski:1974:PCF

[Tur74] L. A. Turski. Pair correlation function for a system with velocitydependent interactions. Journal of Statistical Physics, 11(1):1– 16, July 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01019474.

Tykodi:1971:QA

[Tyk71] R. J. Tykodi. Questions and answers. Journal of Statistical Physics, 3(4):417, Winter 1971. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01008280.

Ueda:1979:RTP

[Ued79] Yoshisuke Ueda. Randomly transitional phenomena in the system governed by Duffing’s equation. Journal of Statistical Physics, 20(2):181–196, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011512.

Varley:1979:ACP

[Var79] R. L. Varley. An alternative construction of the Percus–Yevick equation based on the equilibrium BBGKY hierarchy. Journal of Statistical Physics, 21(1):87–100, July 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011483.

vanBeijeren:1979:KTH

[vBE79] H. van Beijeren and M. H. Ernst. Kinetic theory of hard spheres. Journal of Statistical Physics, 21(2):125–167, August 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01008695.

vandeBraak:1979:UBG

[vdBCGW79] H. P. van de Braak, W. J. Caspers, P. K. H. Gragert, and M. W. M. Willemse. An upper bound for the ground-state energy of an antiferromagnetic chain with next-nearest-neighbor interaction. Journal of Statistical Physics, 20(5):487–497, May 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012896.

vandeBraak:1978:ELP

[vdBCWW78] H. P. van de Braak, W. J. Caspers, F. W. Wiegel, and M. W. M. Willemse. An exact limit for the Peierls distortion energy in an antiferromagnetic chain by means of a renormalization procedure. Journal of Statistical Physics, 18(6):577–584, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014480.

vanderWaals:1979:TTC

[vdW79] J. D. van der Waals. The thermodynamic theory of capillarity under the hypothesis of a continuous variation of density. Journal of Statistical Physics, 20(2):200–244, February 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011514.

Verhagen:1976:ESC

[Ver76] A. M. W. Verhagen. An exactly soluble case of the triangular Ising model in a magnetic field. Journal of Statistical Physics, 15(3):219–231, September 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01012878.

Velarde:1974:FHB

[VH74] M. G. Velarde and E. H. Hauge. Fluctuating hydrodynamics and Brownian motion. II. Note on the slip boundary condition. Journal of Statistical Physics, 10(1):103–105, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011720.

vanHemmen:1978:GRT

[vH78] Leo van Hemmen. A generalization of Rayleigh’s theorem for the infinite harmonic crystal. Journal of Statistical Physics, 18 (1):53–58, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01014670.

Vilenkin:1978:TMH

[Vil78] A. Vilenkin. The theory of melting in heteropolymers. I. Random chains. Journal of Statistical Physics, 19(4):391–404, October 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011757.

Visscher:1978:CHI

[Vis78] P. B. Visscher. Cell hydrodynamics: An information-conserving theory. Journal of Statistical Physics, 18(1):59–89, January 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01014671.

Visscher:1979:DHD

[Vis79] P. B. Visscher. Discrete hydrodynamics: Determination of equations of motion by simulation. Journal of Statistical Physics, 20(6):629–639, June 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01009514.

vanKampen:1974:RLV

[vK74] N. G. van Kampen. Remark on the Lotka–Volterra model. Journal of Statistical Physics, 11(6):475–480, December 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008890.

vanKampen:1977:SMD

[vK77] N. G. van Kampen. A soluble model for diffusion in a bistable potential. Journal of Statistical Physics, 17(2):71–88, August 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01268919.

vanKampen:1978:BRB

[vK78] N. G. van Kampen. Book review: Statistical mechanics and statistical methods in theory and application. Journal of Statistical Physics, 19(5):557–558, November 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011700.

Vieceli:1973:ATC

[VR73a] J. J. Vieceli and H. Reiss. Application of the thermodynamics of curved boundary layers to the scaled particle theory of hard-sphere fluids. Journal of Statistical Physics, 8(4):299–307, August 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01020269.

Vieceli:1973:FAS

[VR73b] J. J. Vieceli and H. Reiss. Further application of the statistical thermodynamics of curved surfaces to scaled particle theory. Journal of Statistical Physics, 7(2):143–159, February 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01024212.

Vstovsky:1976:DBM

[Vst76] V. P. Vstovsky. On the dynamical basis of macroscopic theory. Journal of Statistical Physics, 15(2):105–121, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012030.

vanWeert:1978:DQK

[vWdB78] Ch. G. van Weert and W. P. H. de Boer. On the derivation of quantum kinetic equations. I. Collision expansions. Journal of Statistical Physics, 18(3):271–280, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01018093.

Walgraef:1976:MBQ

[Wal76] D. Walgraef. Macroscopic behavior of a quantum monomode laser. Journal of Statistical Physics, 14(5):399–416, May 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01040701.

Weiss:1975:ASP

[WD75] George H. Weiss and Menachem Dishon. Application of a singular perturbation expansion to the solution of certain Fokker– Planck equations. Journal of Statistical Physics, 13(2):145–155, August 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01221375.

Weiss:1972:SAS

[Wei72] George H. Weiss. Some approximate solutions to the discrete master equation. Journal of Statistical Physics, 6(2–3): 179–188, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01023686.

Weiss:1973:DCT

[Wei73] George H. Weiss. The diffusion constant for two-state brownons. Journal of Statistical Physics, 8(3):221–224, July 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01012566.

Weiss:1976:TSR

[Wei76] George H. Weiss. The two-state random walk. Journal of Statistical Physics, 15(2):157–165, August 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012035.

Weiss:1979:NOT

[Wei79] George H. Weiss. A note on occupation times of random walks. Journal of Statistical Physics, 21(5):609–611, November 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011172.

Wheeler:1972:QA

[Whe72] John C. Wheeler. Questions and answers. Journal of Statistical Physics, 6(2–3):223–224, November 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01023689.

Widom:1978:SIU

[Wid78] B. Widom. Structure of interfaces from uniformity of the chemical potential. Journal of Statistical Physics, 19(6):563–574, December 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01011768.

Widom:1979:MOK

[Wid79] B. Widom. In memoriam: Oscar Knefler Rice, 1903–1978. Journal of Statistical Physics, 21(3):341–344, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01011475.

Wiegel:1973:NBV

[Wie73] F. W. Wiegel. Nonclassical behavior of the van der Waals gas. Journal of Statistical Physics, 7(3):213–224, March 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01030304.

Wiegel:1975:EST

[Wie75] F. W. Wiegel. An exactly solvable two-dimensional biomembrane model. Journal of Statistical Physics, 13(6):515–530, December 1975. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01013148.

Wiegel:1979:BRB

[Wie79] F. W. Wiegel. Book review: Path integrals. Journal of Statistical Physics, 21(3):337–339, September 1979. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01011474.

Wilemski:1976:DSE

[Wil76] Gerald Wilemski. On the derivation of Smoluchowski equations with corrections in the classical theory of Brownian motion. Journal of Statistical Physics, 14(2):153–169, February 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/BF01011764.

West:1978:SNS

[WLS78] Bruce J. West, Katja Lindenberg, and Kurt E. Shuler. Studies in nonlinear stochastic processes. IV. A comparison of statistical linearization, diagrammatic expansion, and projection operator methods. Journal of Statistical Physics, 18(2):217–233, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014312.

Wintle:1978:PBP

[WP78] H. J. Wintle and P. A. Puhach. Probabilistic bond percolation in random arrays. Journal of Statistical Physics, 18(6):557–575, June 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014479.

Weiss:1976:TOS

[WR76] George H. Weiss and Robert J. Rubin. The theory of ordered spans of unrestricted random walks. Journal of Statistical Physics, 14(4):333–350, April 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01030198.

Wu:1978:PPM

[Wu78] F. Y. Wu. Percolation and the Potts model. Journal of Statistical Physics, 18(2):115–123, February 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01014303.

Yasue:1977:NDS

[Yas77] Kunio Yasue. A note on the derivation of the Schrödinger– Langevin equation. Journal of Statistical Physics, 16(1):113–116, January 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/BF01014608.

Yen:1972:GGR

[Yen72] James T. Yen. Generalization of the Gibbsian relation for dilute systems. Journal of Statistical Physics, 5(1–2):1–18, May 1972. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008368.

Yamazaki:1976:CPI

[YK76] Kikuo Yamazaki and Taro Kihara. Core potential of intermolecular forces applied to third virial coefficients and transport coefficients of polyatomic gases. Journal of Statistical Physics, 14 (6):509–519, June 1976. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01012849.

Yamada:1977:FTA

[YOK77] Tomoji Yamada, Takao Ohta, and Kyozi Kawasaki. Fieldtheoretic approach to critical dynamics far from equilibrium. Journal of Statistical Physics, 16(3):281–298, March 1977. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01020383.

Yoshitani:1970:DLC

[Yos70] Reo Yoshitani. On the detectabillity limit of coherent optical signals in thermal radiation. Journal of Statistical Physics, 2 (4):347–378, ???? 1970. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01020443.

Yuge:1977:TDS

[Yug77] Yoshio Yuge. Three-dimensional site percolation problem and effective-medium theory: A computer study. Journal of Statistical Physics, 16(4):339–348, April 1977. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01020426.

Yorke:1979:MCT

[YY79] James A. Yorke and Ellen D. Yorke. Metastable chaos: The transition to sustained chaotic behavior in the Lorenz model. Journal of Statistical Physics, 21(3):263–277, September 1979. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01011469.

Zannetti:1974:SCT

[Zan74] Marco Zannetti. Self-consistent treatment of a phase transition in a system with a vector order parameter. Journal of Statistical Physics, 10(1):1–9, January 1974. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/BF01011713.

Zwanzig:1973:NGL

[Zwa73] Robert Zwanzig. Nonlinear generalized Langevin equations. Journal of Statistical Physics, 9(3):215–220, November 1973. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01008729.